### v1

# Analysis of Shuffling Beyond Pure Local Differential Privacy

Abstract.
Shuffling is a powerful way to amplify privacy of a local randomizer in private distributed data analysis, but existing analyses mostly treat the local differential privacy (DP) parameter $\varepsilon_{0}$ as the only knob and give generic upper bounds that can be loose and do not even characterize how shuffling amplifies privacy for basic mechanisms such as the Gaussian mechanism. We revisit the privacy blanket bound of Balle et al. (the blanket divergence) and develop an asymptotic analysis that applies to a broad class of local randomizers under mild regularity assumptions, without requiring pure local DP. Our key finding is that the leading term of the blanket divergence depends on the local mechanism only through a single scalar parameter $\chi$, which we call the shuffle index. By applying this asymptotic analysis to both upper and lower bounds, we obtain a tight band for $\delta_{n}$ in the shuffled mechanism’s $(\varepsilon_{n},\delta_{n})$-DP guarantee. Moreover, we derive a simple structural necessary and sufficient condition on the local randomizer under which the blanket-divergence-based upper and lower bounds coincide asymptotically. $k$-$R$ families with $k\geq 3$ satisfy this condition, while for generalized Gaussian mechanisms the condition may not hold but the resulting band remains tight. Finally, we complement the asymptotic theory with an FFT-based algorithm for computing the blanket divergence at finite $n$, which offers rigorously controlled relative error and near-linear running time in $n$, providing a practical numerical analysis for shuffle DP.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Shuffling has emerged as one of the most powerful tools for improving the utility–privacy trade-off of differentially private (DP) distributed data analysis (Liew et al., 2022; Luo et al., 2025). In the single-message model, a population of users first applies a local randomizer to their data and then submits the resulting messages to an anonymizing shuffler. A growing body of work has demonstrated that, even when the local randomizer satisfies only relatively weak privacy guarantees, shuffling can significantly amplify privacy at the level of the aggregated mechanism, leading to substantially improved accuracy compared with pure local DP (Feldman et al., 2022; Balle et al., 2019; Erlingsson et al., 2019).
From both a theoretical and a practical perspective, two complementary ingredients are crucial. On the one hand, we need general analytic results that explain how and to what extent shuffling amplifies privacy for a given local mechanism. On the other hand, practitioners ultimately require numerically accurate finite-$n$ estimates of the privacy parameters $(\varepsilon,\delta)$ for concrete protocols. In the central-DP setting, the latter need has led to a shift from coarse composition theorems to tight numerical accountants (for example, for DP-SGD), which in turn has enabled substantially stronger utility–privacy trade-offs in practice (Gopi et al., 2021). In contrast, shuffle DP currently lacks, both from a theoretical and a practical perspective, tools that offer a comparable level of accuracy and flexibility.
Limitations of existing shuffle analyses.
Most existing theoretical and numerical analyses of shuffling focus on local randomizers that satisfy pure local DP (i.e., $(\varepsilon_{0},0)$-local DP) and treat $\varepsilon_{0}$ as the main knob controlling privacy amplification (see, e.g., (Balle et al., 2019; Erlingsson et al., 2019)). This $\varepsilon_{0}$-centric viewpoint has two important limitations.
First, the parameter $\varepsilon_{0}$ summarizes the local privacy guarantee, but it is a rather crude descriptor from the viewpoint of shuffle DP: it largely ignores the structural properties of the local mechanism that actually govern privacy amplification. In particular, existing general bounds typically do not distinguish between, say, a Laplace mechanism with a given noise scale, a $k$-randomized response ($k$-$R$) mechanism with a given $k$, or other natural families of local mechanisms. As a consequence, generic upper bounds expressed solely in terms of $\varepsilon_{0}$ can be quite loose. Indeed, several works have shown that by exploiting the specific structure of a given mechanism one can obtain significantly tighter bounds on the effect of shuffling (e.g., for $k$-$R$ one can show that privacy amplification becomes stronger as $k$ increases, even when $\varepsilon_{0}$ is kept fixed (Feldman et al., 2022; Biswas et al., 2024)). Yet, we currently lack a general understanding of what intrinsic properties of a local mechanism govern its “shuffle efficiency.”
Second, the restriction to pure local DP mechanisms excludes a wide range of natural and practically relevant local randomizers. Pure local DP is a convenient sufficient condition for privacy amplification by shuffling, but it is far from necessary. In practice, one often encounters approximate-DP local mechanisms or mechanisms that do not satisfy any DP condition. Surprisingly, even for the Gaussian mechanism, arguably one of the most prominent mechanisms in central DP, it is not known how to precisely characterize privacy amplification by shuffling. Balle et al. (Balle et al., 2019) already pointed out that the Gaussian case is technically challenging, and subsequent work by Koskela et al. (Koskela et al., 2022) observed that even numerical accounting for Gaussian shuffling is nontrivial. To date, existing results for shuffling of Gaussian mechanisms are essentially limited to lower bounds (Liew and Takahashi, 2022; Chua et al., 2024). While there are attempts to generalize shuffling guarantees to approximate local DP (i.e., $(\varepsilon_{0},\delta_{0}>0)$-local DP) in a mechanism-agnostic fashion (Feldman et al., 2022; Cheu et al., 2019), such general bounds tend to be overly pessimistic; they often show little visible amplification for moderate values of $n$, because they are driven by worst-case mechanisms that do not reflect the structure of the specific mechanism.
From a practical standpoint, this is problematic: insisting on pure local DP may force system designers to forgo local mechanisms with superior utility properties simply because their analytic and numerical tools cannot analyze them under shuffling.
The privacy blanket and blanket divergence.
A central tool in the analysis of shuffle DP is the privacy blanket introduced by Balle et al. (Balle et al., 2019). Given a local randomizer $\mathcal{R}$, the privacy blanket is a data-independent distribution $\mathcal{R}_{\mathrm{BG}}$ defined using $\mathcal{R}$. Balle et al. showed that $\delta$ of the shuffled mechanism can be upper-bounded by a divergence computed under the blanket distribution; we refer to this quantity as the blanket divergence. This blanket-divergence bound is currently the best-known general upper bound on privacy amplification by shuffling (Su et al., 2025).
However, the blanket divergence itself is still poorly understood, largely due to its analytical and numerical intractability. In particular, Balle et al. (Balle et al., 2019) mainly derive generic upper bounds via Hoeffding- and Bennett-type concentration inequalities based on the local DP parameter $\varepsilon_{0}$. Consequently, these bounds can be quite loose and fail to capture how the blanket divergence depends on the structure of the underlying local randomizer, as well as how tight the blanket-based upper bounds are compared with known lower bounds.
This is partly because Balle et al. explicitly aim to obtain finite-$n$ upper bounds. More generally, carrying out a precise finite-$n$ analysis of shuffling through the blanket divergence is intrinsically challenging: DP requires a worst-case guarantee over all neighboring datasets (i.e., taking the supremum of the blanket divergence over all neighboring input pairs $(x_{1},x_{1}^{\prime})$). Identifying the pairs that maximize the blanket divergence at finite $n$ is highly nontrivial in general. In contrast, in the $\varepsilon_{0}$-local DP setting, the single parameter $\varepsilon_{0}$ already summarizes the worst case over all neighboring input pairs, so concentration-based bounds can be expressed directly in terms of $\varepsilon_{0}$ without explicitly resolving the supremum over $(x_{1},x_{1}^{\prime})$.
Our goal and approach.
In light of the above limitations, our goal is to circumvent the apparent finite-$n$ bottlenecks of generic concentration-based analyses by taking a different route, namely an asymptotic analysis of the blanket divergence. While this sacrifices exactness at any fixed $n$, it allows us to address the conceptual and practical limitations discussed above.
On the theoretical side, we replace the $\varepsilon_{0}$-centric view by a single scalar parameter, the shuffle index $\chi$, which encodes the structure of the local randomizer. As shown in Lemma 3.2, by applying the central limit theorem (i.e., normal approximation), for a sequence $\varepsilon_{n}\to 0$ in the regime [^1] This corresponds to the moderate deviation regime required to apply the asymptotic expansion of the central limit theorem. $\varepsilon_{n}=\omega(n^{-1/2})$ and $\varepsilon_{n}=O(\sqrt{\log n/n})$, the blanket divergence for a fixed distinct pair $(x_{1},x_{1}^{\prime})$ admits the asymptotic expansion
$\mathcal{D}^{\mathrm{blanket}}=\varphi\bigl(\chi\,\varepsilon_{n}\sqrt{n}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{\varepsilon_{n}^{2}n^{3/2}}\right)\bigl(1+o(1)\bigr),$
where $\varphi$ denotes the PDF of the standard normal distribution and the shuffle index $\chi$ takes the simple form $\sqrt{\gamma}/\sigma$ associated with the pair $(x_{1},x_{1}^{\prime})$. That is, $\chi$ is the only mechanism-dependent constant appearing in the leading term. Thus $\chi$ fully governs how well a given local randomizer interacts with shuffling, in a way that is both quantitatively precise and conceptually transparent.
Moreover, by analyzing the worst-case value of the shuffle index over neighboring input pairs (i.e., the infimum of $\chi(x_{1},x_{1}^{\prime})$ over $x_{1}\neq x_{1}^{\prime}$), we obtain an asymptotic characterization of the worst-case blanket divergence and hence of the shuffled mechanism’s $(\varepsilon_{n},\frac{\alpha}{n}(1+o(1)))$ asymptotic profile (Theorem 3.3) of DP (i.e., upper bound). Because we can express not only upper bounds but also lower bounds in terms of the blanket divergence, we can quantitatively assess the tightness of the upper bound. In particular, we derive a simple structural necessary and sufficient condition on the local randomizer under which the blanket divergence yields an asymptotically optimal characterization (i.e., matches the upper bound and the lower bound) of shuffling (Theorem 3.4). This condition is satisfied, for example, by the $k$-$R$ family with $k\geq 3$. Moreover, we show that the Laplace and Gaussian mechanisms do not satisfy the condition but the upper and lower bounds differ from the true shuffled privacy $\varepsilon$ by at most about $30\%$. This shows that the blanket divergence, when analyzed through the shuffle index, provides a remarkably tight description of shuffling.
On the practical side, asymptotic expansions alone are not sufficient for numerical accounting at finite $n$. We therefore complement our asymptotic analysis with a finite-$n$ algorithm for computing the blanket divergence at the asymptotically worst-case pair $(x_{1},x_{1}^{\prime})$. Our algorithm is based on an FFT approximation of the distribution (Section 4.1), combined with explicit control of the truncation, discretization, and aliasing errors (Theorem 5). Leveraging the asymptotic expansion from Lemma 3.2, we show that the parameters can be tuned by a parameter $\eta$ so that the relative error is $O(\eta)$, while the running time scales as $\widetilde{O}(n/\eta)$ (Theorem 4.3). In this way, the asymptotic theory not only clarifies the intrinsic “shuffle efficiency” of a local randomizer via the shuffle index, but also guides the design and analysis of practical numerical accountants for shuffle DP.
Related Work and Our Contributions
Another line of work seeks to go beyond the $\varepsilon_{0}$-centric view by exploiting additional structure of the local randomizer, typically in the form of a small number of scalar parameters. Feldman et al. (Feldman et al., 2023) introduced a stronger clone paradigm and showed that, for mechanisms such as $k$-$R$, one can parameterize the effect of shuffling using $\varepsilon_{0}$ together with two additional scalar quantities, yielding substantially tighter bounds than $\varepsilon_{0}$-only analyses, although even in the $k$-$R$ case exact optimality is not established. Building on this, Wang et al. (Wang et al., 2024, 2025) refine the stronger clone-based analysis, using two scalar parameters, and provide sufficient conditions on the local randomizer under which their bounds become optimal. A subtle point is that the proof of Lemma 4.5 in (Wang et al., 2024) implicitly assumes the existence of a nontrivial post-processing map (see the errata of (Feldman et al., 2023)). Outside specific mechanism classes such as $k$-$R$ where no such map is needed (see Theorem 7.1 of (Feldman et al., 2023)), the results should therefore be interpreted with some care.
Several works have proposed numerical methods for evaluating shuffle DP guarantees. Koskela et al. (Koskela et al., 2022) use the clone paradigm together with FFT to approximate the privacy loss of shuffled mechanisms. Their approach, however, is restricted to pure local DP mechanisms and inherits the looseness of the clone paradigm, as shown by Su et al. (Su et al., 2025). Su et al. (Su et al., 2025) propose a numerical analysis of the blanket divergence based on FFT, similar in spirit to ours. However, they still assume pure local DP and do not provide rigorous control of the relative error; in addition, the running time scales as $O(n^{2})$ in the worst case when targeting a constant additive error. Biswas et al. (Biswas et al., 2024) give an exact combinatorial method for computing the privacy guarantees of shuffled $k$-$R$, but their analysis focuses on fixed neighboring database pairs rather than general $(\varepsilon,\delta)$-DP guarantees.
Our contributions.
Compared with the above works, our contributions are as follows:
  • We are the first to give a unified shuffle DP analysis that does not assume pure local DP and applies to arbitrary local randomizers under mild regularity assumptions.
  • We theoretically analyze the blanket divergence including the exact leading constant factor and summarize its leading behavior by a single shuffle index $\chi$, yielding mechanism-aware bounds and a necessary-and-sufficient optimality condition.
  • We develop an FFT-based blanket-divergence accountant with rigorous relative-error $O(\eta)$ guarantees and near-linear $\widetilde{O}(n/\eta)$ running time.
  • We empirically validate our asymptotic and FFT-based analysis, and in a distribution-estimation task we show that generalized Gaussian mechanisms achieve favorable privacy-accuracy trade-offs compared with pure local DP mechanisms.
2. Preliminaries
Notations
$\Phi$ and $\varphi$ denote the CDF and PDF of the standard normal distribution. $\Gamma(\cdot)$ denotes the Gamma function, $\Gamma(t)=\int_{0}^{\infty}u^{t-1}\mathrm{e}^{-u}\,du$ for $t>0$. $W(\cdot)$ denotes the principal branch of the Lambert $W$-function. Throughout the paper, whenever an expression can be written in terms of $\Phi$, $\Gamma$, $W$, and elementary functions, we will also refer to it as being available in closed form. $\mathbb{E}[X]$ and $\mathrm{Var}(X)$ denote the expectation and variance of a random variable $X$. $\mathbb{R}$ and $\mathbb{N}$ denote the set of real numbers and natural numbers, respectively. $\mathbf{1}\{\cdot\}$ denotes the indicator function, which equals $1$ if the condition inside the brackets is satisfied, and 0 otherwise. We use standard asymptotic notations $O(\cdot)$, $o(\cdot)$, $\Omega(\cdot)$, $\omega(\cdot)$, and $\Theta(\cdot)$.
2.1. Differential Privacy (DP)
We first recall the notion of DP and express it in a form based on the hockey-stick divergence.
Definition 2.1 (Hockey-stick divergence).
Let $\mathcal{Y}$ be an output space equipped with a reference measure (counting measure in the discrete case, Lebesgue measure in the continuous case). Let $P$ and $Q$ be probability distributions on $\mathcal{Y}$ that admit densities $p$ and $q$ with respect to this reference measure, and fix a parameter $\alpha\geq 1$. The hockey-stick divergence of $P$ from $Q$ of order $\alpha$ is
$\mathcal{D}_{\alpha}(P\|Q):=\int_{\mathcal{Y}}\bigl[p(y)-\alpha\,q(y)\bigr]_{+}\,dy,$
where $u]_{+}:=\max\{u,0\}$. In the discrete case, the integral should be read as a sum over $y\in\mathcal{Y}$.
Definition 2.2 (Differential privacy via hockey-stick divergence (Dwork and Roth, 2013; Barthe and Olmedo, 2013)).
Let $\mathcal{X}$ be an input domain and $\mathcal{Z}$ an output space. A randomized mechanism is a map $\mathcal{M}:\mathcal{X}^{n}\to\mathcal{Z}$ that takes a dataset $x_{1:n}\in\mathcal{X}^{n}$ as input and outputs a random element of $\mathcal{Z}$. For each dataset $x_{1:n}$ we write $\mathcal{M}(x_{1:n})$ for the corresponding output distribution on $\mathcal{Z}$, and we denote its density by the same symbol when convenient. Two datasets $x_{1:n},x^{\prime}_{1:n}\in\mathcal{X}^{n}$ are said to be neighboring ($x_{1:n}\simeq x^{\prime}_{1:n}$) if they differ in exactly one element. For a mechanism $\mathcal{M}$, $\varepsilon\geq 0$, and $\delta\in[0,1]$, $\mathcal{M}$ is $(\varepsilon,\delta)$-DP if and only if
$\sup_{x_{1:n}\simeq x^{\prime}_{1:n}}\mathcal{D}_{e^{\varepsilon}}\bigl(\mathcal{M}(x_{1:n})\,\big\|\,\mathcal{M}(x^{\prime}_{1:n})\bigr)\leq\delta.$
Sometimes, we say that $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-DP for $x_{1:n}\simeq x^{\prime}_{1:n}$ if $\mathcal{D}_{e^{\varepsilon}}\bigl(\mathcal{M}(x_{1:n})\,\big\|\,\mathcal{M}(x^{\prime}_{1:n})\bigr)\leq\delta$ (Sommer et al., 2019).
In the shuffle model, the mechanism $\mathcal{M}$ is typically constructed by applying a local randomizer independently to each user’s input.
Definition 2.3 (Local randomizer and blanket distribution).
Let $\mathcal{X}$ be an input space and $\mathcal{Y}$ an output space. A local randomizer is a randomized mechanism $\mathcal{R}:\mathcal{X}\to\mathcal{Y}.$ For each input $x\in\mathcal{X}$ we write $\mathcal{R}_{x}$ for the distribution of the random output $\mathcal{R}(x)$. We assume that there is a reference measure on $\mathcal{Y}$ (counting measure in the discrete case, Lebesgue measure in the continuous case) such that each $\mathcal{R}_{x}$ admits a density, and we write $\mathcal{R}_{x}(y)$ for this probability mass function or density, depending on the context. The blanket distribution $\mathcal{R}_{\mathrm{BG}}$ associated with $\mathcal{R}$ is defined for $y\in\mathcal{Y}$ by,
$\underline{\mathcal{R}}(y):=\inf_{x\in\mathcal{X}}\mathcal{R}_{x}(y),\qquad\gamma:=\int_{\mathcal{Y}}\underline{\mathcal{R}}(y)\,dy\in(0,1],\qquad\mathcal{R}_{\mathrm{BG}}(y):=\frac{1}{\gamma}\,\underline{\mathcal{R}}(y).$
where, in the purely discrete case, the integral should be read as a sum over $y\in\mathcal{Y}$. The scalar $\gamma$ is called the blanket mass of $\mathcal{R}$.
Following Balle et al. (Balle et al., 2019), we work in the single-message randomize-then-shuffle model.That is, we assume black-box access to an ideal functionality shuffler [^2] In contrast to the multi-message shuffling model, which typically requires access to a stronger ideal functionality $\mathcal{S}$ capable of splitting or decorrelating multiple messages from the same user (for example by hiding timing and grouping information), the single-message shuffling model only relies on a comparatively simple “anonymization” ideal functionality that permutes user identities. This makes the single-message model significantly easier to implement in practice, and therefore in this work we restrict attention to the single-message shuffling setting..
Definition 2.4 (Single-message shuffling and shuffled mechanism).
Let $\mathcal{R}$ be a local randomizer. Given a dataset $x_{1:n}\in\mathcal{X}^{n}$, each user $i\in[n]$ applies the local randomizer to obtain a single message $Y_{i}:=\mathcal{R}(x_{i})\in\mathcal{Y}.$ The shuffler is a randomized map $\mathcal{S}:\mathcal{Y}^{n}\to\mathcal{Y}^{n}$ which samples a permutation $\pi$ uniformly at random from the symmetric group on $[n]$ and outputs $\mathcal{S}(y_{1},\dots,y_{n}):=(y_{\pi(1)},\dots,y_{\pi(n)}).$ A shuffled mechanism is $\mathcal{M}=\mathcal{S}\circ\mathcal{R}^{n}$, which shuffles the outputs of $\mathcal{R}$.
Shuffled mechanisms are the main object of study in this paper and the following upper bound is known.
Lemma 2.5 (Lemma 5.3 of Balle et al. (Balle et al., 2019)).
Fix $\varepsilon\geq 0$, let $x_{1:n}\simeq x_{1:n}^{\prime}$ be neighboring inputs where only the first element differs (i.e., $x_{1}\neq x^{\prime}_{1}$), and $\mathcal{R}$ has blanket mass $\gamma$. Then, $\mathcal{S}\circ\mathcal{R}^{n}$ satisfies $(\varepsilon,\delta(x_{1:n},x_{1:n}^{\prime}))$-DP for $x_{1:n}$ and $x_{1:n}^{\prime}$, where
$\delta(x_{1:n},x_{1:n}^{\prime})$ $\leq\mathcal{D}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{BG}},\gamma}^{\mathrm{blanket}}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}}):=\frac{1}{n\gamma}\mathbb{E}_{\begin{subarray}{c}Y_{1:M}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{BG}}\\ M\sim\mathrm{Bin}(n,\gamma)\end{subarray}}\left[\max\left(\sum_{i=1}^{M}l_{\varepsilon}(Y_{i}),0\right)\right].$
Here, $\mathrm{Bin}$ is a binomial distribution and $l_{\varepsilon}(y):=\frac{\mathcal{R}_{x_{1}}(y)-e^{\varepsilon}\mathcal{R}_{x_{1}^{\prime}}(y)}{\mathcal{R}_{\mathrm{BG}}(y)}$ is called the privacy amplification random variable when $y=Y\sim\mathcal{R}_{\mathrm{BG}}$.
We term this upper bound blanket divergence $\mathcal{D}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\text{BG}},\gamma}^{\text{blanket}}$. We omit the parameters and simply write $\mathcal{D}^{\text{blanket}}$ when they are clear from the context. Moreover, for the same mechanism, the following lower bound, also expressed in terms of the blanket divergence, is known.
Theorem 2.6 (Theorem 14 of Su et al. (Su et al., 2025)).
For any $x,a_{1},a_{1}^{\prime}\in\mathcal{X},\varepsilon,n$, and local randomizer $\mathcal{R}$, it holds that:
$\sup_{x_{1:n}\simeq x_{1:n}^{\prime}}\mathcal{D}_{\mathrm{e}^{\varepsilon}}(\mathcal{S}\circ\mathcal{R}^{n}(x_{1:n})\|\mathcal{S}\circ\mathcal{R}^{n}(x_{1:n}^{\prime}))\geq\mathcal{D}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{x},1}^{\text{blanket}}(\mathcal{R}_{a_{1}}\|\mathcal{R}_{a_{1}^{\prime}}).$
To provide a unified treatment of the blanket divergences corresponding to the upper and lower bounds, we generalize the definition of the privacy amplification random variable as follows.
Definition 2.7 (Privacy amplification random variable).
Fix the pair $x_{1}\neq x_{1}^{\prime}\in\mathcal{X}$. Let $\mathcal{R}_{\mathrm{ref}}\in\bigl\{\mathcal{R}_{\mathrm{BG}}\bigr\}\,\cup\,\bigl\{\mathcal{R}_{x}:x\in\mathcal{X}\bigr\}$. We define the privacy amplification random variable with respect to $\mathcal{R}_{\mathrm{ref}}$ by
$l_{\varepsilon}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}}):=\frac{\mathcal{R}_{x_{1}}(Y)-e^{\varepsilon}\mathcal{R}_{x_{1}^{\prime}}(Y)}{\mathcal{R}_{\mathrm{ref}}(Y)}.$
Throughout the paper, whenever we write $l_{\varepsilon}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$ without further comment, the random variable $Y$ is understood to be drawn from the reference distribution $\mathcal{R}_{\mathrm{ref}}$. When the dependence on $(x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$ is clear from context, we simply write $l_{\varepsilon}(Y)$ or $l_{\varepsilon}$. Two choices of the reference distribution $\mathcal{R}_{\mathrm{ref}}$ will be particularly important in this work:
  • If $\mathcal{R}_{\mathrm{ref}}=\mathcal{R}_{\mathrm{BG}}$ is the blanket distribution, then $l_{\varepsilon}(Y)$ is the privacy amplification random variable underlying the blanket-divergence upper bound of Balle et al. (Balle et al., 2019).
  • If $\mathcal{R}_{\mathrm{ref}}=\mathcal{R}_{x}$ for some $x\in\mathcal{X}$, then $l_{\varepsilon}(Y)$ is the privacy amplification random variable appearing in the lower bound of Su et al. (Su et al., 2025).
2.2. Local Randomizers Beyond Pure Local DP
Most existing analyses of shuffling focus on local randomizers that satisfy pure local DP (i.e., $(\varepsilon_{0},0)$-DP at the user level). One of the main goals of this paper is to remove this restriction and develop a more general analysis that applies to arbitrary local randomizers satisfying the following mild regularity assumptions, without requiring pure local DP at the local level.
Assumption 2.8 (Regularity conditions for the local randomizer).
Consider the local randomizer $\mathcal{R}:\mathcal{X}\to\mathcal{Y}$. We assume that there exists $\rho_{0}>0$ such that, for every pair $x_{1}\neq x_{1}^{\prime}\in\mathcal{X}$ and every reference distribution $\mathcal{R}_{\mathrm{ref}}\in\bigl\{\mathcal{R}_{\mathrm{BG}}\bigr\}\,\cup\,\bigl\{\mathcal{R}_{x}:x\in\mathcal{X}\bigr\},$ the following conditions hold uniformly over all $|\varepsilon|\leq\rho_{0}$.
  1. (1)
    Uniform moment bounds. For every integer $k\geq 1$, $\mathbb{E}_{Y\sim\mathcal{R}_{\mathrm{ref}}}\bigl[\,|l_{\varepsilon}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})|^{k}\,\bigr]<\infty.$
  2. (2)
    Non-degenerate variance. The variance $\sigma^{2}:=\mathrm{Var}_{Y\sim\mathcal{R}_{\mathrm{ref}}}\bigl(l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})\bigr)$ is strictly positive, and the following conditions hold: $\int\frac{\mathcal{R}_{x_{1}}(y)^{2}}{\mathcal{R}_{\mathrm{ref}}(y)}\,dy<\infty$ and $\int\frac{\mathcal{R}_{x_{1}^{\prime}}(y)^{2}}{\mathcal{R}_{\mathrm{ref}}(y)}\,dy<\infty.$
  3. (3)
    Structural condition. Either of the following holds.
    1. (Cont)
      $l_{\varepsilon}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$ has a nontrivial absolutely continuous component with respect to Lebesgue measure whose support contains a fixed nondegenerate bounded interval $I\subset\mathbb{R}$ independent of $\varepsilon$.
    2. (Bound)
      There exists a constant $C<\infty$ such that $|l_{\varepsilon}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})|\leq C\quad\text{almost surely}.$
Assumption 2.8 is mild: any nontrivial local randomizer that satisfies pure local DP automatically satisfies these conditions, and even when pure local DP does not hold, if the privacy amplification random variable $l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$ is non-degenerate (i.e., not almost surely constant), then one can enforce the bounded case (Bound) in (3) by truncating the output space to a large but bounded interval (i.e., by slightly modifying the local randomizer).
Accordingly, we formulate our results in a mechanism-agnostic way and do not commit to a specific local randomizer. Nevertheless, it is useful to keep in mind a concrete running example satisfying Assumption 2.8 without pure local DP and truncation. In particular, we will often refer to the generalized Gaussian mechanism, which forms a flexible family of location mechanisms (parameterized by the shape $\beta$) and simultaneously encompasses the Laplace mechanism ($\beta=1$) and the Gaussian mechanism ($\beta=2$). The latter case is especially noteworthy, as the Gaussian mechanism has been explicitly highlighted in prior work (e.g., Balle et al. (Balle et al., 2019) and Koskela et al. (Koskela et al., 2022)) as technically challenging to analyze in the shuffle model.
Definition 2.9 (Generalized Gaussian local randomizer).
Let $\mathcal{X}:=[0,1]$ be the input domain and fix parameters $c>0$ and $\beta\in[1,2]$. The generalized Gaussian local randomizer $\mathcal{R}:\mathcal{X}\to\mathcal{Y}$ is defined by $\mathcal{Y}:=\mathbb{R},\mathcal{R}(x)=Y,\ \text{where }Y\sim\phi_{x}^{\beta}$ and $\phi_{x}^{\beta}$ denotes the generalized Gaussian density
$\phi_{x}^{\beta}(y):=\frac{\beta}{2c\,\Gamma(1/\beta)}\exp\!\left(-\Bigl|\frac{y-x}{c}\Bigr|^{\beta}\right),\qquad y\in\mathbb{R}.$
In other words, for each input $x\in\mathcal{X}$, the output $\mathcal{R}(x)$ is obtained by adding generalized Gaussian noise (with parameters $c,\beta$) to the input $x$.
3. Analysis of Shuffling via the Central Limit Theorem (CLT)
Lemma 2.5 and Theorem 2.6 show that, for the shuffling mechanism, the DP parameter $\delta$ is bounded above and below by expressions in terms of the blanket divergence. Consequently, by analyzing the blanket divergence, we can assess the tightness of the privacy guarantees obtained via the blanket divergence.
3.1. Analysis of Blanket Divergence
Our goal here is to analyze the blanket divergence by deriving an asymptotic expansion and identifying the leading term, including the explicit constant factor. The key observation is that the blanket divergence can be expressed in terms of sums of i.i.d. random variables (even though the common distribution depends on $n$), which allows us to apply CLT-type asymptotic expansions (see Theorem A.1 and Theorem A.3 for details). To prepare for this, we first rewrite $\mathcal{D}^{\mathrm{blanket}}$ and normalize it into forms that are more amenable to the application of these theorems.
Lemma 3.1 (Transformations for the blanket divergence).
Let $\mathcal{R}$ be a local randomizer and a fixed pair $x_{1},x_{1}^{\prime}\in\mathcal{X}$. Let the reference distribution be $\mathcal{R}_{\mathrm{ref}}\in\bigl\{\mathcal{R}_{\mathrm{BG}}\bigr\}\,\cup\,\bigl\{\mathcal{R}_{x}:x\in\mathcal{X}\bigr\}.$ For $n\in\mathbb{N}$, $\gamma\in(0,1]$, and $\varepsilon\geq 0$, define the following random variables: $Y_{1},\dots,Y_{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}$, $X_{i}:=l_{\varepsilon}(Y_{i};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$ for $i=1,\dots,n$, $B_{1},\dots,B_{n}$ i.i.d. $\mathrm{Bernoulli}(\gamma)$, independent of $(Y_{i})_{i=1}^{n}$, and $Z_{i}:=B_{i}X_{i}$ and $S_{m}:=\sum_{i=1}^{m}Z_{i}$ for $m\geq 1$. Let $\mu_{\varepsilon}:=\mathbb{E}[Z_{1}]$ and $\sigma_{\varepsilon}^{2}:=\mathrm{Var}(Z_{1})\in(0,\infty).$ For each $m\geq 1$ define the normalized sum $T_{m}:=\frac{S_{m}-m\mu_{\varepsilon}}{\sigma_{\varepsilon}\sqrt{m}}$ and its CDF $F_{m}(x):=\Pr\left[T_{m}\leq x\right]$ together with the thresholds $t_{m}:=-\frac{\mu_{\varepsilon}\sqrt{m}}{\sigma_{\varepsilon}},$ and $\eta_{m}(z):=\frac{z}{\sigma_{\varepsilon}\sqrt{m}}.$
Then the blanket divergence admits the following two equivalent representations:
(1) $\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})$ $=\frac{1}{\gamma}\,\mathbb{E}\Bigl[Z_{1}\,\Pr\bigl[S_{n-1}>-Z_{1}\bigr]\Bigr]$ $=\frac{1}{\gamma}\,\mathbb{E}\!\left[Z_{1}\,\bigl\{1-F_{n-1}\bigl(t_{n-1}-\eta_{n-1}(Z_{1})\bigr)\bigr\}\right].$
(2) $\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})$ $=\frac{1}{n\gamma}\int_{0}^{\infty}\Pr[S_{n}>u]\,du$ $=\frac{\sigma_{\varepsilon}}{\sqrt{n}\gamma}\,\int_{t_{n}}^{\infty}\bigl(1-F_{n}(x)\bigr)\,dx.$
The full proof is deferred to Appendix A.1. We retain both transformations because the conditions required to identify the leading term differ between the continuous (Cont) and bounded settings (Bound). In the continuous case of Assumption 2.8, we apply Theorem A.1 to the representation (1). In the bounded case (Bound), we instead apply Theorem A.3 to the representation (2). The resulting leading terms coincide, and we obtain the following result.
Lemma 3.2 (Asymptotic expansion of the blanket divergence).
Assume that the local randomizer $\mathcal{R}$ satisfies Assumption 2.8. Fix a pair $x_{1}\neq x_{1}^{\prime}\in\mathcal{X}$ and let $\mathcal{R}_{\mathrm{ref}}\in\bigl\{\mathcal{R}_{\mathrm{BG}}\bigr\}\,\cup\,\bigl\{\mathcal{R}_{x}:x\in\mathcal{X}\bigr\}.$ Set $\sigma^{2}:=\mathrm{Var}_{Y\sim\mathcal{R}_{\mathrm{ref}}}\bigl(l_{\varepsilon=0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})\bigr)$ and define $\chi:=\frac{\sqrt{\gamma}}{\sigma},$ where $\gamma\in(0,1]$ is the blanket mass of $\mathcal{R}$ when $\mathcal{R}_{\mathrm{ref}}=\mathcal{R}_{\mathrm{BG}}$ and $\gamma=1$ when $\mathcal{R}_{\mathrm{ref}}=\mathcal{R}_{x}$ for some $x$.
Assume that $\varepsilon_{n}\to 0$ as $n\to\infty$ and that $\varepsilon_{n}=\omega(n^{-1/2})$ and $\varepsilon_{n}=O\!\left(\sqrt{\frac{\log n}{n}}\right).$ Then, as $n\to\infty$,
(3) $\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon_{n}},n,\mathcal{R}_{\mathrm{ref}},\gamma}\;=\;\varphi\bigl(\chi\bigl(\mathrm{e}^{\varepsilon_{n}}-1\bigr)\sqrt{n}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{\bigl(\mathrm{e}^{\varepsilon_{n}}-1\bigr)^{2}n^{3/2}}\right)\bigl(1+o(1)\bigr)\;=\;\varphi\bigl(\chi\varepsilon_{n}\sqrt{n}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{\varepsilon_{n}^{2}n^{3/2}}\right)\bigl(1+o(1)\bigr).$
The $o(1)$ term depends only on the constants appearing in Assumption 2.8.
See Appendix A.2 for the full proof. The leading term in (3) is obtained by replacing $F_{n-1}$ with the standard normal CDF $\Phi$ in (1) and expanding $\frac{1}{\gamma}\,\mathbb{E}\!\left[Z_{1}\,\bigl\{1-\Phi\bigl(t_{n-1}-\eta_{n-1}(Z_{1})\bigr)\bigr\}\right]$ via Mills’ ratio and a Taylor expansion. In other words, Lemma 3.2 shows that the normal approximation not only captures the correct rate, but is asymptotically exact at the level of the leading constant. The bulk of the lengthy proof is devoted to verifying that all other error contributions are asymptotically negligible compared to this main term. A crucial structural fact is that, for any local randomizer $\mathcal{R}$, one has the identity for $\mu_{\varepsilon}$ in Lemma 3.1, $\mu_{\varepsilon}=\mathbb{E}\left[Z_{1}\right]=\gamma\bigl(1-{\rm e}^{\varepsilon}\bigr),$ so that the dependence on $\varepsilon_{n}$ in the leading term of (3) takes the particularly simple form shown above.
Our key conceptual finding is that the leading constant depends on the local randomizer only through the single parameter $\chi=\frac{\sqrt{\gamma}}{\sigma}.$ In particular, the relationship between the local randomizer $\mathcal{R}$ and the blanket divergence is fully encoded, at the level of the leading asymptotics, by the one-dimensional index $\chi$. Owing to its central role, we refer to $\chi$ as the shuffle index.
3.2. Main Result
Up to this point we have analyzed the blanket divergence for a fixed pair $(x_{1},x_{1}^{\prime})$ of inputs. In contrast, ($\varepsilon_{n},\delta_{n}$)-DP requires a guarantee that holds uniformly over all neighboring datasets $x_{1:n}\simeq x^{\prime}_{1:n}$ (i.e., $\delta_{n}=\sup_{x_{1:n}\simeq x^{\prime}_{1:n}}\mathcal{D}_{e^{\varepsilon_{n}}}(\mathcal{M}(x_{1:n})\|\mathcal{M}(x^{\prime}_{1:n}))$). Our next goal is therefore to extend the fixed-pair asymptotics of Lemma 3.2 to an asymptotically worst-case guarantee over all neighboring datasets.
Recall that Lemma 3.2 identifies, for each fixed pair $(x_{1},x_{1}^{\prime})$ and each choice of reference distribution $\mathcal{R}_{\mathrm{ref}}$, a shuffle index $\chi(x_{1},x_{1}^{\prime};\mathcal{R}_{\mathrm{ref}})$ such that the corresponding blanket divergence has the asymptotic form (3). In particular, the dependence on the local randomizer enters the leading term only through the single scalar $\chi$. Consequently, the problem of upper bounding $\delta_{n}$ reduces to understanding the worst value of $\chi$ over all pairs $(x_{1},x_{1}^{\prime})$ under $\mathcal{R}_{\mathrm{ref}}=\mathcal{R}_{\mathrm{BG}}$. Establishing the tightest lower bound on $\delta_{n}$ corresponds to finding the pair $(x_{1},x_{1}^{\prime})$ and the reference distribution $\mathcal{R}_{\mathrm{ref}}\in\{\mathcal{R}_{x}:x\in\mathcal{X}\}$ that minimize the shuffle index. We now state the result formally.
Theorem 3.3 (Shuffling delta band).
[…]
See Appendix A.3 for the full proof. Informally, Theorem 3.3 shows that for any fixed target constant $\alpha>0$ we construct a sequence $(\varepsilon_{n}^{\ast})_{n\geq 2}$ whose values lie between two explicit curves $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{up}})$ and $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})$, and for which $\delta_{n}=\frac{\alpha}{n}(1+o(1))$. [^4] More generally, one can obtain any prescribed polynomial rate $\delta_{n}=\Theta(n^{-\beta})$ with $\beta>1$ in the leading term. In this work we focus on the regime $\delta_{n}=\alpha/n(1+o(1))$ because in the DP literature, setting $\delta_{n}$ to be $\alpha/n$ is a common practice. In other words, the band between $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{up}})$ and $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})$ precisely captures $\varepsilon^{\ast}_{n}$ that realizes $\delta_{n}\approx\alpha/n$.
3.3. Shuffle Index
In this subsection we investigate shuffle indices $\chi$. Since DP is inherently concerned with worst-case guarantees, we focus on the upper bound (i.e., the curve $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})$) provided by Theorem 3.3. To quantify how tight this upper bound is, we compare it with the lower curve $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{up}})$ and examine the width of the resulting band. As $n\to\infty$, this width is asymptotically captured by the ratio $\frac{\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})}{\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{up}})}\;=\;\frac{\underline{\chi}_{\mathrm{lo}}}{\underline{\chi}_{\mathrm{up}}}\,(1+o(1)),$ so that the quantity $\underline{\chi}_{\mathrm{lo}}/\underline{\chi}_{\mathrm{up}}$ provides a convenient measure of the tightness of our blanket-divergence analysis for each local randomizer. We term $\underline{\chi}_{\mathrm{lo}}$ the lower shuffle index and $\underline{\chi}_{\mathrm{up}}$ the upper shuffle index.
Our next result characterizes when $\underline{\chi}_{\mathrm{up}}$ and $\underline{\chi}_{\mathrm{lo}}$ coincide so that the blanket-divergence analysis is asymptotically tight.
Theorem 3.4 (Necessary and sufficient condition for tightness).
Assume the local randomizer $\mathcal{R}$ with blanket mass $\gamma$ satisfies Assumption 2.8. For a pair $x_{1}\neq x_{1}^{\prime}\in\mathcal{X}$ define the disagreement set $A(x_{1},x_{1}^{\prime}):=\bigl\{y\in\mathcal{Y}:\mathcal{R}_{x_{1}}(y)\neq\mathcal{R}_{x_{1}^{\prime}}(y)\bigr\},$ that is, the set of outputs where the two local output distributions differ. Assume that the infima in the definitions of $\underline{\chi}_{\mathrm{lo}}$ and $\underline{\chi}_{\mathrm{up}}$ are attained by some pairs $(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})$ and $(x_{1}^{\mathrm{up}},x_{1}^{\prime\,\mathrm{up}})$, respectively. Then, $\underline{\chi}_{\mathrm{up}}=\underline{\chi}_{\mathrm{lo}}$ if and only if there exists a pair $(x_{1}^{\ast},x_{1}^{\prime\,\ast})$ and an input $x^{\ast}\in\mathcal{X}$ such that
$\chi_{\mathrm{lo}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\underline{\chi}_{\mathrm{lo}}\qquad\text{and}\qquad\mathcal{R}_{x^{\ast}}(y)=\gamma\mathcal{R}_{\mathrm{BG}}(y)\ \text{ for almost every }y\in A(x_{1}^{\ast},x_{1}^{\prime\,\ast}).$
In particular, when these equivalent conditions hold, the lower and upper curves in Theorem 3.3 coincide asymptotically, and there exists a unique asymptotic privacy curve $\varepsilon_{n}^{\ast}(\alpha)$ realizing $\delta_{n}=\frac{\alpha}{n}(1+o(1))$ such that $\varepsilon_{n}^{\ast}(\alpha)=\varepsilon_{n}\bigl(\alpha,\underline{\chi}_{\mathrm{lo}}\bigr)\,(1+o(1))$ as $n\to\infty.$
We defer the proof of Theorem 3.4 to Appendix A.4. In particular, the theorem applies to the case of $k$-$R$ ($k\geq 3$), and hence to mechanisms that are built on top of $k$-$R$ or share the same blanket structure, such as local hashing (Wang et al., 2017) and (for sufficiently small caps) PrivUnit (Asi et al., 2022). Thus, in many practically relevant shuffle protocols, the blanket-divergence analysis is asymptotically optimal in the sense of Theorem 3.4. For more complex families such as generalized Gaussian mechanisms, the upper and lower shuffle indices do not coincide in general, but across a wide range of parameters we empirically observe that their ratio $\underline{\chi}_{\mathrm{lo}}/\underline{\chi}_{\mathrm{up}}$ remains above $0.7$. The detailed calculations of the shuffle indices and their numerical values are deferred to Appendix B.
4. Analysis of Blanket Divergence via the Fast Fourier Transform (FFT)
The results in the previous section provide an asymptotic characterization of the blanket divergence. However, in applications one is ultimately interested in the finite-$n$ behavior of the shuffled mechanism, and asymptotic expansions alone do not directly yield numerically accurate values of the privacy parameter $\delta$ for concrete choices of $n$.
In this section, we therefore develop a complementary, non-asymptotic approach for computing the blanket divergence at finite $n$. Our goal is to design an algorithm that, for a given local randomizer $\mathcal{R}$ and parameters $(n,\varepsilon)$, produces rigorous upper and lower bounds on $\mathcal{D}^{\mathrm{blanket}}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})$ with explicitly controlled error and provable running time guarantees. The main idea is to exploit the same structure that underlies our CLT-based analysis: the blanket divergence can be expressed in terms of sums of i.i.d. privacy amplification random variables. This observation allows us to approximate the distribution of the sum by means of the Fast Fourier Transform (FFT). To obtain rigorous bounds rather than heuristic approximations, we carefully track the errors introduced by truncation of the summands, discretization on a grid, and FFT wrap-around. By combining these error bounds with the asymptotic expansion from Lemma 3.2, we can both tune the numerical parameters (window size, grid resolution, truncation level) and analyze the overall complexity of the method.
The starting point is another representation of the blanket divergence.
Lemma 4.1 (Transformation of the blanket divergence for FFT).
Let $M\sim 1+\mathrm{Bin}(n-1,\gamma)$. Then, $\mathcal{D}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}^{\mathrm{blanket}}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\Pr_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}}\\ Y_{2:n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}\end{subarray}}\left[\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\right]-\mathrm{e}^{\varepsilon}\Pr_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}^{\prime}}\\ Y_{2:n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}\end{subarray}}\left[\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\right].$
The proof is provided in Appendix A.5. Compared with the transformations in Lemma 3.1, here we work with a representation of the blanket divergence that does not involve explicit expectations or integrals, which is particularly convenient for deriving rigorous bounds on the truncation error. In particular, in our FFT-based computation we must truncate the support of the random variable in order to approximate an a priori unbounded distribution on a finite grid. This issue is especially relevant for non-pure local DP (see, e.g., the discussion of the Gaussian mechanism in Balle et al. (Balle et al., 2019)); hence this formulation is crucial for our purposes.
Su et al. (Su et al., 2025) also propose an FFT-based accountant for the blanket divergence, but their running time scales as $O(n^{2})$ in the number of users when targeting a constant additive error. In contrast, by adopting the concentration-inequality-based error control of Gopi et al. (Gopi et al., 2021) with our results in Section 3, we can tune the truncation and discretization parameters so that our algorithm runs in near-linear time $\widetilde{O}(n)$ while achieving a prescribed relative error. Algorithms 1 and 2 together form our FFT-based procedure.
4.1. FFT Algorithm
Algorithms 1 and 2 together form our FFT-based procedure. As we show in Theorem 5 below, the output of these algorithms yields rigorous upper and lower bounds on the blanket divergence.
Algorithm 1 CalculatePMF
1:
[…]
5:$h$: Bin width for discretization.
[…]
10:Let $Z^{\mathrm{tr}}$ be a random variable for a single summand conditioned on being in $s,s+w^{\mathrm{in}}]$.
11:The CDF of $Z^{\mathrm{tr}}$ is $F_{Z^{\mathrm{tr}}}(t)=\left(F_{\mathrm{ref}}(t)-F_{\mathrm{ref}}(s)\right)/p_{\mathrm{in}}$.
[…]
13:Define a discrete grid $\mathcal{G}=\{x_{j}=j\cdot h\mid j\in\mathbb{Z}\}$.
14:For each grid point $x_{j}$ in the truncated range, compute the PMF of the discretized variable $Z^{\mathrm{di}}$: $\mathbf{p}_{Z^{\mathrm{di}}}[j]\leftarrow F_{Z^{\mathrm{tr}}}(x_{j}+h/2)-F_{Z^{\mathrm{tr}}}(x_{j}-h/2)$.
[…]
18:Set $N\leftarrow\mathrm{Round}(w^{\mathrm{out}}/h)$, construct zero-padded $\mathbf{p}_{\mathrm{pad}}$ of length NN from $\mathbf{p}_{Z^{\mathrm{di}}}$, and let $\omega_{k}=\frac{2\pi k}{Nh}$.
[…]
22:return $\mathbf{p}_{S-\mu_{S^{\mathrm{di}}}}$, $\mu_{S^{\mathrm{tr}}}$
Algorithm 2 CalculateMainTerm
1:
[…]
3:$F(t)$: The CDF of $l_{\varepsilon}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$ under $\mathcal{R}_{x_{1}}$ or $\mathcal{R}_{x_{1}^{\prime}}$.
4:$c$: The constant for the discretization error bound.
[…]
6:for each grid point $z_{j}$ with corresponding PMF value $\mathbf{p}_{S-\mu_{S^{\mathrm{di}}}}[j]$ do
7:  $p_{\text{main}}\leftarrow p_{\text{main}}+\mathbf{p}_{S-\mu_{S^{\mathrm{di}}}}[j]\cdot\left(1-F(-c-z_{j}-\mu_{S^{\mathrm{tr}}})\right)$.
8:return $p_{\text{main}}$
Theorem 4.2 (Error bounds for FFT-based blanket divergence computation).
 [^5] Our error analysis assumes exact real arithmetic. While practical implementations relying on floating-point arithmetic introduce round-off errors, analyzing their propagation is orthogonal to the algorithmic errors characterized in this work.
Fix a distinct pair $x_{1},x_{1}^{\prime}\in\mathcal{X}$ and let $\mathcal{R}$ be a local randomizer. Let the reference distribution be $\mathcal{R}_{\mathrm{ref}}\in\bigl\{\mathcal{R}_{\mathrm{BG}}\bigr\}\,\cup\,\bigl\{\mathcal{R}_{x}:x\in\mathcal{X}\bigr\}.$ For $Y\sim\mathcal{R}_{\mathrm{ref}}$, denote by $F_{\mathrm{ref}}$ the CDF of the privacy amplification random variable $l_{\varepsilon}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$, and for $j\in\{x_{1},x_{1}^{\prime}\}$ denote by $F_{j}$ the CDF of $l_{\varepsilon}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$ under $Y\sim\mathcal{R}_{j}$. Let $(B_{i})_{i=2}^{n}$ be i.i.d. $\mathrm{Bernoulli}(\gamma)$ and $(Y_{i})_{i=2}^{n}$ be i.i.d. with common distribution $\mathcal{R}_{\mathrm{ref}}$, independent of $(B_{i})_{i=2}^{n}$.
Given parameters $(n,\gamma,F_{\mathrm{ref}},s,w^{\mathrm{in}},h,w^{\mathrm{out}})$, we define the truncation event $\perp:=\Bigl\{\exists i\in\{2,\dots,n\}:B_{i}=1,\ l_{\varepsilon}\bigl(Y_{i};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}}\bigr)\notin[s,s+w^{\mathrm{in}}]\Bigr\}$ and let $\bigl(\mathbf{p}_{S-\mu_{S}^{\mathrm{di}}},\mu_{S}^{\mathrm{tr}}\bigr):=\textsc{CalculatePMF}\bigl(n,\gamma,F_{\mathrm{ref}},s,w^{\mathrm{in}},h,w^{\mathrm{out}}\bigr).$ For $x\in\{x_{1},x_{1}^{\prime}\}$ and $c\geq 0$, define the main probability term
$P(c;x):=\bigl(1-\Pr[\perp]\bigr)\,\textsc{CalculateMainTerm}\bigl(\mathbf{p}_{S-\mu_{S}^{\mathrm{di}}},\mu_{S}^{\mathrm{tr}},F_{x},c\bigr).$
Then, for any $c\geq 0$,
$P(-c;x_{1})-\mathrm{e}^{\varepsilon}P(c;x_{1}^{\prime})-\delta_{\mathrm{err}}^{\mathrm{low}}(c,h,s,w^{\mathrm{in}},w^{\mathrm{out}})\leq D_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}^{\mathrm{blanket}}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})$
$\qquad\leq P(c;x_{1})-\mathrm{e}^{\varepsilon}P(-c;x_{1}^{\prime})+\delta_{\mathrm{err}}^{\mathrm{up}}(c,h,s,w^{\mathrm{in}},w^{\mathrm{out}}),$
where $\delta_{\mathrm{err}}^{\mathrm{low}}$ and $\delta_{\mathrm{err}}^{\mathrm{up}}$ are the total error terms , each of which admits a rigorous explicit bound.
See Appendix A.6 for the full proof. At a high level, Algorithm 1 (CalculatePMF) constructs, via FFT, a discrete approximation to the distribution of the sum $\sum_{i=2}^{M}l_{\varepsilon}(Y_{i}),$ where $M\sim 1+\mathrm{Bin}(n-1,\gamma)$ as in Lemma 4.1 and $Y_{i}$ are i.i.d. draws from $\mathcal{R}_{\mathrm{ref}}$. More precisely, we first truncate the privacy amplification random variable to a bounded interval, discretize it on a uniform grid, and then use the FFT to compute the probability mass function of the resulting discretized sum. This step replaces the sum of i.i.d. (continuous) variables by a finitely supported discrete distribution that approximates $\sum_{i=2}^{M}l_{\varepsilon}(Y_{i})$ up to controlled truncation, discretization and wrap-around errors.
Algorithm 2 (CalculateMainTerm) then uses this discrete distribution approximating $\sum_{i=2}^{M}l_{\varepsilon}(Y_{i})$ together with the CDFs of $l_{\varepsilon}(Y)$ under $\mathcal{R}_{x_{1}}$ and $\mathcal{R}_{x_{1}^{\prime}}$ to approximate the probabilities $\Pr\!\Bigl[\sum_{i=2}^{M}l_{\varepsilon}(Y_{i})>-\,l_{\varepsilon}(Y_{1})\Bigr].$ Because the distribution of $\sum_{i=2}^{M}l_{\varepsilon}(Y_{i})$ has been discretized, these probabilities reduce to finite sums over grid points, and are therefore straightforward to evaluate numerically.
4.2. Analysis of FFT Algorithm
Algorithms 1 and 2 depend on several numerical parameters $(c,s,w^{\mathrm{in}},h,w^{\mathrm{out}})$ that control truncation, discretization, and FFT wrap-around. Using the asymptotic characterization of the blanket divergence from Lemma 3.2, we can tune these parameters to control both the relative error and the computational cost of the FFT-based approximation. The following theorem makes this precise.
Theorem 4.3 (The Analysis of FFT-based Approximation with Tunable Relative Error).
Fix a distinct pair $x_{1},x_{1}^{\prime}\in\mathcal{X}$ and let $\mathcal{R}$ be a local randomizer that satisfies Assumption 2.8. Let the reference distribution be $\mathcal{R}_{\mathrm{ref}}\in\bigl\{\mathcal{R}_{\mathrm{BG}}\bigr\}\,\cup\,\bigl\{\mathcal{R}_{x}:x\in\mathcal{X}\bigr\}.$ Fix target knobs $\eta_{\mathrm{main}},\,\eta_{\mathrm{trunc}},\,\eta_{\mathrm{disc}},\,\eta_{\mathrm{alias}}\in\mathbb{R}.$ Let $\varepsilon_{n}\to 0$ be a sequence of privacy levels such that $\varepsilon_{n}=\Theta\bigl(\sqrt{\log n/n}\bigr).$ We denote by $L_{n}$ and $U_{n}$ the lower and upper bounds from Theorem 5, that is, $L_{n}:=P(-c;x_{1})-\mathrm{e}^{\varepsilon}P(c;x_{1}^{\prime})-\delta^{\mathrm{low}}_{\mathrm{err}}$ and $U_{n}:=P(c;x_{1})-\mathrm{e}^{\varepsilon}P(-c;x_{1}^{\prime})+\delta^{\mathrm{up}}_{\mathrm{err}}$ We define our numerical approximation as the midpoint $\widehat{D}_{n}:=\frac{U_{n}+L_{n}}{2}.$
Then we can tune parameters $(c_{n},s_{n},w_{n}^{\mathrm{in}},h_{n},w_{n}^{\mathrm{out}})$ depending on $(\eta_{\mathrm{main}},\eta_{\mathrm{trunc}},\eta_{\mathrm{disc}},\eta_{\mathrm{alias}})$ such that
$\frac{|\widehat{D}_{n}-\mathcal{D}^{\mathrm{blanket}}|}{\mathcal{D}^{\mathrm{blanket}}}=O\bigl(\eta_{\mathrm{main}}+\eta_{\mathrm{trunc}}+\eta_{\mathrm{disc}}+\eta_{\mathrm{alias}}\bigr)\qquad(n\to\infty),$
when $|\mu_{S^{\mathrm{di}}}-\mu_{S^{\mathrm{tr}}}|=O(c_{n})$ [^6] While this condition is not theoretically guaranteed for discrete mechanisms, we empirically confirmed $|\mu_{S^{\mathrm{di}}}-\mu_{S^{\mathrm{tr}}}|\leq c_{n}$ in the regime of our experiments. and $l_{\varepsilon}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$ is a non-lattice distribution [^7] A finitely supported random variable is lattice iff its support is contained in a set of the form $a+h\mathbb{Z}$ for some $a\in\mathbb{R}$ and $h>0$, that is, all pairwise differences between support points are integer multiples of a common nonzero number. For mechanisms with at least three output symbols this requires special algebraic relations, so for generic choices of parameters $l_{\varepsilon}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$ is non-lattice. The $2$-$R$ is a nongeneric exception: in that case $l_{\varepsilon}$ takes only two values and is lattice.. The overall running time of Algorithms 12 with these parameters is $O\left(\frac{n}{\eta_{\mathrm{main}}}(\log\frac{n}{\eta_{\mathrm{main}}})^{3}\right),$ where $\log\log$ factors are omitted.
See Appendix A.7 for the full proof. In particular, in the regime $\varepsilon_{n}=\Theta(\sqrt{\log n/n})$, the theorem shows that we can prescribe a target relative error level $\eta>0$ and choose the numerical knobs so that the relative error becomes $O(\eta)$, which leads to an overall running time $\widetilde{O}\!\left(\frac{n}{\eta}\right),$ up to polylogarithmic factors in $n$ and $1/\eta$. We will provide details on parameter selection and implementation in Appendix C.
5. Experiments
Refer to caption Refer to caption
(a) Upper ($\circ$) and lower ($\square$) shuffled mechanism’s DP $\delta$ for $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})$ of 3-$R$ with $\varepsilon_{0}=2$ (left) and the Gaussian mechanism with $\sigma=2$ (right), both with annotated $\varepsilon_{n}$; the dotted line indicates $\alpha/n$.
(b) Relative error as a function of the accuracy knob $\eta_{\mathrm{main}}$ with annotated runtime.
(c) Runtime as a function of the number of users $n$ with annotated relative error.
(d) Lower shuffle index as a function of the RMSE for mean estimation with annotated RMSE for variance and 33rd moment estimation.
Figure 1. Experimental validation of the FFT-based blanket-divergence accountant and the shuffle-index analysis.
Four labeled subfigures: two plots illustrating the accuracy–runtime trade-offs of the FFT-based accountant, one combined plot showing the tightness of the shuffle-index-based bands for 3-$R$ and Gaussian mechanisms, and one plot showing the dependence of the lower shuffle index on the effective per-user standard deviation.
In this section, we complement our theoretical results with five experiments. The first two visualize the privacy curves from Theorems 3.3 and 3.4. The next two illustrate the accuracy-runtime trade-offs of our FFT-based accountant from Theorems 5 and 4.3. The final experiment compares generalized Gaussian mechanisms with $k$-$R$ in a distribution-estimation task.
Unless stated otherwise, all computations are carried out in double precision and we choose $(c,s,w^{\mathrm{in}},h,w^{\mathrm{out}})$ according to Theorem 5 and set $\eta_{\mathrm{trunc}},\eta_{\mathrm{disc}},\eta_{\mathrm{alias}}=10^{-3}$. For each mechanism we first identify the asymptotically worst-case neighboring pair $(x_{1},x_{1}^{\prime})$ and the corresponding shuffle indices $\underline{\chi}_{\mathrm{lo}}$ and $\underline{\chi}_{\mathrm{up}}$ as in Theorem 3.3, and then run the FFT-based accountant at this pair.
5.1. Privacy Curves and Shuffle Index
We empirically evaluate the privacy guarantees of Theorem 3.3. For a fixed target constant $\alpha>0$ and a given local randomizer we set $\varepsilon_{n}=\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})$ and explore the behavior of the shuffled mechanism’s hockey-stick divergence $\delta_{n}$ as a function of $n$ for two mechanisms. That is, we use the blanket-based upper and lower bounds on the true shuffled hockey-stick divergence: the Balle et al. upper bound (Lemma 2.5) and the Su et al. lower bound (Theorem 2.6), both evaluated via our FFT accountant at the asymptotically worst-case pair.
The left panel of Figure 1(a) shows the results for the case of $3$-$R$ with $\varepsilon_{0}=2$. For each $n$ we plot the upper and lower bounds on the true hockey-stick divergence $\delta_{n}$ of the shuffled mechanism corresponding to $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})$. Across the entire range of $n$ the two curves are practically indistinguishable. This behavior is exactly what Theorem 3.4 predicts for $k$-$R$ with $k\geq 3$. On the theoretical side, the lower and upper shuffle indices coincide, $\underline{\chi}_{\mathrm{lo}}=\underline{\chi}_{\mathrm{up}}$, so the asymptotic band in Theorem 3.3 collapses. Empirically, we see that this collapse already occurs for moderate values of $n$, confirming that the blanket-divergence analysis is effectively optimal for this mechanism.
The right panel of Figure 1(a) reports the same experiment for the Gaussian local randomizer with noise parameter $\sigma=2$. Here the upper and lower curves form a visible but relatively narrow band. This matches our theoretical findings: the Gaussian mechanism does not satisfy the structural condition of Theorem 3.4, so the shuffle indices $\underline{\chi}_{\mathrm{lo}}$ and $\underline{\chi}_{\mathrm{up}}$ need not coincide. Nevertheless, their ratio stays close to one (above $0.7$ in our parameter range), and the band remains tight across all $n$.
5.2. Accuracy and Runtime of the FFT Accountant
We empirically validate the error and complexity guarantees of Theorems 5 and 4.3. Throughout this subsection we fix the local randomizer to be $3$-$R$ with $\varepsilon_{0}=2$ and vary either the accuracy knob $\eta_{\mathrm{main}}$ or the number of users $n$.
Figure 1(b) shows the effect of the parameter $\eta_{\mathrm{main}}$ on both accuracy and runtime. For each value of $\eta_{\mathrm{main}}$ we run the FFT-based algorithm for a fixed $(n,\varepsilon)$. Theorem 5 provides rigorous lower and upper bounds on the blanket divergence, $D_{\mathrm{low}}\;\leq\;\mathcal{D}^{\mathrm{blanket}}\;\leq\;D_{\mathrm{up}},$ so the true value is guaranteed to lie in the interval $D_{\mathrm{low}},D_{\mathrm{up}}]$. As our notion of relative error, we plot the certified relative uncertainty bandwidth $\frac{D_{\mathrm{up}}-D_{\mathrm{low}}}{D_{\mathrm{up}}}.$ As predicted by Theorem 4.3, the relative error shrinks approximately linearly as $\eta_{\mathrm{main}}$ decreases, while the runtime grows roughly like $1/\eta_{\mathrm{main}}$.
In Figure 1(c) we fix $\eta_{\mathrm{main}}$ and vary the number of users $n$. For each $n$ we again compute the interval $D_{\mathrm{low}},D_{\mathrm{up}}]$ and plot both the runtime and the relative error. The relative error remains essentially flat as $n$ increases, confirming that our parameter selection keeps the relative error under control uniformly in $n$. At the same time, the runtime grows almost linearly in $n$, with only mild logarithmic deviations. This matches the theoretical $\widetilde{O}(n/\eta)$ complexity of Theorem 4.3.
5.3. Distribution Estimation
Finally, we compare the generalized Gaussian mechanisms with $k$-$R$ in a distribution estimation task. We fix the data-generating distribution to be $X\sim\mathrm{Unif}[0,1]$ and set the number of users to $n=10^{5}$. For each local mechanism, we evaluate the empirical root-mean-squared error (RMSE) of the mean estimation. Figure 1(d) plots the RMSE on the horizontal axis and the lower shuffle index $\underline{\chi}_{\mathrm{lo}}$ on the vertical axis. For generalized Gaussian mechanisms we additionally annotate each point with the RMSEs for variance and third-moment estimation, whereas for $k$-$R$ we only plot the mean-accuracy vs. shuffle-index trade-off.
In this experiment we focus on $k$-$R$ with $k=3$, which we found to yield the best shuffle index among the values of $k$ under the same mean-estimation accuracy. Under this choice, the figure shows that, for the same level of mean RMSE, $3$-$R$ achieves the largest lower shuffle index in our parameter range, and hence provides the strongest privacy. However, a crucial caveat is that $3$-$R$ (and, more generally, $k$-$R$) does not natively support estimation of higher-order moments: estimating the variance or third central moment typically requires designing a separate mechanism (and hence additional privacy budget). In contrast, generalized Gaussian mechanisms simultaneously support estimation of multiple moments from the same shuffled output. Figure 1(d) reveals that the generalized Gaussian family exhibits a nontrivial dependence on the shape parameter $\beta$: the value of $\beta$ that optimizes the shuffle index varies with the target mean-estimation accuracy. In particular, in more noisy regimes (larger local noise), the Gaussian mechanism ($\beta=2$) attains the largest lower shuffle index among the generalized Gaussian family. Moreover, over a wide range of mean-accuracy levels, the Gaussian mechanism also achieves the smallest RMSE for variance and third-moment estimation among the generalized Gaussian mechanisms, as indicated by the annotations. Taken together, these results suggest that generalized Gaussian mechanisms ($\beta>1$) provide a better utility-privacy trade-off than pure local DP mechanisms. Importantly, this kind of optimization is made possible by our unified framework beyond pure local DP.
6. Limitations and Conclusion
Our analysis has a number of limitations that we highlight as directions for future work.
First, when studying the generalized Gaussian mechanism we restricted attention to the one-dimensional case. This restriction is not conceptually essential. Our asymptotic and FFT-based arguments extend in principle to higher-dimensional local randomizers; however, in high dimensions the blanket mass $\gamma$ of the Gaussian mechanism decays exponentially in the dimension. As a consequence, the lower shuffle index $\underline{\chi}_{\mathrm{lo}}$ shrinks rapidly with the dimension, and the blanket-divergence upper bound becomes essentially vacuous for practically relevant values of $n$. This is particularly problematic when analyzing locally private federated learning or shuffled DP-SGD through our framework. At the same time, the upper shuffle index $\underline{\chi}_{\mathrm{up}}$ exhibits essentially the same behavior as in the one-dimensional case, so that the gap between upper and lower shuffle indices widens with the dimension. Closing this gap in high dimensions appears to require techniques that go beyond the blanket divergence, which remains an open problem.
Second, our moderate-deviation guarantees are formulated in terms of an asymptotically worst-case neighboring pair, that is, a pair $(x_{1}^{\ast},x_{1}^{\prime\ast})$ that attains the worst-case shuffle index in the limit $n\to\infty$. While this is natural from the point of view of our CLT-based analysis, it does not rule out the possibility that, for finite $n$, special pairs might induce slightly worse $(\varepsilon,\delta)$ trade-offs than those induced by the asymptotic worst-case pair. Establishing conditions under which the asymptotically worst-case pair is also worst-case (or near worst-case) at finite $n$ is a practically relevant problem that we leave for future work.
Despite these limitations, our work is, to the best of our knowledge, the first to go beyond the pure local DP viewpoint and organize privacy amplification by shuffling via the shuffle index. Our framework provides a unified way to reason about shuffle amplification and to obtain tight $(\varepsilon,\delta)$ guarantees. We believe that these conceptual and technical advantages are valuable both for the theoretical understanding of shuffle DP and for its practical deployment in real systems.
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.
Appendix A Missing Proofs
A.1. Proofs of Lemma 3.1
See 3.1
Proof.
We start from Balle et al.’s representation (Lemma 2.5), adapted to our notation. Let $Y_{1},\dots,Y_{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}$, let $X_{i}:=l_{\varepsilon}(Y_{i};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$, and let $B_{1},\dots,B_{n}$ be i.i.d. $\mathrm{Bernoulli}(\gamma)$, independent of $(Y_{i})_{i=1}^{n}$. Set $Z_{i}:=B_{i}X_{i}$ and
$S_{n}:=\sum_{i=1}^{n}Z_{i}.$
Then,
(4) $\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\frac{1}{n\gamma}\,\mathbb{E}\bigl[(S_{n})_{+}\bigr],\qquad(S_{n})_{+}:=\max\{S_{n},0\}.$
Indeed, if $M\sim\mathrm{Bin}(n,\gamma)$ and $(Y_{i})_{i\geq 1}$ are i.i.d. $\mathcal{R}_{\mathrm{ref}}$ then $\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})$ has the same distribution as $\sum_{i=1}^{n}Z_{i}$, so (4) is equivalent to Balle et al.’s original formula.
We now show that (4) implies the two representations in the statement.
Proof of (1). We have
$(S_{n})_{+}=S_{n}\mathbf{1}\{S_{n}>0\}=\Bigl(\sum_{i=1}^{n}Z_{i}\Bigr)\mathbf{1}\{S_{n}>0\}=\sum_{i=1}^{n}Z_{i}\,\mathbf{1}\{S_{n}>0\}.$
Taking expectations and using linearity of expectation,
$\mathbb{E}[(S_{n})_{+}]=\sum_{i=1}^{n}\mathbb{E}\bigl[Z_{i}\,\mathbf{1}\{S_{n}>0\}\bigr].$
Since $(Z_{i})_{i=1}^{n}$ are i.i.d., all summands are equal, hence
$\mathbb{E}[(S_{n})_{+}]=n\,\mathbb{E}\bigl[Z_{1}\,\mathbf{1}\{S_{n}>0\}\bigr].$
Substituting this into (4) yields
(5) $\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\frac{1}{\gamma}\,\mathbb{E}\bigl[Z_{1}\,\mathbf{1}\{S_{n}>0\}\bigr].$
Next, decompose the sum $S_{n}$ as
$S_{n}=Z_{1}+\sum_{i=2}^{n}Z_{i}.$
Define
$S_{2:n}:=\sum_{i=2}^{n}Z_{i}.$
Then
$\{S_{n}>0\}=\{Z_{1}+S_{2:n}>0\}.$
Using the definition of conditional expectation and the fact that $Z_{1}$ is independent of $S_{2:n}$,
$\mathbb{E}\bigl[Z_{1}\mathbf{1}\{S_{n}>0\}\bigr]$ $=\mathbb{E}\Bigl[Z_{1}\mathbf{1}\{Z_{1}+S_{2:n}>0\}\Bigr]$
$=\mathbb{E}\Bigl[Z_{1}\,\Pr\bigl[Z_{1}+S_{2:n}>0\,\big|\,Z_{1}\bigr]\Bigr]$
$=\mathbb{E}\Bigl[Z_{1}\,\Pr\bigl[S_{2:n}>-Z_{1}\bigr]\Bigr].$
Since $(Z_{i})_{i=1}^{n}$ are i.i.d., we have
$S_{2:n}=\sum_{i=2}^{n}Z_{i}\;\stackrel{{\scriptstyle d}}{{=}}\;\sum_{i=1}^{n-1}Z_{i}=:S_{n-1},$
and $S_{2:n}$ is independent of $Z_{1}$. Hence
$\Pr\bigl[S_{2:n}>-Z_{1}\bigr]=\Pr\bigl[S_{n-1}>-Z_{1}\bigr],$
and therefore
(6) $\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\frac{1}{\gamma}\,\mathbb{E}\bigl[Z_{1}\,\Pr\bigl[S_{n-1}>-Z_{1}\bigr]\bigr].$
By definition,
$\mu_{\varepsilon}:=\mathbb{E}[Z_{1}],\qquad\sigma_{\varepsilon}^{2}:=\mathrm{Var}(Z_{1})\in(0,\infty),$
and
$T_{n-1}:=\frac{S_{n-1}-(n-1)\mu_{\varepsilon}}{\sigma_{\varepsilon}\sqrt{n-1}},\qquad F_{n-1}(x):=\Pr\left[T_{n-1}\leq x\right].$
We can rewrite $S_{n-1}$ as
$S_{n-1}=(n-1)\mu_{\varepsilon}+\sigma_{\varepsilon}\sqrt{n-1}\,T_{n-1}.$
For any real $z$,
$\{S_{n-1}>-z\}$ $\iff\bigl\{(n-1)\mu_{\varepsilon}+\sigma_{\varepsilon}\sqrt{n-1}\,T_{n-1}>-z\bigr\}$
$\iff\left\{T_{n-1}>\frac{-z-(n-1)\mu_{\varepsilon}}{\sigma_{\varepsilon}\sqrt{n-1}}\right\}.$
Define
$t_{n-1}:=-\frac{\mu_{\varepsilon}\sqrt{n-1}}{\sigma_{\varepsilon}},\qquad\eta_{n-1}(z):=\frac{z}{\sigma_{\varepsilon}\sqrt{n-1}},$
so that
$\frac{-z-(n-1)\mu_{\varepsilon}}{\sigma_{\varepsilon}\sqrt{n-1}}=t_{n-1}-\eta_{n-1}(z).$
Therefore,
$\Pr\left[S_{n-1}>-z\right]=\Pr\left(T_{n-1}>t_{n-1}-\eta_{n-1}(z)\right)=1-F_{n-1}\left(t_{n-1}-\eta_{n-1}(z)\right).$
Applying this with $z=Z_{1}$ in (6), we obtain
$\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})$ $=\frac{1}{\gamma}\,\mathbb{E}\Bigl[Z_{1}\,\bigl\{1-F_{n-1}\bigl(t_{n-1}-\eta_{n-1}(Z_{1})\bigr)\bigr\}\Bigr],$
which is exactly (1).
Proof of (2). For any real number $x$ we have the identity
$x_{+}=\max\{x,0\}=\int_{0}^{\infty}\mathbf{1}\{x>u\}\,du.$
Applying this with $x=S_{n}$ and using Tonelli’s theorem (Fubini’s theorem for nonnegative functions) gives
$\mathbb{E}[(S_{n})_{+}]$ $=\mathbb{E}\Bigl[\int_{0}^{\infty}\mathbf{1}\{S_{n}>u\}\,du\Bigr]$
$=\int_{0}^{\infty}\Pr\left[S_{n}>u\right]\,du.$
Substituting into (4),
(7) $\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\frac{1}{n\gamma}\int_{0}^{\infty}\Pr\left[S_{n}>u\right]\,du.$
We now express the tail probability in terms of the normalized sum $T_{n}$. By definition,
$T_{n}:=\frac{S_{n}-n\mu_{\varepsilon}}{\sigma_{\varepsilon}\sqrt{n}},\qquad F_{n}(x):=\Pr\left[T_{n}\leq x\right],$
so that
$S_{n}=n\mu_{\varepsilon}+\sigma_{\varepsilon}\sqrt{n}\,T_{n}.$
For any $u\in\mathbb{R}$,
$\Pr\left[S_{n}>u\right]$ $=\Pr\Bigl[n\mu_{\varepsilon}+\sigma_{\varepsilon}\sqrt{n}\,T_{n}>u\Bigr]$
$=\Pr\Bigl[T_{n}>\frac{u-n\mu_{\varepsilon}}{\sigma_{\varepsilon}\sqrt{n}}\Bigr]$
$=1-F_{n}\Bigl(\frac{u-n\mu_{\varepsilon}}{\sigma_{\varepsilon}\sqrt{n}}\Bigr).$
Perform the change of variables
$x=\frac{u-n\mu_{\varepsilon}}{\sigma_{\varepsilon}\sqrt{n}},\qquad u=n\mu_{\varepsilon}+\sigma_{\varepsilon}\sqrt{n}\,x,\qquad du=\sigma_{\varepsilon}\sqrt{n}\,dx.$
When $u=0$, the corresponding lower limit is
$x=\frac{0-n\mu_{\varepsilon}}{\sigma_{\varepsilon}\sqrt{n}}=-\frac{\mu_{\varepsilon}\sqrt{n}}{\sigma_{\varepsilon}}=:t_{n}.$
Thus
$\int_{0}^{\infty}\Pr[S_{n}>u]\,du$ $=\int_{0}^{\infty}\Bigl\{1-F_{n}\Bigl(\frac{u-n\mu_{\varepsilon}}{\sigma_{\varepsilon}\sqrt{n}}\Bigr)\Bigr\}\,du$
$=\int_{t_{n}}^{\infty}\bigl(1-F_{n}(x)\bigr)\,\sigma_{\varepsilon}\sqrt{n}\,dx$
$=\sigma_{\varepsilon}\sqrt{n}\int_{t_{n}}^{\infty}\bigl(1-F_{n}(x)\bigr)\,dx.$
Substituting this into (7) gives
$\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})$ $=\frac{1}{n\gamma}\,\sigma_{\varepsilon}\sqrt{n}\int_{t_{n}}^{\infty}\bigl(1-F_{n}(x)\bigr)\,dx$
$=\frac{\sigma_{\varepsilon}}{\sqrt{n}\,\gamma}\,\int_{t_{n}}^{\infty}\bigl(1-F_{n}(x)\bigr)\,dx,$
which is (2). This completes the proof. ∎
A.2. The proof of Lemma 3.2
See 3.2
We analyze the blanket divergence by deriving an asymptotic expansion based on central limit theorems (CLT). The general CLT-based results used in our analysis are stated below in a form tailored to our setting, with the level of abstraction kept to the minimum necessary. In particular, we treat the continuous case (Cont) and the bounded case (Bound) separately, using a triangular-array Edgeworth expansion in the former and a triangular-array moderate deviation theorem in the latter.
Theorem A.1 (Triangular-array Edgeworth expansion after Angst–Poly (Angst and Poly, 2017)).
Let $(X_{i,n})_{1\leq i\leq n,\,n\geq 1}$ be a triangular array of real-valued random variables such that, for each fixed $n\geq 1$, the variables $X_{1,n},\dots,X_{n,n}$ are independent and identically distributed (row-wise i.i.d.), while the common distribution may depend on $n$. For each $n$ set $S_{n}:=\sum_{i=1}^{n}X_{i,n},$ $B_{n}^{2}:=\mathrm{Var}(S_{n}),$ and $F_{n}(x):=\Pr\left[\frac{S_{n}}{B_{n}}\leq x\right].$
Fix an integer $s\geq 3$. Assume: (i) $\forall r\in\mathbb{N},\sup_{n}\mathbb{E}[|X_{1,n}|^{r}]<\infty$. (ii) For all $n$, $\mathbb{E}[X_{1,n}]=0$ and $\inf_{n}\mathrm{Var}(X_{1,n})\geq C>0$. (iii) $X_{1,n}$ satisfies Cramér’s condition. That is, $\sup_{n}\limsup_{|t|\to\infty}\bigl|\mathbb{E}\big[e^{it\cdot X_{1,n}}\big]\bigr|<1.$ For $r\geq 3$ denote by $\kappa_{r,n}:=\frac{1}{B_{n}^{r}}\sum_{i=1}^{n}\kappa_{r}(X_{i,n})$ the normalized cumulants of $S_{n}/B_{n}$, and let $P_{1,n},\dots,P_{s-2,n}$ be polynomials in $x$, whose coefficients are universal polynomials in $\kappa_{3,n},\dots,\kappa_{s,n}$.
Then we have the Edgeworth expansion
$\sup_{x\in\mathbb{R}}\biggl|F_{n}(x)-\Phi(x)-\sum_{j=1}^{s-2}P_{j,n}(x)\,\varphi(x)\biggr|=o\bigl(n^{-(s-2)/2}\bigr),\qquad n\to\infty.$
Remark A.2.
Theorem A.1 is obtained by specializing the general triangular-array Edgeworth expansion of Angst and Poly (Angst and Poly, 2017, Thm. 4.3) to the one-dimensional row-wise i.i.d. case. Their theorem covers much more general triangular arrays (including non-identically distributed and vector-valued arrays); in this paper we only need the resulting triangular-array extension of the classical Cramér-type Edgeworth expansion of Petrov (Petrov, 1975, Ch. VI, Thms. 3–4). Compared with the hypotheses of (Angst and Poly, 2017, Thm. 4.3), we adopt stronger assumptions in the spirit of Petrov: we require the existence of moments of all integer orders and replace the weak mean Cramér condition by the classical Cramér condition for the common distribution of $X_{1,n}$. When the distribution of $X_{1,n}$ does not depend on $n$, Theorem A.1 reduces to Petrov’s one-dimensional Edgeworth expansion for sums of i.i.d. random variables.
Theorem A.3 (Triangular-array moderate deviations after Slastnikov (Slastnikov, 1979)).
Let $(X_{i,n})_{1\leq i\leq n,\,n\geq 1}$ be a triangular array of real-valued random variables such that, for each fixed $n\geq 1$, the variables $X_{1,n},\dots,X_{n,n}$ are independent and identically distributed (row-wise i.i.d.), while the common distribution may depend on $n$. For each $n$ set $S_{n}:=\sum_{i=1}^{n}X_{i,n},$ $B_{n}^{2}:=\mathrm{Var}(S_{n}),$ and $F_{n}(x):=\Pr\!\left[\frac{S_{n}}{B_{n}}\leq x\right].$ Assume: (i) $\mathbb{E}[X_{1,n}]=0$ for all $n\geq 1$; (ii) there exists a constant $\sigma_{0}>0$ such that $\mathrm{Var}(X_{1,n})\geq\sigma_{0}^{2}$ for all $n\geq 1$ (equivalently, $B_{n}^{2}=\Theta(n)$); (iii) the array is uniformly bounded: there exists $K<\infty$ such that $|X_{1,n}|\leq K$ almost surely for all $n\geq 1$. Then for every fixed $c>0$ we have, as $n\to\infty$,
$\sup_{0\leq x\leq c\sqrt{\log n}}\left|\frac{1-F_{n}(x)}{1-\Phi(x)}-1\right|\;\longrightarrow\;0.$
Remark A.4.
Theorem A.3 is a specialization of Slastnikov’s moderate deviation theorem for triangular arrays (Slastnikov, 1979, Thm. 1.1) to the row-wise i.i.d. case, under the uniform boundedness assumption (iii). In Slastnikov’s notation, assumption (iii) implies his moment condition for any fixed $c>0$, and (i)–(iii) together yield $B_{n}^{2}=\mathrm{Var}(S_{n})=\Theta(n)$. Hence his conclusion $1-F_{n}(x)\sim 1-\Phi(x)$ holds uniformly in the moderate deviation region $0\leq x\leq c\sqrt{\log B_{n}}$ for any fixed $c>0$. Since $\log B_{n}=\tfrac{1}{2}\log n+O(1)$, this is equivalent to uniformity on $0\leq x\leq c\sqrt{\log n}$ as stated in Theorem A.3.
Proof.
We first explain the structure of the proof. The argument proceeds in three steps.
First, in order to work with bounded summands that simplify the analysis, we introduce a mean-preserving truncation of the privacy amplification variable in the (Cont) case and define the corresponding truncated blanket divergence. We show that the truncation error is negligible compared to the main term; see Lemma A.6.
Second, for the truncated variables we analyze the expectation appearing in the transformation formula of Lemma 3.1. In the continuous case (Cont) we apply a triangular-array Edgeworth expansion (Theorem A.1) to the normalized sums and derive a precise asymptotic expression for the expectation; this is summarized in Lemma A.7. In the discrete bounded case (Bound) we instead rely on Slastnikov’s triangular-array moderate deviation theorem (Theorem A.3); see Corollary A.8.
Finally, we express the resulting asymptotic formula in terms of the variance $\sigma^{2}=\mathrm{Var}(l_{0}(Y))$ of the underlying privacy amplification variable and the parameter $\chi=\sqrt{\gamma}/\sigma.$ This is achieved by showing that $\sigma_{\varepsilon_{n}}\to\sqrt{\gamma}\,\sigma$ and that the Gaussian tail factor $\varphi(t_{n})$ can be reparametrized as $\varphi(\chi(\mathrm{e}^{\varepsilon_{n}}-1)\sqrt{n})$ up to a relative $o(1)$ error. Putting the three steps together yields the claimed asymptotic expansion in Lemma 3.2.
Definition A.5 (Truncated blanket divergence).
 [^8] The truncation step is not essential for the validity of the asymptotic expansion. One could work directly with $l_{\varepsilon_{n}}(Y)$ by splitting expectations over the events $\{|l_{\varepsilon_{n}}(Y)|\leq\tau_{n}\}$ and its complement, and controlling the tail contribution via higher moments and Markov’s inequality. We introduce the truncated variables $X_{i,n}^{\tau_{n}}$ only for technical convenience: the uniform bound $|X_{i,n}^{\tau_{n}}|\leq C\tau_{n}$ allows us to treat the error terms in the Edgeworth and Taylor expansions in a simpler and more uniform way.
Let $(\varepsilon_{n})_{n\geq 1}$ be a sequence of positive numbers and $(\tau_{n})_{n\geq 1}$ a sequence of truncation levels with $\tau_{n}>0$. For each $n\geq 1$, let $Y_{1},Y_{2},\dots$ be i.i.d. with distribution $\mathcal{R}_{\mathrm{ref}}$ and define
$X_{i,n}:=l_{\varepsilon_{n}}(Y_{i};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}}),\qquad\mu_{n}:=\mathbb{E}[X_{1,n}]=1-\mathrm{e}^{\varepsilon_{n}}.$
Let $B_{1},B_{2},\dots$ be i.i.d. $\mathrm{Bernoulli}(\gamma)$, independent of $(Y_{i})_{i\geq 1}$, and set
$Z_{i,n}:=B_{i}X_{i,n},\qquad i\geq 1.$
We introduce the truncated random variable at level $\tau_{n}$ for $l_{\varepsilon_{n}}$ where $Y\sim\mathcal{R}_{\mathrm{ref}}$ by mean preserving truncation:
$\tilde{l}_{\varepsilon_{n},\tau_{n}}(Y):=l_{\varepsilon_{n}}(Y)\,\mathbf{1}\bigl\{|l_{\varepsilon_{n}}(Y)|\leq\tau_{n}\bigr\},$
$l_{\varepsilon_{n},\tau_{n}}(Y):=\tilde{l}_{\varepsilon_{n},\tau_{n}}(Y)-\mathbb{E}\bigl[\tilde{l}_{\varepsilon_{n},\tau_{n}}(Y_{1})\bigr]+\mu_{n},$
and write
$X_{i,n}^{\tau_{n}}:=l_{\varepsilon_{n},\tau_{n}}(Y_{i}),\qquad Z_{i,n}^{\tau_{n}}:=B_{i}X_{i,n}^{\tau_{n}},\qquad i\geq 1.$
For each $n$, define the partial sums
$S_{n}:=\sum_{i=1}^{n}Z_{i,n},\qquad S_{n}^{\tau_{n}}:=\sum_{i=1}^{n}Z_{i,n}^{\tau_{n}}.$
Using the Bernoulli–selector representation of the blanket divergence, we write
$\mathcal{D}^{\mathrm{blanket}}:=\frac{1}{n\gamma}\,\mathbb{E}\bigl[(S_{n})_{+}\bigr],\qquad\mathcal{D}^{\mathrm{tblanket}}:=\frac{1}{n\gamma}\,\mathbb{E}\bigl[(S_{n}^{\tau_{n}})_{+}\bigr].$
We call $\mathcal{D}^{\mathrm{tblanket}}$ the truncated blanket divergence at level $\tau_{n}$.
Lemma A.6 (Truncation error for the blanket divergence).
Assume that the local randomizer $\mathcal{R}$ satisfies Assumption 2.8 with (Cont). Then, for every integer $p\geq 2$ there exists a constant $C<\infty$ such that
(8) $\bigl|\mathcal{D}^{\mathrm{blanket}}-\mathcal{D}^{\mathrm{tblanket}}\bigr|\;\leq\;\frac{C}{\tau_{n}^{p-1}},\qquad n\geq 1.$
In particular, if $\tau_{n}=n^{\alpha}$ for some $0<\alpha<1/2$, then
$\bigl|\mathcal{D}^{\mathrm{blanket}}-\mathcal{D}^{\mathrm{tblanket}}\bigr|=O\bigl(n^{-\alpha(p-1)}\bigr),n\to\infty.$
so by choosing $p$ sufficiently large the truncation error can be made smaller than any prescribed polynomial order in $n$.
Moreover, the truncated random variables $Z_{i,n}^{\tau_{n}}:=B_{i}l_{\varepsilon_{n},\tau_{n}}(Y_{i})$ also satisfy the conditions of Assumption 2.8 (possibly with different constants in that assumption) uniformly in $n$. In addition, there exist constants $C<\infty$ and $n_{0}\in\mathbb{N}$ such that
$|Z_{i,n}^{\tau_{n}}|\leq C\tau_{n}\qquad\text{almost surely for all }n\geq n_{0},\;i\geq 1.$
Proof.
Write
$\Delta_{i,n}:=Z_{i,n}-Z_{i,n}^{\tau_{n}}=B_{i}\bigl(X_{i,n}-X_{i,n}^{\tau_{n}}\bigr),\qquad i\geq 1.$
By the definition of the blanket divergence and the truncated blanket divergence,
$\mathcal{D}^{\mathrm{blanket}}-\mathcal{D}^{\mathrm{tblanket}}=\frac{1}{n\gamma}\,\mathbb{E}\bigl[(S_{n})_{+}-(S_{n}^{\tau_{n}})_{+}\bigr].$
Since $x\mapsto x_{+}$ is $1$-Lipschitz, we have
$\bigl|(S_{n})_{+}-(S_{n}^{\tau_{n}})_{+}\bigr|\leq\bigl|S_{n}-S_{n}^{\tau_{n}}\bigr|=\biggl|\sum_{i=1}^{n}\Delta_{i,n}\biggr|.$
Therefore
(9) $\bigl|\mathcal{D}^{\mathrm{blanket}}-\mathcal{D}^{\mathrm{tblanket}}\bigr|\leq\frac{1}{n\gamma}\,\mathbb{E}\biggl|\sum_{i=1}^{n}\Delta_{i,n}\biggr|.$
By the triangle inequality,
$\mathbb{E}\biggl|\sum_{i=1}^{n}\Delta_{i,n}\biggr|\leq\sum_{i=1}^{n}\mathbb{E}\bigl[|\Delta_{i,n}|\bigr]=n\,\mathbb{E}\bigl[|\Delta_{1,n}|\bigr].$
Substituting this into (9) yields
$\bigl|\mathcal{D}^{\mathrm{blanket}}-\mathcal{D}^{\mathrm{tblanket}}\bigr|\leq\frac{1}{\gamma}\,\mathbb{E}\bigl[|\Delta_{1,n}|\bigr].$
It remains to estimate $\mathbb{E}[|\Delta_{1,n}|]$. By the definition of the mean-preserving truncation,
$X_{1,n}^{\tau_{n}}=X_{1,n}\mathbf{1}\{|X_{1,n}|\leq\tau_{n}\}-\mathbb{E}\bigl[X_{1,n}\mathbf{1}\{|X_{1,n}|\leq\tau_{n}\}\bigr]+\mu_{n},$
and since $\mathbb{E}[X_{1,n}]=\mu_{n}$ we obtain, after rearranging,
$X_{1,n}^{\tau_{n}}=X_{1,n}\mathbf{1}\{|X_{1,n}|\leq\tau_{n}\}+\mathbb{E}\bigl[X_{1,n}\mathbf{1}\{|X_{1,n}|>\tau_{n}\}\bigr].$
Hence
$X_{1,n}-X_{1,n}^{\tau_{n}}=X_{1,n}\mathbf{1}\{|X_{1,n}|>\tau_{n}\}-\mathbb{E}\bigl[X_{1,n}\mathbf{1}\{|X_{1,n}|>\tau_{n}\}\bigr].$
Using $|B_{1}|\leq 1$, we get
$|\Delta_{1,n}|=|B_{1}|\;|X_{1,n}-X_{1,n}^{\tau_{n}}|\leq|X_{1,n}|\mathbf{1}\{|X_{1,n}|>\tau_{n}\}+\mathbb{E}\bigl[|X_{1,n}|\mathbf{1}\{|X_{1,n}|>\tau_{n}\}\bigr].$
Taking expectations gives
$\mathbb{E}\bigl[|\Delta_{1,n}|\bigr]\leq 2\,\mathbb{E}\bigl[|X_{1,n}|\mathbf{1}\{|X_{1,n}|>\tau_{n}\}\bigr].$
Let $p\geq 2$ be fixed. By Markov’s inequality and the uniform moment bound in Assumption 2.8, we have
$\mathbb{E}\bigl[|X_{1,n}|\mathbf{1}\{|X_{1,n}|>\tau_{n}\}\bigr]\leq\frac{\mathbb{E}[|X_{1,n}|^{p}]}{\tau_{n}^{p-1}}\leq\frac{C_{p}}{\tau_{n}^{p-1}},$
for some constant $C_{p}<\infty$ independent of $n$. Combining the last three displays yields
$\mathbb{E}\bigl[|\Delta_{1,n}|\bigr]\leq\frac{2C_{p}}{\tau_{n}^{p-1}}.$
Recalling that $\bigl|\mathcal{D}^{\mathrm{blanket}}-\mathcal{D}^{\mathrm{tblanket}}\bigr|\leq\gamma^{-1}\mathbb{E}[|\Delta_{1,n}|]$, we obtain (8) with $C:=2C_{p}/\gamma$, which proves the first claim.
For the second claim, take $\tau_{n}=n^{\alpha}$ with $0<\alpha<1/2$. Then
$\bigl|\mathcal{D}^{\mathrm{blanket}}-\mathcal{D}^{\mathrm{tblanket}}\bigr|\leq Cn^{-\alpha(p-1)}.$
Given any $K>0$, choose $p$ large enough so that $\alpha(p-1)>K$. Then $n^{K}\bigl|\mathcal{D}^{\mathrm{blanket}}-\mathcal{D}^{\mathrm{tblanket}}\bigr|\to 0$ as $n\to\infty$, i.e., the truncation error decays faster than the rate $n^{-K}$.
It remains to check that the truncated random variables $Z_{i,n}^{\tau_{n}}:=B_{i}l_{\varepsilon_{n},\tau_{n}}(Y_{i})$ inherit Assumption 2.8. Write $X_{n}:=l_{\varepsilon_{n}}(Y_{1})$ and note that
$X_{1,n}^{\tau_{n}}=X_{n}\mathbf{1}\{|X_{n}|\leq\tau_{n}\}+\mathbb{E}[X_{n}\mathbf{1}\{|X_{n}|>\tau_{n}\}],$
so
$|X_{1,n}^{\tau_{n}}|\leq|X_{n}|+\mathbb{E}|X_{n}|.$
Thus all moment bounds required in Assumption 2.8 for $X_{n}$ carry over to $X_{1,n}^{\tau_{n}}$ (up to different constants), and a fortiori to $Z_{1,n}^{\tau_{n}}=B_{1}X_{1,n}^{\tau_{n}}$ since $|B_{1}|\leq 1$.
For the continuity part of Assumption 2.8, the distribution of $X_{n}$ has a common absolutely continuous component on a fixed bounded interval $I=[a,b]$, with mass and density bounded away from zero and infinity, uniformly in $n$. Since $Z_{1,n}=B_{1}X_{1,n}$ with $B_{1}$ independent of $X_{1,n}$, the distribution of $Z_{1,n}$ is a mixture of an atom at 0 (coming from $B_{1}=0$) and the distribution of $X_{1,n}$ (coming from $B_{1}=1$), so $Z_{1,n}$ and $Z_{1,n}^{\tau_{n}}$ inherit the same absolutely continuous component on $I$ (up to a constant factor and, for $Z_{1,n}^{\tau_{n}}$, up to a translation).
Finally, by the above display and the uniform moment bounds, there exist constants $C<\infty$ and $n_{0}\in\mathbb{N}$ such that
$|X_{1,n}^{\tau_{n}}|\leq C\tau_{n}\quad\text{for all }n\geq n_{0},$
and therefore $|Z_{1,n}^{\tau_{n}}|=|B_{1}||X_{1,n}^{\tau_{n}}|\leq C\tau_{n}$ as well. The same bound holds for all $i\geq 1$ by identical distribution, which completes the proof. ∎
We now proceed to the second step using the truncated blanket divergence.
Lemma A.7 (Evaluation of the expectation).
Assume the setting of Assumption 2.8 and Definition 8. Let $(\varepsilon_{n})_{n\geq 1}$ be a sequence with $\varepsilon_{n}>0$, $\varepsilon_{n}\to 0$, $\varepsilon_{n}=\omega(n^{-1/2})$, and $\varepsilon_{n}=O\!\bigl(\sqrt{(\log n)/n}\bigr)$.
For each $n$ let $Y_{1},Y_{2},\dots$ be i.i.d. with distribution $\mathcal{R}_{\mathrm{ref}}$, let $B_{1},B_{2},\dots$ be i.i.d. $\mathrm{Bernoulli}(\gamma)$ independent of $(Y_{i})_{i\geq 1}$, and define the truncated variables
$Z_{i,n}^{\tau_{n}}:=B_{i}\,l_{\varepsilon_{n},\tau_{n}}(Y_{i}),\qquad i\geq 1.$
Write
$Z:=Z_{1,n}^{\tau_{n}},\qquad\mu_{n}:=\mathbb{E}[Z],\qquad\sigma_{\varepsilon_{n}}^{2}:=\mathrm{Var}(Z)\in(0,\infty).$
For $n\geq 2$ define
$S_{n-1}:=\sum_{i=2}^{n}Z_{i,n}^{\tau_{n}},\qquad T_{n-1}:=\frac{S_{n-1}-(n-1)\mu_{n}}{\sigma_{\varepsilon_{n}}\sqrt{n-1}},$
and let $F_{n-1}$ denote the distribution function of $T_{n-1}$. Set
$t_{n}:=-\frac{\mu_{n}\sqrt{n-1}}{\sigma_{\varepsilon_{n}}},\qquad\eta_{n}(z):=\frac{z}{\sigma_{\varepsilon_{n}}\sqrt{n-1}}.$
Then there exists a sequence $(r_{n})_{n\geq 1}$ with $r_{n}\in\mathbb{R}$ such that
(10) $\mathbb{E}\Bigl[Z\bigl(1-F_{n-1}\bigl(t_{n}-\eta_{n}(Z)\bigr)\bigr)\Bigr]=\frac{\varphi(t_{n})}{\sqrt{n-1}}\,\frac{\sigma_{\varepsilon_{n}}}{t_{n}^{2}}\,\bigl(1+r_{n}\bigr),$
and $r_{n}\to 0$ as $n\to\infty$.
Moreover, the sequence $(r_{n})$ depends only on the constants appearing in Assumption 2.8.
Proof.
Fix $n\geq 2$. Throughout the proof we suppress the dependence on $n$ in the notation whenever this causes no confusion, and write $Z_{i}:=Z_{i,n}^{\tau_{n}}$.
Step 1: Edgeworth expansion for the truncated array. For each $n$ the variables $Z_{1},Z_{2},\dots$ are i.i.d. with
$\mu_{n}:=\mathbb{E}[Z_{1}],\qquad\sigma_{\varepsilon_{n}}^{2}:=\mathrm{Var}(Z_{1}).$
By Assumption 2.8 and the truncation Lemma A.6, the triangular array $(Z_{i,n}^{\tau_{n}})_{1\leq i\leq n,\,n\geq 1}$ satisfies the hypotheses of Theorem A.1 (uniformly in $n$). In particular, for any fixed integer $s\geq 3$ there exist polynomials $P_{1,k},\dots,P_{s-2,k}$ and constants $C_{s}<\infty$ and $k_{0}\in\mathbb{N}$ such that, for all $k\geq k_{0}$, the distribution $F_{k}$ of
$T_{k}:=\frac{S_{k}-k\mu_{n}}{\sigma_{\varepsilon_{n}}\sqrt{k}},\qquad S_{k}:=\sum_{i=1}^{k}Z_{i},$
satisfies
$\sup_{u\in\mathbb{R}}\biggl|F_{k}(u)-\Phi(u)-\sum_{j=1}^{s-2}k^{-j/2}P_{j,k}(u)\varphi(u)\biggr|\leq C_{s}k^{-(s-2)/2}.$
We apply this with k=n1k=n-1 (for $n$ large enough so that $n-1\geq k_{0}$). Thus,
(11) $\sup_{u\in\mathbb{R}}\biggl|F_{n-1}(u)-\Phi(u)-\sum_{j=1}^{s-2}(n-1)^{-j/2}P_{j,n-1}(u)\varphi(u)\biggr|\leq C_{s}(n-1)^{-(s-2)/2}.$
The polynomials $P_{j,n-1}$ and the constant $C_{s}$ may be chosen independently of $n$.
In the sequel we use the shorthand
$t_{n}:=-\frac{\mu_{n}\sqrt{n-1}}{\sigma_{\varepsilon_{n}}},\qquad\eta_{n}(z):=\frac{z}{\sigma_{\varepsilon_{n}}\sqrt{n-1}}.$
Step 2: Decomposition of the inner expectation. We start from
$\mathbb{E}\Bigl[Z\bigl(1-F_{n-1}(t_{n}-\eta_{n}(Z))\bigr)\Bigr].$
Using the expansion (11) at the point $u=t_{n}-\eta_{n}(Z)$, we write
$1-F_{n-1}(u)=(1-\Phi(u))-\sum_{j=1}^{s-2}(n-1)^{-j/2}P_{j,n-1}(u)\varphi(u)+R_{n-1,s}(u),$
where the remainder $R_{n-1,s}$ satisfies $|R_{n-1,s}(u)|\leq C_{s}(n-1)^{-(s-2)/2}$ for all $u\in\mathbb{R}$. Substituting $u=t_{n}-\eta_{n}(Z)$ and taking the expectation with respect to $Z$ yields
$\mathbb{E}\Bigl[Z\bigl(1-F_{n-1}(t_{n}-\eta_{n}(Z))\bigr)\Bigr]$ $=A_{n}-\sum_{j=1}^{s-2}B_{n,j}+E_{n},$
where we set
$A_{n}$ $:=\mathbb{E}\Bigl[Z\bigl(1-\Phi(t_{n}-\eta_{n}(Z))\bigr)\Bigr],$
$B_{n,j}$ $:=(n-1)^{-j/2}\,\mathbb{E}\Bigl[ZP_{j,n-1}\bigl(t_{n}-\eta_{n}(Z)\bigr)\varphi\bigl(t_{n}-\eta_{n}(Z)\bigr)\Bigr],$
$E_{n}$ $:=\mathbb{E}\Bigl[ZR_{n-1,s}\bigl(t_{n}-\eta_{n}(Z)\bigr)\Bigr].$
Step 3: Bound on the Edgeworth remainder $E_{n}$. By Lemma A.6 there exist constants $C,h_{0}>0$ such that, for all $n$ large enough, $|Z|\leq C\tau_{n}$ almost surely. Using the bound on $R_{n-1,s}$ and this truncation, we obtain
$|E_{n}|\leq\mathbb{E}\Bigl[|Z|\,\bigl|R_{n-1,s}\bigl(t_{n}-\eta_{n}(Z)\bigr)\bigr|\Bigr]\leq C\tau_{n}C_{s}(n-1)^{-(s-2)/2}=O\bigl(\tau_{n}(n-1)^{-(s-2)/2}\bigr).$
Step 4: Evaluation of $A_{n}$. Define $g(z):=1-\Phi(z)$. Then
$g^{\prime}(z)=-\varphi(z),\qquad g^{\prime\prime}(z)=z\varphi(z).$
By Taylor’s theorem with Lagrange remainder, for each $z$ there exists $\xi=\xi(z)$ between $t_{n}$ and $t_{n}-\eta_{n}(z)$ such that
(12) $1-\Phi\bigl(t_{n}-\eta_{n}(z)\bigr)=g(t_{n})-g^{\prime}(t_{n})\eta_{n}(z)+\frac{1}{2}g^{\prime\prime}(\xi)\eta_{n}(z)^{2}.$
We now control $g^{\prime\prime}(\xi)$. By the explicit formula for the Gaussian density,
$\frac{\varphi(\xi)}{\varphi(t_{n})}=\exp\Bigl(-\frac{\xi^{2}-t_{n}^{2}}{2}\Bigr)=\exp\Bigl(t_{n}\theta\eta_{n}(z)-\tfrac{1}{2}\theta^{2}\eta_{n}(z)^{2}\Bigr),$
for some $\theta\in[0,1]$. Under our scaling assumptions, $\varepsilon_{n}=O(\sqrt{(\log n)/n})$ and $\varepsilon_{n}=\omega(n^{-1/2})$, and using that $\mu_{n}$ is of order $\varepsilon_{n}$ and $\sigma_{\varepsilon_{n}}$ is bounded away from 0 and $\infty$ (by Assumption 2.8 and the truncation lemma), we have
$t_{n}=-\frac{\mu_{n}\sqrt{n-1}}{\sigma_{\varepsilon_{n}}}=\Theta\bigl(\varepsilon_{n}\sqrt{n}\bigr),$
so $t_{n}\to\infty$ and $t_{n}^{2}=O(\log n)$.
Moreover, choosing $\tau_{n}=n^{\alpha}$ for some fixed $0<\alpha<1/2$, we have $|z|\leq C\tau_{n}$ and hence
$|\eta_{n}(z)|=O\Bigl(\frac{\tau_{n}}{\sqrt{n}}\Bigr)=O\bigl(n^{\alpha-1/2}\bigr)=o(1).$
Thus $|t_{n}\eta_{n}(z)|=o(1)$ and therefore $\varphi(\xi)\leq C_{\xi}\varphi(t_{n})$ for some constant $C_{\xi}$ independent of $n$. Furthermore, $|\xi|\leq|t_{n}|+|\eta_{n}(z)|\leq C(1+t_{n})$, so altogether
$|g^{\prime\prime}(\xi)|=|\xi|\varphi(\xi)\leq C^{\prime}(1+t_{n})\varphi(t_{n}).$
Substituting (12) into the definition of $A_{n}$ and using $\mathbb{E}[Z]=\mu_{n}$ gives
$A_{n}$ $=\mathbb{E}\Bigl[Z\bigl(1-\Phi(t_{n}-\eta_{n}(Z))\bigr)\Bigr]$
$=\mu_{n}(1-\Phi(t_{n}))+\varphi(t_{n})\mathbb{E}[Z\eta_{n}(Z)]+\frac{1}{2}\mathbb{E}\bigl[Zg^{\prime\prime}(\xi)\eta_{n}(Z)^{2}\bigr].$
Note that
$\mathbb{E}[Z\eta_{n}(Z)]=\frac{1}{\sigma_{\varepsilon_{n}}\sqrt{n-1}}\mathbb{E}[Z^{2}]=\frac{\sigma_{\varepsilon_{n}}^{2}+\mu_{n}^{2}}{\sigma_{\varepsilon_{n}}\sqrt{n-1}}.$
Using the bound on $g^{\prime\prime}(\xi)$ and the uniform boundedness of the third moment, we obtain
$\Bigl|\mathbb{E}\bigl[Zg^{\prime\prime}(\xi)\eta_{n}(Z)^{2}\bigr]\Bigr|\leq C^{\prime\prime}\frac{\varphi(t_{n})(1+t_{n})}{n-1}.$
Therefore
(13) $A_{n}=\mu_{n}(1-\Phi(t_{n}))+\frac{\sigma_{\varepsilon_{n}}^{2}+\mu_{n}^{2}}{\sigma_{\varepsilon_{n}}\sqrt{n-1}}\varphi(t_{n})+O\Bigl(\frac{\varphi(t_{n})(1+t_{n})}{n-1}\Bigr).$
Step 5: Evaluation of $B_{n,1}$. In one dimension, the first Edgeworth polynomial has the explicit form
$P_{1,n-1}(u)=\frac{\kappa_{3,n-1}}{6}(u^{2}-1),$
where $\kappa_{3,n-1}$ is the normalized third cumulant of $T_{n-1}$. Since the summands are i.i.d., we have
$\kappa_{3,n-1}=\frac{\kappa_{3}(Z_{1})}{\sigma_{\varepsilon_{n}}^{3}\sqrt{n-1}}=:\frac{\kappa_{3}}{\sigma_{\varepsilon_{n}}^{3}\sqrt{n-1}},$
where $\kappa_{3}$ is the third cumulant of $Z_{1}$, uniformly bounded in $n$ by the moment assumptions. Hence
$B_{n,1}=\frac{\kappa_{3}}{6\sigma_{\varepsilon_{n}}^{3}\sqrt{n-1}}\,\mathbb{E}\Bigl[Z\varphi(u)(u^{2}-1)\Bigr],\qquad u:=t_{n}-\eta_{n}(Z).$
Define $D(u):=\varphi(u)(u^{2}-1)$. A direct computation gives
$D^{\prime}(u)=-u\varphi(u)(u^{2}-3),\qquad D^{\prime\prime}(u)=\varphi(u)(u^{4}-6u^{2}+3).$
By Taylor’s theorem,
$D(u)=D(t_{n})+D^{\prime}(t_{n})\eta_{n}(Z)+\tfrac{1}{2}D^{\prime\prime}(\xi^{\prime})\eta_{n}(Z)^{2}$
for some $\xi^{\prime}=\xi^{\prime}(Z)$ between $t_{n}$ and $u$. Using the same argument as for $g^{\prime\prime}(\xi)$, we obtain $|D^{\prime\prime}(\xi^{\prime})|\leq C_{D}\varphi(t_{n})(1+t_{n}^{4})$ for all large $n$. Therefore
$D(u)$ $=\varphi(t_{n})(t_{n}^{2}-1)+\varphi(t_{n})t_{n}(t_{n}^{2}-3)\eta_{n}(Z)+O\bigl(\varphi(t_{n})t_{n}^{4}\eta_{n}(Z)^{2}\bigr),$
and hence
$B_{n,1}$ $=\frac{\kappa_{3}}{6\sigma_{\varepsilon_{n}}^{3}\sqrt{n-1}}\mathbb{E}\bigl[ZD(u)\bigr]$
$=\frac{\kappa_{3}\mu_{n}}{6\sigma_{\varepsilon_{n}}^{3}\sqrt{n-1}}\varphi(t_{n})(t_{n}^{2}-1)+\frac{\kappa_{3}}{6\sigma_{\varepsilon_{n}}^{3}\sqrt{n-1}}\varphi(t_{n})t_{n}(t_{n}^{2}-3)\mathbb{E}[Z\eta_{n}(Z)]$
(14) $\quad+O\Bigl(\frac{\varphi(t_{n})t_{n}^{4}}{\sigma_{\varepsilon_{n}}^{2}(n-1)}\mathbb{E}[Z^{3}]\Bigr).$
Using $\mathbb{E}[Z\eta_{n}(Z)]$ as above and bounded third moment, we obtain
(15) $B_{n,1}=\frac{\kappa_{3}\mu_{n}}{6\sigma_{\varepsilon_{n}}^{3}\sqrt{n-1}}\varphi(t_{n})(t_{n}^{2}-1)+\frac{\kappa_{3}(\sigma_{\varepsilon_{n}}^{2}+\mu_{n}^{2})}{6\sigma_{\varepsilon_{n}}^{4}}\varphi(t_{n})t_{n}(t_{n}^{2}-3)\frac{1}{n-1}+O\Bigl(\frac{\varphi(t_{n})(1+t_{n}^{4})}{(n-1)^{3/2}}\Bigr).$
Step 6: Bounds for $B_{n,j}$ with $j\geq 2$. By standard properties of Edgeworth polynomials (cf. Theorem A.1), for each $j\geq 2$ there exist constants $C_{j}<\infty$ and integers $k_{j}\leq 3j$ such that
$|P_{j,n-1}(u)|\leq C_{j}\bigl(1+|u|^{k_{j}}\bigr),\qquad u\in\mathbb{R}.$
As above, we have $t_{n}^{2}=O(\log n)$ and $|\eta_{n}(Z)|=O(\tau_{n}/\sqrt{n})=o(1)$, so $|u|=|t_{n}-\eta_{n}(Z)|\leq C(1+t_{n})$ and $\varphi(u)\leq C^{\prime}\varphi(t_{n})$ for all large $n$. Since $|Z|\leq C\tau_{n}$, we obtain
$|B_{n,j}|$ $\leq(n-1)^{-j/2}\,\mathbb{E}\bigl[|Z||P_{j,n-1}(u)|\varphi(u)\bigr]$
(16) $\leq C_{j}^{\prime\prime}(n-1)^{-j/2}\,\varphi(t_{n})\,(1+t_{n}^{k_{j}}),$
uniformly in $n$ for $2\leq j\leq s-2$.
Step 7: Extraction of the main term. Combining (13), (15), (16), and the bound on $E_{n}$, we obtain
$\mathbb{E}\Bigl[Z\bigl(1-F_{n-1}(t_{n}-\eta_{n}(Z))\bigr)\Bigr]$ $=A_{n}-B_{n,1}-\sum_{j=2}^{s-2}B_{n,j}+E_{n}$
$=\mu_{n}(1-\Phi(t_{n}))+\frac{\sigma_{\varepsilon_{n}}^{2}+\mu_{n}^{2}}{\sigma_{\varepsilon_{n}}\sqrt{n-1}}\varphi(t_{n})$
$\quad-\frac{\kappa_{3}\mu_{n}}{6\sigma_{\varepsilon_{n}}^{3}\sqrt{n-1}}\varphi(t_{n})(t_{n}^{2}-1)$
$\quad+O\Bigl(\frac{\varphi(t_{n})(1+t_{n}^{4})}{n-1}\Bigr)+O\Bigl(\sum_{j=2}^{s-2}(n-1)^{-j/2}\varphi(t_{n})t_{n}^{k_{j}}\Bigr)$
(17) $\quad+O\bigl(\tau_{n}(n-1)^{-(s-2)/2}\bigr).$
We now use the Mills ratio expansion
$1-\Phi(t_{n})=\frac{\varphi(t_{n})}{t_{n}}\Bigl(1-\frac{1}{t_{n}^{2}}+O\Bigl(\frac{1}{t_{n}^{4}}\Bigr)\Bigr),\qquad t_{n}\to\infty.$
Under our scaling assumptions,
$t_{n}=-\frac{\mu_{n}\sqrt{n-1}}{\sigma_{\varepsilon_{n}}}=\Theta\bigl(\varepsilon_{n}\sqrt{n}\bigr),$
so $t_{n}\to\infty$ and $t_{n}^{2}=O(\log n)$.
Substituting the Mills ratio into the first term of (17) and using the identity $t_{n}=-\mu_{n}\sqrt{n-1}/\sigma_{\varepsilon_{n}}$, a straightforward algebraic simplification shows that the terms of order $O(\varphi(t_{n})/\sqrt{n})$ cancel out, and the leading contribution from $A_{n}$ is
$\frac{\varphi(t_{n})}{\sqrt{n-1}}\,\frac{\sigma_{\varepsilon_{n}}}{t_{n}^{2}},$
while the remaining terms in (17) are of smaller order.
More precisely, we can write
(18) $\mathbb{E}\Bigl[Z\bigl(1-F_{n-1}(t_{n}-\eta_{n}(Z))\bigr)\Bigr]=\frac{\varphi(t_{n})}{\sqrt{n-1}}\,\frac{\sigma_{\varepsilon_{n}}}{t_{n}^{2}}\,\bigl(1+r_{n}\bigr),$
where the error term $r_{n}$ collects:
  • (a)
    the higher-order terms in the Mills ratio expansion (of order $O(t_{n}^{-2})$);
  • (b)
    the $O(\cdot)$-terms in (13) and (15);
  • (c)
    the sum over $j\geq 2$ in (16);
  • (d)
    the remainder $E_{n}$.
Using $t_{n}^{2}=O(\log n)$, $\tau_{n}=n^{\alpha}$ with $0<\alpha<1/2$, and choosing $s$ sufficiently large in the Edgeworth expansion, we deduce from the bounds in (17), (16), and the explicit form of the Mills ratio remainder that $r_{n}\to 0$ as $n\to\infty$. All constants involved depend only on the constants in Assumption 2.8, and the preceding estimates are uniform over the distinct pair $(x_{1},x_{1}^{\prime})$ and the admissible choices of $\mathcal{R}_{\mathrm{ref}}$.
Thus (18) proves (10), completing the proof. ∎
Final replacement of $\sigma_{\varepsilon_{n}}$ by $\sigma$.
We now relate the variance $\sigma_{\varepsilon_{n}}^{2}:=\mathrm{Var}(Z_{1,n}^{\tau_{n}})$ of the truncated variable appearing in Lemma A.7 to the limiting variance $\sigma^{2}=\mathrm{Var}(l_{0}(Y))$.
First, let $\sigma_{\mathrm{untr}}^{2}(\varepsilon):=\mathrm{Var}(B_{1}l_{\varepsilon}(Y_{1}))$ denote the variance of the original untruncated privacy amplification variable. From the explicit integral representation
$\mathbb{E}\bigl[l_{\varepsilon}(Y)^{2}\bigr]=\int\frac{\bigl(\mathcal{R}_{x_{1}}(y)-\mathrm{e}^{\varepsilon}\mathcal{R}_{x_{1}^{\prime}}(y)\bigr)^{2}}{\mathcal{R}_{\mathrm{ref}}(y)}\,dy,$
we observe that $\varepsilon\mapsto\sigma_{\mathrm{untr}}^{2}(\varepsilon)$ is continuously differentiable. Specifically,
$\sigma_{\mathrm{untr}}^{2}(\varepsilon)=\gamma\bigl(a_{1}+a_{2}\mathrm{e}^{2\varepsilon}-2a_{3}\mathrm{e}^{\varepsilon}\bigr)-\gamma^{2}(1-\mathrm{e}^{\varepsilon})^{2},$
where $a_{1},a_{2},a_{3}$ depend only on the reference distributions and they are uniformly bounded from Assumption 2.8. Differentiation shows that the derivative is uniformly bounded for $\varepsilon$ near 0. Since $\sigma_{\mathrm{untr}}^{2}(0)=\gamma\sigma^{2}$, the Lipschitz continuity implies
$\bigl|\sigma_{\mathrm{untr}}(\varepsilon_{n})-\sqrt{\gamma}\,\sigma\bigr|\leq C|\varepsilon_{n}|,$
for some constant $C<\infty$.
Second, we compare the truncated variance $\sigma_{\varepsilon_{n}}^{2}$ with the untruncated variance $\sigma_{\mathrm{untr}}^{2}(\varepsilon_{n})$. Recall that the truncated variable $Z_{1,n}^{\tau_{n}}$ is obtained by a mean-preserving truncation at level $\tau_{n}=n^{\alpha}$. By the Cauchy–Schwarz inequality and the moment bounds from Assumption 2.8, the difference is dominated by the tail contribution:
$\bigl|\sigma_{\varepsilon_{n}}^{2}-\sigma_{\mathrm{untr}}^{2}(\varepsilon_{n})\bigr|\leq\mathbb{E}\bigl[Z_{1,n}^{2}\mathbf{1}\{|Z_{1,n}|>\tau_{n}\}\bigr]\leq\frac{C_{p}}{\tau_{n}^{p-2}}.$
By choosing $p$ sufficiently large, this truncation error decays faster than any polynomial in $n$, and in particular is $o(\varepsilon_{n})$. Combining this with the Lipschitz bound yields
(19) $\sigma_{\varepsilon_{n}}=\sqrt{\gamma}\,\sigma+O(\varepsilon_{n}).$
Finally, we simplify the Gaussian factor $\varphi(t_{n})$. Define the parameter $\chi:=\sqrt{\gamma}/\sigma$. Using the definition $t_{n}=-\mu_{n}\sqrt{n-1}/\sigma_{\varepsilon_{n}}$ and the expansion (19), we write
$t_{n}=-\frac{\mu_{n}\sqrt{n-1}}{\sqrt{\gamma}\,\sigma(1+O(\varepsilon_{n}))}=-\frac{\mu_{n}\sqrt{n-1}}{\sqrt{\gamma}\,\sigma}+r_{n},$
where the error term $r_{n}$ satisfies $r_{n}=O(\mu_{n}\sqrt{n-1}\cdot\varepsilon_{n})=O(\sqrt{n}\varepsilon_{n}^{2})$. Since we assume $\varepsilon_{n}=O(\sqrt{(\log n)/n})$, we have
$t_{n}=-\chi(\mathrm{e}^{\varepsilon_{n}}-1)\sqrt{n-1}(1+r_{n})=-\chi(\mathrm{e}^{\varepsilon_{n}}-1)\sqrt{n-1}(1+o(1)).$
A Taylor expansion of the Gaussian density $\varphi(t)$ around $\bar{t}_{n}:=-\mu_{n}\sqrt{n-1}/(\sqrt{\gamma}\sigma)$ gives
$\varphi(t_{n})=\varphi(\bar{t}_{n})\bigl(1+o(1)\bigr)=\varphi\bigl(\chi(\mathrm{e}^{\varepsilon_{n}}-1)\sqrt{n-1}\bigr)\bigl(1+o(1)\bigr).$
By continuity of $x\mapsto x^{3}$ on $(0,\infty)$ we obtain
(20) $\sigma_{\varepsilon_{n}}^{3}=\sigma_{0}^{3}\bigl(1+o(1)\bigr)=\bigl(\sqrt{\gamma}\,\sigma\bigr)^{3}\bigl(1+o(1)\bigr)=\gamma^{3/2}\sigma^{3}\bigl(1+o(1)\bigr).$
From the Edgeworth-based analysis with $Z_{i,n}^{\tau_{n}}$ (i.e., (18)) we previously obtained an expansion of the form
$\mathcal{D}^{\mathrm{blanket}}=\frac{1}{\gamma}\frac{\varphi(t_{n})}{\sqrt{n-1}}\,\frac{\sigma_{\varepsilon_{n}}}{t_{n}^{2}}\,\bigl(1+o(1)\bigr),$
where the factor $1/\gamma$ comes from the transformation $\mathcal{D}^{\mathrm{blanket}}=\gamma^{-1}\mathbb{E}[Z_{1}\mathbf{1}\{T_{n-1}>-Z_{1}\}]$. Using (19) and (20), and recalling $\mu_{n}=\gamma(1-{\rm e}^{\varepsilon_{n}})$, we obtain
$\mathcal{D}^{\mathrm{blanket}}=\varphi\bigl(\chi(\mathrm{e}^{\varepsilon_{n}}-1)\sqrt{n-1}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{\bigl(\mathrm{e}^{\varepsilon_{n}}-1\bigr)^{2}(n-1)^{3/2}}\right)\bigl(1+o(1)\bigr).$
Finally, we may replace n1n-1 by $n$ everywhere in the display, at the cost of modifying the $o(1)$ term, which yields the desired asymptotic expression for the blanket divergence.
Next, we state the proof for the (Bound) case in the similar way as above.
Corollary A.8 (Discrete bounded case).
Under the assumptions of Lemma 3.2, the same conclusion holds if Assumption 2.8 is replaced by its bounded version (Bound).
Proof.
We use the integral representation (i.e., (ii) of Lemma 3.1) [^9] We deliberately work with the integral representation (ii) in Lemma 3.1 rather than the linear form (i). In the discrete bounded case, starting from (i) would lead to terms of the form $\mathbb{E}[Z_{1}(1-F_{n-1}(u))\rho_{n}(u)]$, and controlling the contribution of the Slastnikov remainder $\rho_{n}(u)$ relative to the main term would require an explicit convergence rate for the moderate deviation error, which is not available in the version we use. By contrast, the integral representation keeps the factor $(1+\rho_{n}(x))$ inside the same tail integral and allows us to obtain a relative $o(1)$ error without an explicit rate. Conversely, in the continuous case we avoid representation (ii), because our Edgeworth-based proof works under moment assumptions only and does not assume the existence of exponential moments, which are convenient for the Chernoff-type tail bounds needed in the discrete bounded analysis. together with Bernoulli selectors. For each $n$ let $Y_{1},\dots,Y_{n}$ be i.i.d. with distribution $\mathcal{R}_{\mathrm{ref}}$, let $B_{1},\dots,B_{n}$ be i.i.d. $\mathrm{Bernoulli}(\gamma)$ independent of $(Y_{i})_{i\geq 1}$. That is,
(21) $\mathcal{D}^{\mathrm{blanket}}_{e^{{\varepsilon_{n}}},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\frac{\sigma_{\varepsilon_{n}}}{\sqrt{n}\gamma}\,\int_{t_{n}}^{\infty}\bigl(1-F_{n}(x)\bigr)\,dx.$
By assumption (Bound) and the definition of $Z_{i}$ there exists $K<\infty$ such that $|Z_{i}|\leq K$ almost surely, uniformly in $n$ and in the choice of $(x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$.
Recall
$\mu_{n}:=\mathbb{E}[Z_{1}]=\gamma\,\mathbb{E}[X_{1}]=\gamma(1-{\rm e}^{\varepsilon_{n}}),\qquad\sigma_{\varepsilon_{n}}^{2}:=\mathrm{Var}(Z_{1}),$
and
$T_{n}:=\frac{S_{n}-n\mu_{n}}{\sigma_{\varepsilon_{n}}\sqrt{n}},\qquad F_{n}(x):=\Pr\left[T_{n}\leq x\right]$
Step 1: Moderate deviations for $T_{n}$. The array $(Z_{i})_{1\leq i\leq n,\,n\geq 1}$ is row-wise i.i.d., uniformly bounded, and satisfies $\mathbb{E}[Z_{1}]=\mu_{n},\;\mathrm{Var}(Z_{1})=\sigma_{\varepsilon_{n}}^{2}\geq C>0$ for all $n$, where $C$ depends only on the constants in Assumption 2.8. Hence Slastnikov’s moderate deviation theorem for triangular arrays (Theorem A.3) applies to the standardized sums $T_{n}$.
In particular, for every fixed $c>0$ there exists a sequence $\Delta_{n}\downarrow 0$ such that
(22) $\sup_{0\leq x\leq c\sqrt{\log n}}\left|\frac{\Pr\left[T_{n}>x\right]}{1-\Phi(x)}-1\right|\leq\Delta_{n},\qquad n\to\infty.$
Equivalently, there exists a function $\rho_{n}:[0,\infty)\to\mathbb{R}$ such that
$\Pr\left[T_{n}>x\right]=(1-\Phi(x))\bigl(1+\rho_{n}(x)\bigr),$
and
$\sup_{0\leq x\leq c\sqrt{\log n}}|\rho_{n}(x)|\leq\Delta_{n},\qquad\Delta_{n}\xrightarrow[n\to\infty]{}0.$
Under the scaling assumptions $\varepsilon_{n}=\omega(n^{-1/2}),\varepsilon_{n}=O(\sqrt{(\log n)/n})$, we have $\mu_{n}=\gamma(1-{\rm e}^{\varepsilon_{n}})=O(\varepsilon_{n})$ and $\sigma_{\varepsilon_{n}}$ bounded away from 0 and $\infty$, so
$t_{n}=-\frac{\mu_{n}\sqrt{n}}{\sigma_{\varepsilon_{n}}}=\Theta\bigl(\varepsilon_{n}\sqrt{n}\bigr)=O\bigl(\sqrt{\log n}\bigr),$
and in particular $t_{n}\to\infty$ and, for suitable $c>0$ and all large $n$,
$t_{n}\leq c\sqrt{\log n}.$
Step 2: Extraction of the main term. Substituting the representation of $\Pr\left[T_{n}>x\right]$ into (21) yields
$\mathcal{D}^{\mathrm{blanket}}=\frac{\sigma_{\varepsilon_{n}}}{\gamma\sqrt{n}}\int_{t_{n}}^{\infty}(1-\Phi(x))\bigl(1+\rho_{n}(x)\bigr)\,dx.$
Define
$\mathrm{Main}_{n}:=\int_{t_{n}}^{\infty}(1-\Phi(x))\,dx,\qquad\mathrm{Err}_{n}:=\int_{t_{n}}^{\infty}(1-\Phi(x))\rho_{n}(x)\,dx,$
so that
(23) $\mathcal{D}^{\mathrm{blanket}}=\frac{\sigma_{\varepsilon_{n}}}{\gamma\sqrt{n}}\,\bigl(\mathrm{Main}_{n}+\mathrm{Err}_{n}\bigr).$
Fix $B>1$ and split
$\mathrm{Err}_{n}=\int_{t_{n}}^{B\sqrt{\log n}}(1-\Phi(x))\rho_{n}(x)\,dx+\int_{B\sqrt{\log n}}^{\infty}(1-\Phi(x))\rho_{n}(x)\,dx=:I_{1}+I_{2}.$
For $I_{1}$, note that for all large $n$ we have $t_{n}\leq B\sqrt{\log n}\leq c\sqrt{\log n}$, so by (22),
$\sup_{x\in[t_{n},B\sqrt{\log n}]}|\rho_{n}(x)|\leq\Delta_{n}.$
Hence
$|I_{1}|\leq\Delta_{n}\int_{t_{n}}^{B\sqrt{\log n}}(1-\Phi(x))\,dx\leq\Delta_{n}\int_{t_{n}}^{\infty}(1-\Phi(x))\,dx=\Delta_{n}\,\mathrm{Main}_{n}.$
For $I_{2}$ we use crude exponential bounds on the tails. Since $Z_{i}$ are bounded, a Chernoff bound yields the existence of $c_{1}>0$ such that
$\Pr\left[T_{n}>x\right]\leq{\rm e}^{-c_{1}x^{2}},\qquad x\geq 0.$
Moreover, by Mills’ bound there exists $C_{1}>0$ such that
$1-\Phi(x)\leq C_{1}{\rm e}^{-x^{2}/2},\qquad x\geq 0.$
From the definition of $\rho_{n}$ we have $(1-\Phi(x))\rho_{n}(x)=\Pr\left[T_{n}>x\right]-(1-\Phi(x))$, and thus
$|(1-\Phi(x))\rho_{n}(x)|\leq\Pr\left[T_{n}>x\right]+(1-\Phi(x))\leq{\rm e}^{-c_{1}x^{2}}+C_{1}{\rm e}^{-x^{2}/2}.$
Therefore
$|I_{2}|$ $\leq\int_{B\sqrt{\log n}}^{\infty}\bigl({\rm e}^{-c_{1}x^{2}}+C_{1}{\rm e}^{-x^{2}/2}\bigr)\,dx$
$\leq C_{2}\,n^{-K(B)},$
for some finite constant $C_{2}<\infty$ and $K(B):=\min\{c_{1}B^{2},\tfrac{1}{2}B^{2}\}>0$.
On the other hand, by the standard asymptotic expansion of the Mills integral,
(24) $\mathrm{Main}_{n}=\int_{t_{n}}^{\infty}(1-\Phi(x))\,dx=\frac{\varphi(t_{n})}{t_{n}^{2}}\bigl(1+O(t_{n}^{-2})\bigr),\qquad t_{n}\to\infty.$
Since $t_{n}^{2}=\Theta(\log n)$, we have $\mathrm{Main}_{n}=\Theta(\varphi(t_{n})/t_{n}^{2})$, which decays only polynomially in $n$. Thus, by choosing $B>0$ large enough, we can ensure that $n^{-K(B)}=o(\mathrm{Main}_{n})$, and hence $|I_{2}|=o(\mathrm{Main}_{n})$.
Combining the bounds for $I_{1}$ and $I_{2}$, we obtain
$\mathrm{Err}_{n}=\mathrm{Main}_{n}\cdot o(1),$
and therefore, by (24),
$\mathrm{Main}_{n}+\mathrm{Err}_{n}=\frac{\varphi(t_{n})}{t_{n}^{2}}\bigl(1+o(1)\bigr).$
Substituting back into (23) gives
(25) $\mathcal{D}^{\mathrm{blanket}}=\frac{\sigma_{\varepsilon_{n}}}{\gamma\sqrt{n}}\,\frac{\varphi(t_{n})}{t_{n}^{2}}\,\bigl(1+o(1)\bigr),\qquad t_{n}:=-\frac{\mu_{n}\sqrt{n}}{\sigma_{\varepsilon_{n}}}.$
Using $t_{n}^{2}=\mu_{n}^{2}n/\sigma_{\varepsilon_{n}}^{2}$, we can rewrite this as
$\mathcal{D}^{\mathrm{blanket}}=\varphi(t_{n})\,\frac{\sigma_{\varepsilon_{n}}^{3}}{\gamma\,\mu_{n}^{2}n^{3/2}}\,\bigl(1+o(1)\bigr).$
Step 3: Final replacement of $\sigma_{\varepsilon_{n}}$ by $\sigma$. It remains to express the right-hand side in terms of the variance $\sigma^{2}=\mathrm{Var}(l_{0}(Y))$ and the parameter $\chi=\sqrt{\gamma}/\sigma$. As in the “final replacement” paragraph for the Bernoulli–weighted variables $Z_{\varepsilon}=B\,l_{\varepsilon}(Y)$, one checks that:
  • the map $\varepsilon\mapsto\sigma_{\varepsilon}^{2}:=\mathrm{Var}(Z_{\varepsilon})$ is $C^{1}$ and Lipschitz on $\varepsilon_{0},\varepsilon_{0}]$, so $|\sigma_{\varepsilon_{n}}-\sigma_{0}(Z)|\leq C|\varepsilon_{n}|$ with $\sigma_{0}(Z)^{2}=\gamma\sigma^{2}$;
  • consequently $\sigma_{\varepsilon_{n}}=\sigma_{0}(Z)\bigl(1+o(1)\bigr)=\sqrt{\gamma}\,\sigma\bigl(1+o(1)\bigr)$ and $\sigma_{\varepsilon_{n}}^{3}=\gamma^{3/2}\sigma^{3}\bigl(1+o(1)\bigr);$
  • with $t:=-\chi\bigl({\rm e}^{\varepsilon_{n}}-1\bigr)\sqrt{n}$ and $\mu_{n}=\gamma(1-{\rm e}^{\varepsilon_{n}})$, we have $t_{n}-t=o(1)$ and hence $\varphi(t_{n})=\varphi(t)\bigl(1+o(1)\bigr)$.
Substituting these into (25) yields
$\mathcal{D}^{\mathrm{blanket}}=\varphi\bigl(\chi({\rm e}^{\varepsilon_{n}}-1)\sqrt{n}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{\bigl({\rm e}^{\varepsilon_{n}}-1\bigr)^{2}n^{3/2}}\right)\bigl(1+o(1)\bigr),$
and this is exactly (3). The $o(1)$ term depends only on the constants in Assumption 2.8. ∎
A.3. The proof of Theorem 3.3
See 3.3
Proof of Theorem 3.3.
For each $n\geq 1$ and $\varepsilon\geq 0$, define the global privacy profile
$\delta_{n}(\varepsilon):=\sup_{x_{1:n}\simeq x^{\prime}_{1:n}}\mathcal{D}_{e^{\varepsilon}}\bigl(\mathcal{M}(x_{1:n})\,\big\|\,\mathcal{M}(x^{\prime}_{1:n})\bigr),\qquad\mathcal{M}=\mathcal{S}\circ\mathcal{R}^{n}.$
By definition of the hockey-stick divergence, the map $\alpha\mapsto\mathcal{D}_{\alpha}(P\|Q)$ is nonincreasing, hence $\varepsilon\mapsto\delta_{n}(\varepsilon)$ is nonincreasing for each fixed $n$.
Step 1: A uniform asymptotic form for the blanket divergence. For each pair $x_{1}\neq x_{1}^{\prime}\in\mathcal{X}$ and each reference distribution $\mathcal{R}_{\mathrm{ref}}\in\{\mathcal{R}_{\mathrm{BG}}\}\cup\{\mathcal{R}_{x}:x\in\mathcal{X}\}$, Lemma 3.2 gives an asymptotic expansion of the corresponding blanket divergence in the moderate-deviation regime. It is convenient to isolate the leading term as
$g_{n}(\varepsilon;\chi):=\varphi\bigl(\chi(\mathrm{e}^{\varepsilon}-1)\sqrt{n}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{(\mathrm{e}^{\varepsilon}-1)^{2}n^{3/2}}\right).$
By Assumption 2.8, the constants in the Edgeworth and moderate-deviation bounds can be chosen uniformly over all pairs $(x_{1},x_{1}^{\prime})$ and all admissible $\mathcal{R}_{\mathrm{ref}}$. Hence there exists a sequence $\rho_{n}\downarrow 0$ such that, for every choice of $(x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$ and every sequence $(\varepsilon_{n})$ with
$\varepsilon_{n}\to 0,\qquad\varepsilon_{n}=\omega(n^{-1/2}),\qquad\varepsilon_{n}=O\!\Bigl(\sqrt{\tfrac{\log n}{n}}\Bigr),$
we have
(26) $\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon_{n}},n,\mathcal{R}_{\mathrm{ref}},\gamma}\;=\;g_{n}\bigl(\varepsilon_{n};\chi(x_{1},x_{1}^{\prime};\mathcal{R}_{\mathrm{ref}})\bigr)\,\bigl(1+\theta_{n}(x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})\bigr),\qquad\bigl|\theta_{n}(x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})\bigr|\leq\rho_{n},$
where $\chi(x_{1},x_{1}^{\prime};\mathcal{R}_{\mathrm{ref}})$ is the shuffle index associated with $(x_{1},x_{1}^{\prime})$ and $\mathcal{R}_{\mathrm{ref}}$ as in Lemma 3.2.
A direct computation shows that for each fixed $n$ and $\varepsilon>0$ the map $\chi\mapsto g_{n}(\varepsilon;\chi)$ is strictly decreasing on $(0,\infty)$. Indeed, if we set $u:=\chi(\mathrm{e}^{\varepsilon}-1)\sqrt{n}$, then
$g_{n}(\varepsilon;\chi)=\frac{1}{\sqrt{2\pi}}\exp\!\Bigl(-\frac{u^{2}}{2}\Bigr)\cdot\frac{1}{\chi u^{2}\sqrt{n}},$
and the right-hand side is strictly decreasing in $\chi$.
Step 2: Solving the leading equation via the Lambert $W$-function. Fix $\alpha>0$ and $\chi>0$. We now solve the equation $g_{n}(\varepsilon;\chi)=\alpha/n$ explicitly. As above, write $u:=\chi(\mathrm{e}^{\varepsilon}-1)\sqrt{n}>0$. Then
$g_{n}(\varepsilon;\chi)=\frac{1}{\sqrt{2\pi}}\exp\!\Bigl(-\frac{u^{2}}{2}\Bigr)\cdot\frac{1}{\chi u^{2}\sqrt{n}}.$
The equation $g_{n}(\varepsilon;\chi)=\alpha/n$ is thus equivalent to
$\frac{1}{\sqrt{2\pi}}\exp\!\Bigl(-\frac{u^{2}}{2}\Bigr)\frac{1}{\chi u^{2}\sqrt{n}}=\frac{\alpha}{n},$
which can be rearranged as
$\exp\!\Bigl(\frac{u^{2}}{2}\Bigr)=\frac{\sqrt{n}}{\alpha\sqrt{2\pi}\,\chi u^{2}}.$
Multiplying both sides by $u^{2}/2$ yields
$\frac{u^{2}}{2}\,\exp\!\Bigl(\frac{u^{2}}{2}\Bigr)=\frac{\sqrt{n}}{2\alpha\chi\sqrt{2\pi}}.$
Setting $z:=u^{2}/2$ we obtain $z\,e^{z}=\frac{\sqrt{n}}{2\alpha\chi\sqrt{2\pi}}$, hence
$z=W\!\left(\frac{\sqrt{n}}{2\alpha\chi\sqrt{2\pi}}\right),\qquad u^{2}=2\,W\!\left(\frac{\sqrt{n}}{2\alpha\chi\sqrt{2\pi}}\right),$
where $W(\cdot)$ is the principal branch of the Lambert $W$-function. It follows that
$\mathrm{e}^{\varepsilon}-1=\frac{u}{\chi\sqrt{n}}=\sqrt{\frac{2}{\chi^{2}n}\,W\!\left(\frac{\sqrt{n}}{2\alpha\chi\sqrt{2\pi}}\right)},$
and hence
$\varepsilon_{n}(\alpha,\chi):=\log\left(1+\sqrt{\frac{2}{\chi^{2}n}\,W\!\left(\frac{\sqrt{n}}{2\alpha\chi\sqrt{2\pi}}\right)}\right)$
is the unique solution of $g_{n}(\varepsilon;\chi)=\alpha/n$ with $\varepsilon>0$.
From the standard asymptotics $W(z)=\log z-\log\log z+O(1)$ as $z\to\infty$ we see that
$W\!\left(\frac{\sqrt{n}}{2\alpha\chi\sqrt{2\pi}}\right)=\Theta(\log n),$
and therefore
$\mathrm{e}^{\varepsilon_{n}(\alpha,\chi)}-1=\Theta\left(\sqrt{\frac{\log n}{n}}\right),\qquad\varepsilon_{n}(\alpha,\chi)=\Theta\left(\sqrt{\frac{\log n}{n}}\right),$
which implies $\varepsilon_{n}(\alpha,\chi)\to 0$, $\varepsilon_{n}(\alpha,\chi)=\omega(n^{-1/2})$ and $\varepsilon_{n}(\alpha,\chi)=O(\sqrt{\log n/n})$. Thus $\varepsilon_{n}(\alpha,\chi)$ lies in the moderate-deviation regime where Lemma 3.2 is applicable. Moreover, from the explicit expression it is clear that $\chi\mapsto\varepsilon_{n}(\alpha,\chi)$ is nonincreasing.
Step 3: An upper bound via the blanket divergence. We first derive an asymptotic upper bound on $\delta_{n}(\varepsilon)$. By Lemma 2.5, for any neighboring datasets $x_{1:n}\simeq x^{\prime}_{1:n}$ differing in the first coordinate we have
$\mathcal{D}_{e^{\varepsilon}}\bigl(\mathcal{M}(x_{1:n})\,\big\|\,\mathcal{M}(x^{\prime}_{1:n})\bigr)\;\leq\;\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{BG}},\gamma}(\mathcal{R}_{x_{1}}\,\|\,\mathcal{R}_{x_{1}^{\prime}}).$
Taking the supremum over $x_{1:n}\simeq x^{\prime}_{1:n}$ yields
$\delta_{n}(\varepsilon)\leq\sup_{x_{1}\neq x_{1}^{\prime}}\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{BG}},\gamma}(\mathcal{R}_{x_{1}}\,\|\,\mathcal{R}_{x_{1}^{\prime}}).$
For each pair $(x_{1},x_{1}^{\prime})$ define
$\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime}):=\sqrt{\frac{\gamma}{\mathrm{Var}_{\mathcal{R}_{\mathrm{BG}}}\bigl(l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{BG}})\bigr)}},$
and let
$\underline{\chi}_{\mathrm{lo}}:=\inf_{x_{1}\neq x_{1}^{\prime}}\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime}).$
Assumption 2.8 implies that the variance in the denominator is finite and bounded away from zero uniformly over all pairs, so that $0<\underline{\chi}_{\mathrm{lo}}<\infty$.
Applying the uniform asymptotic (26) with $\mathcal{R}_{\mathrm{ref}}=\mathcal{R}_{\mathrm{BG}}$ and $\varepsilon_{n}=\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})$ (which lies in the moderate-deviation regime by Step 2) and using the monotonicity of $g_{n}$ in $\chi$ yields
$\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon_{n}},n,\mathcal{R}_{\mathrm{BG}},\gamma}(\mathcal{R}_{x_{1}}\,\|\,\mathcal{R}_{x_{1}^{\prime}})$ $=g_{n}\bigl(\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}});\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})\bigr)\,\bigl(1+\theta_{n}(x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{BG}})\bigr)$
$\leq(1+\rho_{n})\,g_{n}\bigl(\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}});\underline{\chi}_{\mathrm{lo}}\bigr),$
uniformly over all $x_{1}\neq x_{1}^{\prime}$. By construction of $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})$ in Step 2 we have
$g_{n}\bigl(\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}});\underline{\chi}_{\mathrm{lo}}\bigr)=\frac{\alpha}{n},$
hence
$\sup_{x_{1}\neq x_{1}^{\prime}}\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon_{n}},n,\mathcal{R}_{\mathrm{BG}},\gamma}(\mathcal{R}_{x_{1}}\,\|\,\mathcal{R}_{x_{1}^{\prime}})\leq\frac{\alpha}{n}\,(1+\rho_{n}).$
Combining this with the upper bound on $\delta_{n}$ gives
$\delta_{n}\bigl(\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})\bigr)\leq\frac{\alpha}{n}\,(1+\rho_{n})=\frac{\alpha}{n}\,(1+o(1)).$
Therefore the mechanism $\mathcal{M}$ satisfies
$\bigl(\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}}),\tfrac{\alpha}{n}(1+o(1))\bigr)\text{-DP}.$
Step 4: A lower bound via Su et al.’s blanket representation. We now show how the lower bound. By Theorem 2.6, for any $x,a_{1},a_{1}^{\prime}\in\mathcal{X}$ and any $\varepsilon\geq 0$,
$\delta_{n}(\varepsilon)\geq\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{x},1}(\mathcal{R}_{a_{1}}\,\|\,\mathcal{R}_{a_{1}^{\prime}}).$
For each pair $(x_{1},x_{1}^{\prime})$ define
$\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime}):=\inf_{x\in\mathcal{X}}\sqrt{\frac{1}{\mathrm{Var}_{\mathcal{R}_{x}}\bigl(l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{x})\bigr)}},$
and set
$\underline{\chi}_{\mathrm{up}}:=\inf_{x_{1}\neq x_{1}^{\prime}}\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime}).$
Again, Assumption 2.8 guarantees $0<\underline{\chi}_{\mathrm{up}}<\infty$.
Fix $\eta>0$. By definition of the infimum, there exist $x^{(\eta)}\in\mathcal{X}$ and a pair $(a_{1}^{(\eta)},a_{1}^{(\eta)}{}^{\prime})$ with $a_{1}^{(\eta)}\neq a_{1}^{(\eta)}{}^{\prime}$ such that
$\chi\bigl(a_{1}^{(\eta)},a_{1}^{(\eta)}{}^{\prime};\mathcal{R}_{x^{(\eta)}}\bigr)\leq\underline{\chi}_{\mathrm{up}}+\eta,$
where the left-hand side is the shuffle index associated with the reference distribution $\mathcal{R}_{x^{(\eta)}}$. Applying the asymptotic (26) with $\mathcal{R}_{\mathrm{ref}}=\mathcal{R}_{x^{(\eta)}}$ and $\varepsilon_{n}=\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{up}})$, and using that $g_{n}(\varepsilon;\chi)$ is decreasing in $\chi$, we obtain
$\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon_{n}},n,\mathcal{R}_{x^{(\eta)}},1}\bigl(\mathcal{R}_{a_{1}^{(\eta)}}\,\big\|\,\mathcal{R}_{a_{1}^{(\eta)}{}^{\prime}}\bigr)$ $=g_{n}\bigl(\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{up}});\chi(a_{1}^{(\eta)},a_{1}^{(\eta)}{}^{\prime};\mathcal{R}_{x^{(\eta)}})\bigr)\,\bigl(1+\theta_{n}(x^{(\eta)},a_{1}^{(\eta)},a_{1}^{(\eta)}{}^{\prime})\bigr)$
$\geq(1-\rho_{n})\,g_{n}\bigl(\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{up}});\underline{\chi}_{\mathrm{up}}+\eta\bigr),$
for all sufficiently large $n$.
By construction of $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{up}})$ we have
$g_{n}\bigl(\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{up}});\underline{\chi}_{\mathrm{up}}\bigr)=\frac{\alpha}{n}.$
The continuity of $g_{n}(\varepsilon;\chi)$ in $\chi$ and the fact that $g_{n}$ is smooth in $\chi$ on $(0,\infty)$ imply that, for each fixed $n$,
$g_{n}\bigl(\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{up}});\underline{\chi}_{\mathrm{up}}+\eta\bigr)=\frac{\alpha}{n}\bigl(1+\kappa_{n}(\eta)\bigr),$
where $\kappa_{n}(\eta)\to 0$ as $\eta\downarrow 0$, uniformly in $n$ on compact subsets of $(0,\infty)$ for $\underline{\chi}_{\mathrm{up}}$. Hence, for any fixed $\eta>0$,
$\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon_{n}},n,\mathcal{R}_{x^{(\eta)}},1}\bigl(\mathcal{R}_{a_{1}^{(\eta)}}\,\big\|\,\mathcal{R}_{a_{1}^{(\eta)}{}^{\prime}}\bigr)\geq\frac{\alpha}{n}\bigl(1-\rho_{n}\bigr)\bigl(1+\kappa_{n}(\eta)\bigr).$
Combining this with Theorem 2.6 gives
$\delta_{n}\bigl(\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{up}})\bigr)\;\geq\;\frac{\alpha}{n}\bigl(1-\rho_{n}\bigr)\bigl(1+\kappa_{n}(\eta)\bigr).$
Letting $n\to\infty$ and then $\eta\downarrow 0$ shows that there exists a sequence $\xi_{n}\to 0$ such that
$\delta_{n}\bigl(\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{up}})\bigr)\geq\frac{\alpha}{n}\,(1-\xi_{n})=\frac{\alpha}{n}\,(1+o(1)).$
Step 5: Existence of a true optimal privacy curve. Fix $\alpha>0$. We now show that, for all sufficiently large $n$, there exists an optimal privacy parameter $\varepsilon_{n}^{\ast}$ satisfying
$\delta_{n}\bigl(\varepsilon_{n}^{\ast}\bigr)=\frac{\alpha}{n}.$
The hockey-stick divergence is continuous, nonincreasing in $\varepsilon$, and converges to 0 as $\varepsilon\to\infty$ because $\mathcal{R}$ satisfies Assumption 2.8; the distributions are mutually absolutely continuous. Thus, for each fixed $n$, $\lim_{\varepsilon\to\infty}\delta_{n}(\varepsilon)=0.$
Fix $\alpha>0$. Since $\alpha/n\to 0$ as $n\to\infty$, for all sufficiently large $n$ we have
$\delta_{n}(0)>\frac{\alpha}{n}\qquad\text{and}\qquad\lim_{\varepsilon\to\infty}\delta_{n}(\varepsilon)=0<\frac{\alpha}{n}.$
By monotonicity, the set
$S_{n}(\alpha):=\Bigl\{\varepsilon\geq 0:\delta_{n}(\varepsilon)\leq\frac{\alpha}{n}\Bigr\}$
is nonempty and contained in $0,\infty)$. Define
$\varepsilon_{n}^{\ast}(\alpha):=\inf S_{n}(\alpha)\in[0,\infty).$
The right-continuity of $\delta_{n}$ implies $\delta_{n}(\varepsilon_{n}^{\ast}(\alpha))\leq\alpha/n$, while the definition of the infimum and the strict inequality $\delta_{n}(0)>\alpha/n$ imply $\delta_{n}(\varepsilon_{n}^{\ast}(\alpha))\geq\alpha/n$. Hence, for all sufficiently large $n$,
$\delta_{n}\bigl(\varepsilon_{n}^{\ast}(\alpha)\bigr)=\frac{\alpha}{n}.$
Thus we have constructed an optimal privacy curve $(\varepsilon_{n}^{\ast}(\alpha))_{n\geq 1}$: for each large $n$ it is the smallest privacy parameter for which the shuffled mechanism $\mathcal{M}$ satisfies $(\varepsilon,\alpha/n)$-DP. Comparing this curve with the explicit bounds obtained in Steps 3 and 4 shows that it lies asymptotically within the band
$\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{up}})\;\leq\;\varepsilon_{n}^{\ast}(\alpha)\;\leq\;\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})\quad\text{up to a vanishing relative error as }n\to\infty,$
which completes the argument.
A.4. The proof of Theorem 3.4
See 3.4
Proof.
In the purely discrete case, all integrals over $\mathcal{Y}$ should be read as sums over $y\in\mathcal{Y}$. Recall the notation from Section 3.3. For $x\in\mathcal{X}$ we write $\mathcal{R}_{x}(y)$ for the density (or pmf) of $\mathcal{R}_{x}$ with respect to the reference measure on $\mathcal{Y}$. We set
$\underline{\mathcal{R}}(y):=\inf_{z\in\mathcal{X}}\mathcal{R}_{z}(y),\qquad\gamma:=\int_{\mathcal{Y}}\underline{\mathcal{R}}(y)\,dy,\qquad\mathcal{R}_{\mathrm{BG}}(y):=\frac{1}{\gamma}\,\underline{\mathcal{R}}(y).$
Thus $\underline{\mathcal{R}}(y)=\gamma\,\mathcal{R}_{\mathrm{BG}}(y)$ for all $y\in\mathcal{Y}$. For $x_{1}\neq x_{1}^{\prime}\in\mathcal{X}$ we define the disagreement set
$A(x_{1},x_{1}^{\prime}):=\bigl\{y\in\mathcal{Y}:\mathcal{R}_{x_{1}}(y)\neq\mathcal{R}_{x_{1}^{\prime}}(y)\bigr\}.$
We first prove the following lemma.
Lemma A.9 (Pairwise ordering and equality condition).
Let $x_{1}\neq x_{1}^{\prime}\in\mathcal{X}$ be fixed and consider the pairwise shuffle indices
$\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime}):=\sqrt{\frac{\gamma}{\mathrm{Var}_{Y\sim\mathcal{R}_{\mathrm{BG}}}\bigl(l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{BG}})\bigr)}},\qquad\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime}):=\inf_{x\in\mathcal{X}}\sqrt{\frac{1}{\mathrm{Var}_{Y\sim\mathcal{R}_{x}}\bigl(l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{x})\bigr)}}.$
Then
$\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime})\;\geq\;\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})\qquad\text{for all }x_{1}\neq x_{1}^{\prime}.$
Moreover, equality holds if and only if there exists an input $x^{\ast}\in\mathcal{X}$ such that
$\mathcal{R}_{x^{\ast}}(y)=\underline{\mathcal{R}}(y)\qquad\text{for almost every }y\in A(x_{1},x_{1}^{\prime}).$
Proof.
Fix $x_{1}\neq x_{1}^{\prime}\in\mathcal{X}$ and abbreviate
$\Delta(y):=\mathcal{R}_{x_{1}}(y)-\mathcal{R}_{x_{1}^{\prime}}(y),\qquad y\in\mathcal{Y}.$
Note that $\int_{\mathcal{Y}}\Delta(y)\,dy=0$, since both $\mathcal{R}_{x_{1}}$ and $\mathcal{R}_{x_{1}^{\prime}}$ are probability distributions with respect to the same reference measure.
Step 1: Variance under $\mathcal{R}_{\mathrm{BG}}$. For $Y\sim\mathcal{R}_{\mathrm{BG}}$ we have
$l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{BG}})=\frac{\mathcal{R}_{x_{1}}(Y)-\mathcal{R}_{x_{1}^{\prime}}(Y)}{\mathcal{R}_{\mathrm{BG}}(Y)}=\frac{\Delta(Y)}{\mathcal{R}_{\mathrm{BG}}(Y)}.$
Since $\mathbb{E}_{\mathcal{R}_{\mathrm{BG}}}[l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{BG}})]=0$, its variance is
$\sigma_{\mathrm{BG}}^{2}(x_{1},x_{1}^{\prime})$ $:=\mathrm{Var}_{Y\sim\mathcal{R}_{\mathrm{BG}}}\bigl(l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{BG}})\bigr)$
$=\int_{\mathcal{Y}}\frac{\Delta(y)^{2}}{\mathcal{R}_{\mathrm{BG}}(y)}\,dy.$
On the complement of $A(x_{1},x_{1}^{\prime})$ we have $\Delta(y)=0$, so
$\sigma_{\mathrm{BG}}^{2}(x_{1},x_{1}^{\prime})=\int_{A(x_{1},x_{1}^{\prime})}\frac{\Delta(y)^{2}}{\mathcal{R}_{\mathrm{BG}}(y)}\,dy.$
Step 2: Variance under $\mathcal{R}_{x}$. Similarly, for any $x\in\mathcal{X}$ and $Y\sim\mathcal{R}_{x}$,
$l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{x})=\frac{\mathcal{R}_{x_{1}}(Y)-\mathcal{R}_{x_{1}^{\prime}}(Y)}{\mathcal{R}_{x}(Y)}=\frac{\Delta(Y)}{\mathcal{R}_{x}(Y)},$
and again the mean is zero. Hence
$\sigma_{x}^{2}(x_{1},x_{1}^{\prime})$ $:=\mathrm{Var}_{Y\sim\mathcal{R}_{x}}\bigl(l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{x})\bigr)$
$=\int_{\mathcal{Y}}\frac{\Delta(y)^{2}}{\mathcal{R}_{x}(y)}\,dy=\int_{A(x_{1},x_{1}^{\prime})}\frac{\Delta(y)^{2}}{\mathcal{R}_{x}(y)}\,dy.$
Step 3: Comparison of variances. By definition of the blanket, for all $x\in\mathcal{X}$ and $y\in\mathcal{Y}$
$\underline{\mathcal{R}}(y)\leq\mathcal{R}_{x}(y).$
Since $\mathcal{R}_{\mathrm{BG}}(y)=\underline{\mathcal{R}}(y)/\gamma$, this implies
$\mathcal{R}_{x}(y)\;\geq\;\underline{\mathcal{R}}(y)=\gamma\,\mathcal{R}_{\mathrm{BG}}(y)\qquad\text{for all }x\in\mathcal{X},\ y\in\mathcal{Y}.$
Consequently, for all $x\in\mathcal{X}$ and all $y$ with $\Delta(y)\neq 0$,
$\frac{\Delta(y)^{2}}{\mathcal{R}_{x}(y)}\;\leq\;\frac{\Delta(y)^{2}}{\gamma\,\mathcal{R}_{\mathrm{BG}}(y)}.$
Integrating over $A(x_{1},x_{1}^{\prime})$ yields
$\sigma_{x}^{2}(x_{1},x_{1}^{\prime})=\int_{A(x_{1},x_{1}^{\prime})}\frac{\Delta(y)^{2}}{\mathcal{R}_{x}(y)}\,dy\;\leq\;\frac{1}{\gamma}\int_{A(x_{1},x_{1}^{\prime})}\frac{\Delta(y)^{2}}{\mathcal{R}_{\mathrm{BG}}(y)}\,dy=\frac{1}{\gamma}\,\sigma_{\mathrm{BG}}^{2}(x_{1},x_{1}^{\prime}).$
Taking the supremum over $x\in\mathcal{X}$ gives
$\sup_{x\in\mathcal{X}}\sigma_{x}^{2}(x_{1},x_{1}^{\prime})\;\leq\;\frac{1}{\gamma}\,\sigma_{\mathrm{BG}}^{2}(x_{1},x_{1}^{\prime}).$
By definition,
$\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})=\sqrt{\frac{\gamma}{\sigma_{\mathrm{BG}}^{2}(x_{1},x_{1}^{\prime})}},\qquad\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime})=\inf_{x\in\mathcal{X}}\sqrt{\frac{1}{\sigma_{x}^{2}(x_{1},x_{1}^{\prime})}}=\frac{1}{\sqrt{\sup_{x\in\mathcal{X}}\sigma_{x}^{2}(x_{1},x_{1}^{\prime})}}.$
Using the inequality on the variances we obtain
$\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime})=\frac{1}{\sqrt{\sup_{x}\sigma_{x}^{2}(x_{1},x_{1}^{\prime})}}\;\geq\;\sqrt{\frac{\gamma}{\sigma_{\mathrm{BG}}^{2}(x_{1},x_{1}^{\prime})}}=\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime}).$
Step 4: Characterization of equality. We now analyze when equality $\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime})=\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})$ holds. By the definitions above this is equivalent to
$\sup_{x\in\mathcal{X}}\sigma_{x}^{2}(x_{1},x_{1}^{\prime})=\frac{1}{\gamma}\,\sigma_{\mathrm{BG}}^{2}(x_{1},x_{1}^{\prime}).$
From the pointwise inequality
$\frac{\Delta(y)^{2}}{\mathcal{R}_{x}(y)}\;\leq\;\frac{\Delta(y)^{2}}{\gamma\,\mathcal{R}_{\mathrm{BG}}(y)},\qquad y\in A(x_{1},x_{1}^{\prime}),$
we see that for any fixed $x\in\mathcal{X}$ we have
$\sigma_{x}^{2}(x_{1},x_{1}^{\prime})=\frac{1}{\gamma}\,\sigma_{\mathrm{BG}}^{2}(x_{1},x_{1}^{\prime})\quad\Longleftrightarrow\quad\mathcal{R}_{x}(y)=\gamma\,\mathcal{R}_{\mathrm{BG}}(y)=\underline{\mathcal{R}}(y)\quad\text{for almost every }y\in A(x_{1},x_{1}^{\prime}).$
Indeed, if $\mathcal{R}_{x}(y)>\gamma\,\mathcal{R}_{\mathrm{BG}}(y)$ on a subset of $A(x_{1},x_{1}^{\prime})$ of positive measure where $\Delta(y)\neq 0$, the inequality would be strict on that set and hence strict after integration.
Thus equality $\sup_{x}\sigma_{x}^{2}(x_{1},x_{1}^{\prime})=\sigma_{\mathrm{BG}}^{2}(x_{1},x_{1}^{\prime})/\gamma$ holds if and only if there exists $x^{\ast}\in\mathcal{X}$ such that $\sigma_{x^{\ast}}^{2}(x_{1},x_{1}^{\prime})=\sigma_{\mathrm{BG}}^{2}(x_{1},x_{1}^{\prime})/\gamma$, which in turn is equivalent to
$\mathcal{R}_{x^{\ast}}(y)=\underline{\mathcal{R}}(y)\quad\text{for almost every }y\in A(x_{1},x_{1}^{\prime}).$
Combining this with the previous step proves the lemma. ∎
Assume that the infima in the definitions of $\underline{\chi}_{\mathrm{lo}}$ and $\underline{\chi}_{\mathrm{up}}$ are attained by some distinct pairs $(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})$ and $(x_{1}^{\mathrm{up}},x_{1}^{\prime\,\mathrm{up}})$, respectively.
Step 1: $(i)\Rightarrow(ii)$. Suppose first that $\underline{\chi}_{\mathrm{up}}=\underline{\chi}_{\mathrm{lo}}$. By definition,
$\underline{\chi}_{\mathrm{lo}}=\chi_{\mathrm{lo}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}}),\qquad\underline{\chi}_{\mathrm{up}}=\chi_{\mathrm{up}}(x_{1}^{\mathrm{up}},x_{1}^{\prime\,\mathrm{up}}).$
Lemma A.9 implies that for every neighboring pair $(x_{1},x_{1}^{\prime})$,
$\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime})\;\geq\;\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime}).$
In particular,
$\chi_{\mathrm{up}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})\;\geq\;\chi_{\mathrm{lo}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})=\underline{\chi}_{\mathrm{lo}},$
and by the definition of $\underline{\chi}_{\mathrm{up}}$,
$\underline{\chi}_{\mathrm{up}}\leq\chi_{\mathrm{up}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}}).$
Combining these inequalities with $\underline{\chi}_{\mathrm{up}}=\underline{\chi}_{\mathrm{lo}}$ yields
$\chi_{\mathrm{up}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})=\chi_{\mathrm{lo}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})=\underline{\chi}_{\mathrm{lo}}.$
Applying Lemma A.9 to the pair $(x_{1}^{\ast},x_{1}^{\prime\,\ast}):=(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})$, we conclude that there exists $x^{\ast}\in\mathcal{X}$ such that
$\mathcal{R}_{x^{\ast}}(y)=\underline{\mathcal{R}}(y)\quad\text{for almost every }y\in A(x_{1}^{\ast},x_{1}^{\prime\,\ast}),$
and, by construction, $\chi_{\mathrm{lo}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\underline{\chi}_{\mathrm{lo}}$. This is precisely condition (ii).
Step 2: $(ii)\Rightarrow(i)$. Conversely, suppose there exist a distinct pair $(x_{1}^{\ast},x_{1}^{\prime\,\ast})$ and $x^{\ast}\in\mathcal{X}$ such that
$\chi_{\mathrm{lo}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\underline{\chi}_{\mathrm{lo}}\qquad\text{and}\qquad\mathcal{R}_{x^{\ast}}(y)=\underline{\mathcal{R}}(y)\ \text{ for almost every }y\in A(x_{1}^{\ast},x_{1}^{\prime\,\ast}).$
By Lemma A.9, this implies
$\chi_{\mathrm{up}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\chi_{\mathrm{lo}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\underline{\chi}_{\mathrm{lo}}.$
On the other hand, for any distinct pair $(x_{1},x_{1}^{\prime})$ we have
$\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime})\;\geq\;\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})\;\geq\;\underline{\chi}_{\mathrm{lo}}$
by Lemma A.9 and the definition of $\underline{\chi}_{\mathrm{lo}}$. Therefore
$\underline{\chi}_{\mathrm{up}}=\inf_{x_{1}\neq x_{1}^{\prime}}\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime})\;\geq\;\underline{\chi}_{\mathrm{lo}},$
while the existence of $(x_{1}^{\ast},x_{1}^{\prime\,\ast})$ with $\chi_{\mathrm{up}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\underline{\chi}_{\mathrm{lo}}$ shows that
$\underline{\chi}_{\mathrm{up}}\leq\chi_{\mathrm{up}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\underline{\chi}_{\mathrm{lo}}.$
Hence $\underline{\chi}_{\mathrm{up}}=\underline{\chi}_{\mathrm{lo}}$, which is condition (i). This completes the proof.
A.5. The proof of Lemma 4.1
See 4.1
Proof.
Recall the definition of the blanket divergence (generalized to an arbitrary reference distribution $\mathcal{R}_{\mathrm{ref}}$): for $M_{0}\sim\mathrm{Bin}(n,\gamma)$ and $Y_{1},Y_{2},\dots\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}$,
$\mathcal{D}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}^{\text{blanket}}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})\;=\;\frac{1}{n\gamma}\,\mathbb{E}\!\left[\Bigl(\sum_{i=1}^{M_{0}}l_{\varepsilon}(Y_{i})\Bigr)_{+}\right],$
where $a_{+}:=\max\{a,0\}$ and
$l_{\varepsilon}(y)=l_{\varepsilon}(y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}}):=\frac{\mathcal{R}_{x_{1}}(y)-\mathrm{e}^{\varepsilon}\mathcal{R}_{x_{1}^{\prime}}(y)}{\mathcal{R}_{\mathrm{ref}}(y)}.$
We will transform this expression into the claimed probability representation in two steps.
Step 1: Size-biasing the binomial and isolating one summand. Let $X_{i}:=l_{\varepsilon}(Y_{i})$ and
$S_{m}:=\sum_{i=1}^{m}X_{i},\qquad m\geq 1.$
Then, for each fixed $m\geq 1$,
$(S_{m})_{+}=S_{m}\mathbf{1}\{S_{m}>0\}=\sum_{i=1}^{m}X_{i}\,\mathbf{1}\{S_{m}>0\}.$
Taking expectation and using the symmetry of the $X_{i}$,
$\mathbb{E}\bigl[(S_{m})_{+}\bigr]=\sum_{i=1}^{m}\mathbb{E}\bigl[X_{i}\,\mathbf{1}\{S_{m}>0\}\bigr]=m\,\mathbb{E}\bigl[X_{1}\,\mathbf{1}\{S_{m}>0\}\bigr].$
Hence, conditioning on $M_{0}$,
$\mathcal{D}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}^{\text{blanket}}$ $=\frac{1}{n\gamma}\,\mathbb{E}\bigl[(S_{M_{0}})_{+}\bigr]$
$=\frac{1}{n\gamma}\,\mathbb{E}\bigl[\,\mathbb{E}\bigl[(S_{M_{0}})_{+}\bigm|M_{0}\bigr]\,\bigr]$
$=\frac{1}{n\gamma}\,\mathbb{E}\bigl[\,M_{0}\,\mathbb{E}\bigl[X_{1}\,\mathbf{1}\{S_{M_{0}}>0\}\bigm|M_{0}\bigr]\bigr].$
Now recall that the size-biased version of a nonnegative integer-valued random variable $M_{0}$ is defined by
$\mathbb{P}(M=m):=\frac{m\,\mathbb{P}(M_{0}=m)}{\mathbb{E}[M_{0}]}\qquad(m\geq 1).$
When $M_{0}\sim\mathrm{Bin}(n,\gamma)$, its size-biased version is
$M\;\stackrel{{\scriptstyle d}}{{=}}\;1+\mathrm{Bin}(n-1,\gamma),$
and $\mathbb{E}[M_{0}]=n\gamma$. Therefore, for any function $g$,
$\frac{1}{n\gamma}\,\mathbb{E}\bigl[M_{0}g(M_{0})\bigr]=\mathbb{E}\bigl[g(M)\bigr].$
Applying this with $g(m)=\mathbb{E}[X_{1}\,\mathbf{1}\{S_{m}>0\}\mid M_{0}=m]$, we obtain
$\mathcal{D}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}^{\text{blanket}}=\mathbb{E}\bigl[X_{1}\,\mathbf{1}\{S_{M}>0\}\bigr],$
where now $M\sim 1+\mathrm{Bin}(n-1,\gamma)$ and $S_{M}=\sum_{i=1}^{M}X_{i}=\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})$.
In other words,
(27) $\mathcal{D}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}^{\text{blanket}}=\mathbb{E}_{\begin{subarray}{c}M\sim 1+\mathrm{Bin}(n-1,\gamma)\\ Y_{1:n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}\end{subarray}}\Bigl[l_{\varepsilon}(Y_{1})\,\mathbf{1}\Bigl\{\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\Bigr\}\Bigr].$
Step 2: Change of measure on $Y_{1}$. By definition of $l_{\varepsilon}$,
$l_{\varepsilon}(y)\,\mathcal{R}_{\mathrm{ref}}(y)=\mathcal{R}_{x_{1}}(y)-\mathrm{e}^{\varepsilon}\mathcal{R}_{x_{1}^{\prime}}(y).$
Hence, for any bounded measurable function $h$ (possibly depending on other random variables),
$\mathbb{E}_{Y_{1}\sim\mathcal{R}_{\mathrm{ref}}}\bigl[l_{\varepsilon}(Y_{1})\,h(Y_{1})\bigr]$ $=\int h(y)\,l_{\varepsilon}(y)\,\mathcal{R}_{\mathrm{ref}}(y)\,dy$
$=\int h(y)\,\bigl(\mathcal{R}_{x_{1}}(y)-\mathrm{e}^{\varepsilon}\mathcal{R}_{x_{1}^{\prime}}(y)\bigr)\,dy$
$=\mathbb{E}_{Y_{1}\sim\mathcal{R}_{x_{1}}}[h(Y_{1})]-\mathrm{e}^{\varepsilon}\mathbb{E}_{Y_{1}\sim\mathcal{R}_{x_{1}^{\prime}}}[h(Y_{1})].$
In (27), take
$h(Y_{1}):=\mathbf{1}\Bigl\{\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\Bigr\},$
and treat $M$ and $Y_{2:n}$ as part of the outer expectation. Then
$\mathcal{D}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}^{\text{blanket}}$ $=\mathbb{E}_{M,Y_{2:n}}\Bigl[\mathbb{E}_{Y_{1}\sim\mathcal{R}_{\mathrm{ref}}}\bigl[l_{\varepsilon}(Y_{1})\,\mathbf{1}\{\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\}\bigm|M,Y_{2:n}\bigr]\Bigr]$
$=\mathbb{E}_{M,Y_{2:n}}\Bigl[\mathbb{P}_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}}\end{subarray}}\Bigl(\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\Bigm|M,Y_{2:n}\Bigr)$
$\qquad\qquad\qquad\quad-\mathrm{e}^{\varepsilon}\mathbb{P}_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}^{\prime}}\end{subarray}}\Bigl(\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\Bigm|M,Y_{2:n}\Bigr)\Bigr]$
$=\mathbb{P}_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}}\\ Y_{2:n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}\\ M\sim 1+\mathrm{Bin}(n-1,\gamma)\end{subarray}}\Bigl[\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\Bigr]$
$\quad-\mathrm{e}^{\varepsilon}\mathbb{P}_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}^{\prime}}\\ Y_{2:n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}\\ M\sim 1+\mathrm{Bin}(n-1,\gamma)\end{subarray}}\Bigl[\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\Bigr].$
This is exactly the desired expression. ∎
A.6. The proof of Theorem 5
See 4.2
Proof.
This proof is an adaptation of the FFT error analysis of Gopi et al. (Gopi et al., 2021) to our blanket-divergence setting.
Fix $n\in\mathbb{N}$, $\gamma\in(0,1]$, $\varepsilon\in\mathbb{R}$, and distinct inputs $x_{1},x_{1}^{\prime}\in\mathcal{X}$. Let $M\sim 1+\mathrm{Bin}(n-1,\gamma)$ and $Y_{2:n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}},$ all independent of $M$. Define
(28) $p(x):=\Pr_{Y_{1}\sim\mathcal{R}_{x}}\Bigl[l_{\varepsilon}(Y_{1};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})+\sum_{i=2}^{M}l_{\varepsilon}(Y_{i};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})>0\Bigr].$
We write $l(Y_{i})$ as a shorthand for $l_{\varepsilon}(Y_{i};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$ in the following. From the transformation lemma (i.e., Lemma 4.1), the blanket divergence can be written as
$D_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}^{\mathrm{blanket}}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})\;=\;p(x_{1})-\mathrm{e}^{\varepsilon}p(x_{1}^{\prime}).$
Our goal is to compare each $p(x)$ with the corresponding main-term approximation $P(c,F_{x})$ output by the algorithms, and then combine the two inequalities.
Step 1: Truncation error. Fix constants $s,w^{\mathrm{in}}$ and, for $i\geq 2$, let
$Z_{i}^{\mathrm{tr}}\sim l(Y_{i})\ \text{conditioned on}\ l(Y_{i})\in\bigl[s,s+w^{\mathrm{in}}\bigr].$
Then for each $x\in\{x_{1},x_{1}^{\prime}\}$ we have
$p(x)$ $=\Pr\Bigl[l(Y_{1})+\textstyle\sum_{i=2}^{M}l(Y_{i})>0\Bigr]$
(29) $\leq\Pr\Bigl[l(Y_{1})+\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}>0\Bigr]+\Pr\bigl[\perp\bigr],$
(30) $p(x)$ $\geq\Pr\Bigl[l(Y_{1})+\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}>0\Bigr]-\Pr\bigl[\perp\bigr].$
We abbreviate the truncation error by
$E_{\mathrm{trunc}}(x):=\Pr\bigl[\perp\bigr].$
Step 2: Discretization error. Fix a bin width $h>0$, and for $i\geq 2$ let $Z_{i}^{\mathrm{di}}$ be a discretized version of $Z_{i}^{\mathrm{tr}}$ on the grid $\mathcal{G}=\{x_{j}=jh:j\in\mathbb{Z}\}$ such that $|Z_{i}^{\mathrm{tr}}-Z_{i}^{\mathrm{di}}|\leq h/2$ almost surely. Let
$\Delta_{\mathrm{ES}}:=\mathbb{E}\Bigl[\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}-\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}\Bigr].$
Then using the union bound for any $c\geq 0$ we can write
$\Pr\Bigl[l(Y_{1})+\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}>0\Bigr]$
$=\Pr\Bigl[\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}>-l(Y_{1})\Bigr]$
$=\Pr\Bigl[\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}-\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}+\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}-\Delta_{\mathrm{ES}}>-l(Y_{1})-\Delta_{\mathrm{ES}}\Bigr]$
$\leq\Pr\Bigl[\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}-\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}-\Delta_{\mathrm{ES}}>c\Bigr]+\Pr\Bigl[\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}>-c-l(Y_{1})-\Delta_{\mathrm{ES}}\Bigr].$
We therefore define the discretization error
$E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}}):=\Pr\Bigl[\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}-\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}-\Delta_{\mathrm{ES}}>c\Bigr]$
and obtain the upper bound
(31) $\Pr\Bigl[l(Y_{1})+\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}>0\Bigr]\;\leq\;E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})+\Pr\Bigl[\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}>-c-l(Y_{1})-\Delta_{\mathrm{ES}}\Bigr].$
A symmetric argument, using a window shifted by c-c, yields a corresponding lower bound
(32) $\Pr\Bigl[l(Y_{1})+\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}>0\Bigr]\;\geq\;-E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})+\Pr\Bigl[\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}>+c-l(Y_{1})-\Delta_{\mathrm{ES}}\Bigr].$
Step 3: Aliasing error and FFT main term. Fix an outer window size $w^{\mathrm{out}}>0$ and define $L^{\mathrm{out}},L^{\mathrm{out}}]:=[-w^{\mathrm{out}}/2,w^{\mathrm{out}}/2]$. We extend the PMF of the discretized variable $Z_{i}^{\mathrm{di}}$ by zero outside $L^{\mathrm{out}},L^{\mathrm{out}}]$ and use the FFT on a grid of size $N$ such that the period $P=2L^{\mathrm{out}}$ satisfies $P/m=h$ for some integer $m$. This exactly corresponds to computing the distribution of the random sum
$S:=\sum_{i=2}^{M}\widetilde{Z}_{i}^{\mathrm{di}},$
where $\widetilde{Z}_{i}^{\mathrm{di}}$ is the zero-padded periodic extension of $Z_{i}^{\mathrm{di}}$. The difference between $S$ and the true discretized sum $\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}$ is the aliasing error; we define
$E_{\mathrm{alias}}(x;h,s,w^{\mathrm{in}},w^{\mathrm{out}}):=\Pr\Bigl[\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}\neq S\Bigr].$
Then using the union bound for the term appearing in (31) we have
$\Pr\Bigl[\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}>-c-l(Y_{1})-\Delta_{\mathrm{ES}}\Bigr]$
$\leq\Pr\Bigl[\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}\neq S\Bigr]+\Pr\bigl[S>-c-l(Y_{1})-\Delta_{\mathrm{ES}}\bigr]$
(33) $=E_{\mathrm{alias}}(x;h,s,w^{\mathrm{in}},w^{\mathrm{out}})+\sum_{z}\Pr[S=z]\,\Pr\bigl[l(Y_{1})>-c-z-\mathbb{E}[\textstyle\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}]\bigr],$
where in the last equality we have simply re-centered the threshold by the mean of $\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}$, which only affects $\Delta_{\mathrm{ES}}$.
Let $F_{x}$ be the CDF of $l(Y)$ under $Y\sim\mathcal{R}_{x}$. Then the inner probability in (33) is
$\Pr\bigl[l(Y_{1})>-c-z-\mathbb{E}[\textstyle\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}]\bigr]=1-F_{x}\bigl(-c-z-\mathbb{E}[\textstyle\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}]\bigr).$
By construction, the algorithm CalculatePMF computes the PMF $\mathbf{p}_{S-\mu_{S^{\mathrm{di}}}}$ of the centered sum $S-\mu_{S^{\mathrm{di}}}$, where
$\mu_{S^{\mathrm{di}}}:=(n-1)\gamma\,\mu_{Z^{\mathrm{di}}},$
and CalculateMainTerm computes
$\widehat{p}_{c}(x):=\textsc{CalculateMainTerm}(\mathbf{p}_{S-\mu_{S^{\mathrm{di}}}},F_{x},c)=\sum_{j}\mathbf{p}_{S-\mu_{S^{\mathrm{di}}}}[j]\,\Bigl(1-F_{x}\bigl(-c-z_{j}-\mathbb{E}[\textstyle\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}]\bigr)\Bigr),$
which coincides with the right-hand probability in (33). Thus, combining (31) and (33) with the truncation step (29), we obtain
(34) $p(x)$ $\leq\widehat{p}_{c}(x)+E_{\mathrm{alias}}(x;h,s,w^{\mathrm{in}},w^{\mathrm{out}})+E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})+E_{\mathrm{trunc}}(x).$
A completely symmetric argument, using c-c instead of +c+c in the union bounds, gives a second quantity $\widehat{p}_{-c}(x)=\textsc{CalculateMainTerm}(\mathbf{p}_{S-\mu_{S}},F_{x},-c)$ such that
(35) $p(x)$ $\geq\widehat{p}_{-c}(x)-E_{\mathrm{alias}}(x;h,s,w^{\mathrm{in}},w^{\mathrm{out}})-E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})-E_{\mathrm{trunc}}(x).$
Step 4: Incorporating the factor $(1-\Pr[\perp])$. In the theorem statement, the main term is defined as
$P(c,F_{x}):=(1-\Pr[\perp])\,\widehat{p}_{c}(x),\qquad P(-c,F_{x}):=(1-\Pr[\perp])\,\widehat{p}_{-c}(x),$
where $\Pr[\perp]=E_{\mathrm{trunc}}(x)$ is the truncation probability (note that $\perp$ depends only on $(Y_{i})_{i\geq 2}$, hence is the same for $x_{1}$ and $x_{1}^{\prime}$).
Since $0\leq\widehat{p}_{\pm c}(x)\leq 1$, we have
$\widehat{p}_{c}(x)-E_{\mathrm{trunc}}(x)\;\leq\;P(c,F_{x})\;\leq\;\widehat{p}_{c}(x),$
and similarly for c-c. Therefore, from (34) we obtain
(36) $p(x)$ $\leq P(c,F_{x})+E_{\mathrm{alias}}(x;h,s,w^{\mathrm{in}},w^{\mathrm{out}})+E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})+2E_{\mathrm{trunc}}(x),$
and from (35)
(37) $p(x)$ $\geq P(-c,F_{x})-E_{\mathrm{alias}}(x;h,s,w^{\mathrm{in}},w^{\mathrm{out}})-E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})-2E_{\mathrm{trunc}}(x).$
Step 5: Combining the two hypotheses. Apply (37) with $x=x_{1}$ and (36) with $x=x_{1}^{\prime}$:
$p(x_{1})$ $\geq P(-c,F_{x_{1}})-E_{\mathrm{alias}}(x_{1})-E_{\mathrm{disc}}(x_{1})-2E_{\mathrm{trunc}}(x_{1}),$
$p(x_{1}^{\prime})$ $\leq P(c,F_{x_{1}^{\prime}})+E_{\mathrm{alias}}(x_{1}^{\prime})+E_{\mathrm{disc}}(x_{1}^{\prime})+2E_{\mathrm{trunc}}(x_{1}^{\prime}),$
where we have suppressed the explicit dependence on $(c,h,s,w^{\mathrm{in}},w^{\mathrm{out}})$ in the notation of the error terms.
Hence
$D_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}^{\mathrm{blanket}}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})$ $=p(x_{1})-\mathrm{e}^{\varepsilon}p(x_{1}^{\prime})$
$\geq P(-c,F_{x_{1}})-\mathrm{e}^{\varepsilon}P(c,F_{x_{1}^{\prime}})-\delta_{\mathrm{err}}^{\mathrm{low}}(c,h,s,w^{\mathrm{in}},w^{\mathrm{out}}),$
where
$\delta_{\mathrm{err}}^{\mathrm{low}}:=E_{\mathrm{alias}}(x_{1})+E_{\mathrm{disc}}(x_{1})+2E_{\mathrm{trunc}}(x_{1})+\mathrm{e}^{\varepsilon}\bigl(E_{\mathrm{alias}}(x_{1}^{\prime})+E_{\mathrm{disc}}(x_{1}^{\prime})+2E_{\mathrm{trunc}}(x_{1}^{\prime})\bigr).$
Similarly, apply (36) with $x=x_{1}$ and (37) with $x=x_{1}^{\prime}$:
$p(x_{1})$ $\leq P(c,F_{x_{1}})+E_{\mathrm{alias}}(x_{1})+E_{\mathrm{disc}}(x_{1})+2E_{\mathrm{trunc}}(x_{1}),$
$p(x_{1}^{\prime})$ $\geq P(-c,F_{x_{1}^{\prime}})-E_{\mathrm{alias}}(x_{1}^{\prime})-E_{\mathrm{disc}}(x_{1}^{\prime})-2E_{\mathrm{trunc}}(x_{1}^{\prime}),$
so that
$D_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}^{\mathrm{blanket}}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})$ $=p(x_{1})-\mathrm{e}^{\varepsilon}p(x_{1}^{\prime})$
$\leq P(c,F_{x_{1}})-\mathrm{e}^{\varepsilon}P(-c,F_{x_{1}^{\prime}})+\delta_{\mathrm{err}}^{\mathrm{up}}(c,h,s,w^{\mathrm{in}},w^{\mathrm{out}}),$
where
$\delta_{\mathrm{err}}^{\mathrm{up}}:=E_{\mathrm{alias}}(x_{1})+E_{\mathrm{disc}}(x_{1})+2E_{\mathrm{trunc}}(x_{1})+\mathrm{e}^{\varepsilon}\bigl(E_{\mathrm{alias}}(x_{1}^{\prime})+E_{\mathrm{disc}}(x_{1}^{\prime})+2E_{\mathrm{trunc}}(x_{1}^{\prime})\bigr).$
By construction, both $\delta_{\mathrm{err}}^{\mathrm{low}}$ and $\delta_{\mathrm{err}}^{\mathrm{up}}$ are nonnegative and depend only on the numerical parameters $(c,h,s,w^{\mathrm{in}},w^{\mathrm{out}})$ and on the distributions $\mathcal{R}_{\mathrm{ref}},\mathcal{R}_{x_{1}},\mathcal{R}_{x_{1}^{\prime}}$ through the truncation, discretization, and aliasing probabilities.
See Section C for a discussion of how to rigorously control these error terms in practice.
This proves the desired two-sided bounds. ∎
A.7. The proof of Theorem 4.3
See 4.3
Proof.
Since $\mathcal{D}^{\mathrm{blanket}}\in[L_{n},U_{n}]$ from Theorem 5, we obtain
$\bigl|\widehat{D}_{n}-\mathcal{D}^{\mathrm{blanket}}\bigr|$ $\leq\frac{U_{n}-L_{n}}{2}$
$=\frac{1}{2}\Bigl(P(c;x_{1})-P(-c;x_{1})+\mathrm{e}^{\varepsilon}\bigl\{P(c;x_{1}^{\prime})-P(-c;x_{1}^{\prime})\bigr\}\Bigr)+\frac{\delta^{\mathrm{up}}_{\mathrm{err}}+\delta^{\mathrm{low}}_{\mathrm{err}}}{2}$
(38) $\leq\frac{1}{2}\bigl|P(c;x_{1})-P(-c;x_{1})\bigr|+\frac{\mathrm{e}^{\varepsilon}}{2}\bigl|P(c;x_{1}^{\prime})-P(-c;x_{1}^{\prime})\bigr|+\frac{\delta^{\mathrm{up}}_{\mathrm{err}}+\delta^{\mathrm{low}}_{\mathrm{err}}}{2}.$
In particular, the total approximation error decomposes into a main term, given by the differences $P(c;\cdot)-P(-c;\cdot)$, and the numerical error terms $\delta^{\mathrm{up}}_{\mathrm{err}}$ and $\delta^{\mathrm{low}}_{\mathrm{err}}$ (truncation, discretization, and aliasing). We bound each term in (38) relative to the true blanket divergence $\mathcal{D}^{\mathrm{blanket}}$.
We first consider the main term. Recall that $M\sim 1+\mathrm{Bin}(n-1,\gamma)$ and $Y_{2:n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}$, independent of $M$. Let
$S_{n-1}:=\sum_{i=2}^{M}l_{\varepsilon}\bigl(Y_{i};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}}\bigr),$
and let $\widetilde{S}_{n-1}$ denote the truncated, discretized, and periodized version of $S_{n-1}$ that is actually used in the FFT computation. Moreover, for $x\in\{x_{1},x_{1}^{\prime}\}$ we write
$Z_{1}:=l_{\varepsilon}\bigl(Y_{1};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}}\bigr),\qquad Y_{1}\sim\mathcal{R}_{x},$
and $\Delta_{\mathrm{ES}}$ is defined as in the proof of Theorem 5 (i.e., $\mu_{S^{\mathrm{tr}}}-\mu_{S^{\mathrm{di}}}$). With this notation, the main term appearing in (38) can be expressed as
(39) $P(c;x)-P(-c;x)$ $=(1-\Pr[\perp])\,\mathbb{E}_{Y_{1}\sim\mathcal{R}_{x}}\Bigl[\Pr\bigl[-c<\widetilde{S}_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}\leq c\,\big|\,Z_{1}\bigr]\Bigr],$
where $\perp$ is the truncation event defined in Theorem 5. In other words, the quantity P(c;x)$P(c;x)$-P(-c;x) is precisely the probability that the random sum $\widetilde{S}_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}$ falls inside the band (c,c)(-c,c), up to the multiplicative factor $(1-\Pr[\perp])$.
The approximating sum $\widetilde{S}_{n-1}$ is somewhat inconvenient to analyze directly, so we next replace it by the true sum $S_{n-1}$ at the price of an additional error term. Let
$\widetilde{S}_{n-1}=S_{n-1}+E_{n-1},\qquad E_{n-1}:=\widetilde{S}_{n-1}-S_{n-1},$
where $E_{n-1}$ collects the truncation, discretization, and aliasing errors. Fix $c>0$ and define the good event
$G:=\bigl\{|E_{n-1}|\leq c/2\bigr\},\qquad G^{c}:=\bigl\{|E_{n-1}|>c/2\bigr\}.$
Then, for any realization of $Z_{1}$ and on the event $G$, the inclusion
$\bigl\{-c<\widetilde{S}_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}\leq c\bigr\}\subset\bigl\{|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<c+|E_{n-1}|\bigr\}\subset\bigl\{|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c\bigr\}$
holds. Consequently,
$\Pr\bigl[-c<\widetilde{S}_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}\leq c\,\big|\,Z_{1}\bigr]$
(40) $\qquad\leq\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c\,\big|\,Z_{1}\bigr]+\Pr\bigl[|E_{n-1}|>c/2\bigr].$
Inserting this bound into (39) and taking the expectation over $Z_{1}$ yields
(41) $|P(c;x)-P(-c;x)|$ $\leq\mathbb{E}_{Y_{1}\sim\mathcal{R}_{x}}\Bigl[\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c\,\big|\,Z_{1}\bigr]\Bigr]+\Pr\bigl[|E_{n-1}|>c/2\bigr].$
Here, we used $1-\Pr[\perp]\leq 1$.
Next we further decompose the numerical error term $\Pr[|E_{n-1}|>c/2]$ into truncation, discretization, and wrap-around (aliasing) contributions.
Recall that
$S_{n-1}:=\sum_{i=2}^{M}l_{\varepsilon}\bigl(Y_{i};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}}\bigr)$
is the true sum, and that $\widetilde{S}_{n-1}$ denotes the truncated, discretized, and periodized version of $S_{n-1}$ used in the FFT computation. As before we write
$\widetilde{S}_{n-1}=S_{n-1}+E_{n-1},\qquad E_{n-1}:=\widetilde{S}_{n-1}-S_{n-1}.$
We view the construction of $\widetilde{S}_{n-1}$ as a sequence of three approximations:
  • Truncation:
    $S_{n-1}^{\mathrm{tr}}:=\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}},$
    where $Z_{i}^{\mathrm{tr}}$ is coupled with $l(Y_{i})$ such that $Z_{i}^{\mathrm{tr}}=l(Y_{i})$ whenever $l(Y_{i})\in[s,s+w^{\mathrm{in}}]$.
  • Discretization:
    $S_{n-1}^{\mathrm{di}}:=\sum_{i=2}^{M}Z_{i}^{\mathrm{di}},$
    where $Z_{i}^{\mathrm{di}}$ is the discretized version of $Z_{i}^{\mathrm{tr}}$ on the grid $\mathcal{G}=\{x_{j}=jh:j\in\mathbb{Z}\}$ with $|Z_{i}^{\mathrm{tr}}-Z_{i}^{\mathrm{di}}|\leq h/2$ almost surely.
  • Wrap-around (aliasing): $\widetilde{S}_{n-1}$ is obtained from $S_{n-1}^{\mathrm{di}}$ by periodic extension outside the outer window $w^{\mathrm{out}}/2,w^{\mathrm{out}}/2]$.
Accordingly, we decompose the total error into three components:
$E_{\mathrm{Mtrunc}}$ $:=S_{n-1}^{\mathrm{tr}}-S_{n-1},$
$E_{\mathrm{Mdisc}}$ $:=S_{n-1}^{\mathrm{di}}-S_{n-1}^{\mathrm{tr}},$
$E_{\mathrm{Malias}}$ $:=\widetilde{S}_{n-1}-S_{n-1}^{\mathrm{di}},$
so that
$E_{n-1}=E_{\mathrm{Mtrunc}}+E_{\mathrm{Mdisc}}+E_{\mathrm{Malias}}.$
By the triangle inequality we have
$|E_{n-1}|\leq|E_{\mathrm{Mtrunc}}|+|E_{\mathrm{Mdisc}}|+|E_{\mathrm{Malias}}|.$
Hence, for any $c>0$,
$\Bigl\{|E_{n-1}|>\tfrac{c}{2}\Bigr\}\subset\Bigl\{|E_{\mathrm{Mtrunc}}|>\tfrac{c}{6}\Bigr\}\,\cup\,\Bigl\{|E_{\mathrm{Mdisc}}|>\tfrac{c}{6}\Bigr\}\,\cup\,\Bigl\{|E_{\mathrm{Malias}}|>\tfrac{c}{6}\Bigr\},$
and by the union bound this yields
(42) $\Pr\bigl[|E_{n-1}|>c/2\bigr]\leq\Pr\bigl[|E_{\mathrm{Mtrunc}}|>c/6\bigr]+\Pr\bigl[|E_{\mathrm{Mdisc}}|>c/6\bigr]+\Pr\bigl[|E_{\mathrm{Malias}}|>c/6\bigr].$
This decomposes the tail probability of the total numerical error into three contributions corresponding to truncation, discretization, and wrap-around, respectively. In particular, since the error terms $\delta_{\mathrm{err}}^{\mathrm{low}}$ and $\delta_{\mathrm{err}}^{\mathrm{up}}$ in Theorem 5 are also built from these three errors, we evaluate each of them separately in the sequel.
Recall that the error terms in Theorem 5 are given by
$\delta_{\mathrm{err}}^{\mathrm{low}}=\delta_{\mathrm{err}}^{\mathrm{up}}:=E_{\mathrm{alias}}(x_{1})+E_{\mathrm{disc}}(x_{1})+2E_{\mathrm{trunc}}(x_{1})+\mathrm{e}^{\varepsilon}\bigl(E_{\mathrm{alias}}(x_{1}^{\prime})+E_{\mathrm{disc}}(x_{1}^{\prime})+2E_{\mathrm{trunc}}(x_{1}^{\prime})\bigr),$
where, for brevity, we suppress the dependence of $E_{\mathrm{alias}},E_{\mathrm{disc}},E_{\mathrm{trunc}}$ on $(c,h,s,w^{\mathrm{in}},w^{\mathrm{out}})$. Thus $\delta_{\mathrm{err}}^{\mathrm{low}}=\delta_{\mathrm{err}}^{\mathrm{up}}$, and
$\delta_{\mathrm{trunc}}$ $:=2E_{\mathrm{trunc}}(x_{1})+2\mathrm{e}^{\varepsilon}E_{\mathrm{trunc}}(x_{1}^{\prime}),$
$\delta_{\mathrm{disc}}$ $:=E_{\mathrm{disc}}(x_{1})+\mathrm{e}^{\varepsilon}E_{\mathrm{disc}}(x_{1}^{\prime}),$
$\delta_{\mathrm{alias}}$ $:=E_{\mathrm{alias}}(x_{1})+\mathrm{e}^{\varepsilon}E_{\mathrm{alias}}(x_{1}^{\prime}).$
Combining this with (41) and (38), and using that the bound on $\Pr[|E_{n-1}|>c/2]$ does not depend on $x$, we arrive at
$\bigl|\widehat{D}_{n}-\mathcal{D}^{\mathrm{blanket}}\bigr|$ $\leq\frac{1}{2}\,\mathbb{E}_{Y_{1}\sim\mathcal{R}_{x_{1}}}\Bigl[\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c\,\big|\,Z_{1}\bigr]\Bigr]$
$\quad+\frac{\mathrm{e}^{\varepsilon}}{2}\,\mathbb{E}_{Y_{1}\sim\mathcal{R}_{x_{1}^{\prime}}}\Bigl[\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c\,\big|\,Z_{1}\bigr]\Bigr]$
$\quad+\frac{1+\mathrm{e}^{\varepsilon}}{2}\,\Pr\bigl[|E_{\mathrm{Mtrunc}}|>c/6\bigr]+\delta_{\mathrm{trunc}}$
$\quad+\frac{1+\mathrm{e}^{\varepsilon}}{2}\,\Pr\bigl[|E_{\mathrm{Mdisc}}|>c/6\bigr]+\delta_{\mathrm{disc}}$
(43) $\quad+\frac{1+\mathrm{e}^{\varepsilon}}{2}\,\Pr\bigl[|E_{\mathrm{Malias}}|>c/6\bigr]+\delta_{\mathrm{alias}}.$
In other words, the total approximation error splits into a main band term (the first two lines of (43)) plus three groups of numerical error:
  • truncation error: $\delta_{\mathrm{trunc}}$ and $\Pr[|E_{\mathrm{Mtrunc}}|>c/6]$,
  • discretization error: $\delta_{\mathrm{disc}}$ and $\Pr[|E_{\mathrm{Mdisc}}|>c/6]$,
  • aliasing (wrap-around) error: $\delta_{\mathrm{alias}}$ and $\Pr[|E_{\mathrm{Malias}}|>c/6]$.
Thus each of the three numerical effects (truncation, discretization, and wrap-around) contributes via a $\delta$-part (coming from Theorem 5) and an $E_{\mathrm{M}}$-part coming from the replacement of $\widetilde{S}_{n-1}$ by the true sum $S_{n-1}$ in the main term.
Step 1: Main-term error and choice of $c_{n}$. We first evaluate the main term using the following lemma.
Lemma A.10 (Band probability).
Assume the setting of Lemma 3.2, and suppose that the local randomizer $\mathcal{R}$ satisfies Assumption 2.8.
Let $(c_{n})_{n\geq 1}$ be a sequence with $c_{n}>0$ such that
$\frac{c_{n}}{\sigma_{\varepsilon_{n}}\sqrt{n-1}}\longrightarrow 0,$
and suppose that $\Delta_{\mathrm{ES}}=O\left(\frac{\sigma_{\varepsilon_{n}}\sqrt{n}}{\sqrt{\log n}}\right).$ Write $t_{n}:=-\frac{\mu_{n}\sqrt{n-1}}{\sigma_{\varepsilon_{n}}}.$ Then there exists a constant $C<\infty$, depending only on the constants in Assumption 2.8, such that for all sufficiently large $n$ and for every $x\in\{x_{1},x_{1}^{\prime}\}$,
(44) $\mathbb{E}_{Z_{1}\sim\mathcal{R}_{x}}\Bigl[\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c_{n}\,\big|\,Z_{1}\bigr]\Bigr]\;\leq\;C\,\frac{c_{n}}{\sigma_{\varepsilon_{n}}\sqrt{n-1}}\,\varphi(t_{n}).$
We prove this lemma in Section A.7.1. We now evaluate the main-term error. Recall from Lemma A.10 that for any sequence $c_{n}>0$ satisfying $c_{n}/(\sigma_{\varepsilon_{n}}\sqrt{n-1})\to 0$ we have for each $x\in\{x_{1},x_{1}^{\prime}\}$,
(45) $\mathbb{E}_{Z_{1}\sim\mathcal{R}_{x}}\Bigl[\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c_{n}\,\big|\,Z_{1}\bigr]\Bigr]\;\leq\;C\,\frac{c_{n}}{\sigma_{\varepsilon_{n}}\sqrt{n-1}}\,\varphi(t_{n}),$
for some constant $C<\infty$ independent of $n$. Inserting this bound into the midpoint error decomposition (43) yields
[…]
On the other hand, by the asymptotic expansion of the blanket divergence, we have
(47) $\mathcal{D}^{\mathrm{blanket}}\;=\;\varphi\bigl(t_{n}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{\varepsilon_{n}^{2}n^{3/2}}\right)\bigl(1+o(1)\bigr),$
where $\chi>0$ is a constant.
Dividing (46) by (47), and using that $\sigma_{\varepsilon_{n}}\to\sigma>0$ as $n\to\infty$, we obtain
$\frac{\bigl|\text{(main term)}\bigr|}{\mathcal{D}^{\mathrm{blanket}}}$ $=O\!\left(\frac{\dfrac{c_{n}}{\sigma_{\varepsilon_{n}}\sqrt{n}}\varphi(t_{n})}{\varphi(t_{n})\,\dfrac{1}{\chi^{3}}\dfrac{1}{\varepsilon_{n}^{2}n^{3/2}}}\right)(1+o(1))$
$=O\!\left(c_{n}\,\varepsilon_{n}^{2}n\,\frac{\chi^{3}}{\sigma_{\varepsilon_{n}}}\right)(1+o(1))$
(48) $=O\!\left(c_{n}\,\varepsilon_{n}^{2}n\right),$
where the last step uses that $\chi^{3}/\sigma_{\varepsilon_{n}}$ is bounded uniformly in $n$.
We would like the contribution of the main term to the overall relative error to be at most $\eta_{\mathrm{main}}$, that is,
$\frac{\bigl|\text{(main term)}\bigr|}{\mathcal{D}^{\mathrm{blanket}}}\;\lesssim\;\eta_{\mathrm{main}}.$
From (48), a sufficient condition is
(49) $c_{n}\,\varepsilon_{n}^{2}n\;\lesssim\;\eta_{\mathrm{main}}.$
Equivalently, we may choose $c_{n}$ of order
(50) $c_{n}\;=\;\Theta\!\left(\frac{\eta_{\mathrm{main}}}{\varepsilon_{n}^{2}n}\right).$
With such a choice, the main-term contribution to the relative error is $O(\eta_{\mathrm{main}})$.
Finally, we check that the technical assumption of Lemma A.10, namely $c_{n}/(\sigma_{\varepsilon_{n}}\sqrt{n})\to 0$, is satisfied. Using (50), we have
$\frac{c_{n}}{\sigma_{\varepsilon_{n}}\sqrt{n}}=\Theta\!\left(\frac{\eta_{\mathrm{main}}}{\varepsilon_{n}^{2}n^{3/2}}\right).$
Under our standing assumption $\varepsilon_{n}^{2}n\to\infty$ (e.g. $\varepsilon_{n}^{2}=\Theta(\log n/n)$), the right-hand side tends to 0, so the condition of Lemma A.10 indeed holds.
In particular, in the canonical regime $\varepsilon_{n}^{2}=\Theta(\log n/n)$, we obtain the more explicit scaling
$c_{n}\;=\;\Theta\!\left(\frac{\eta_{\mathrm{main}}}{\log n}\right).$
Step 2: Truncation error and choice of $(s_{n},w_{n}^{\mathrm{in}})$. Recall the truncation event
$\perp:=\Bigl\{\exists i\in\{2,\dots,n\}:B_{i}=1,\ l_{\varepsilon_{n}}(Y_{i};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})\notin[s_{n},s_{n}+w_{n}^{\mathrm{in}}]\Bigr\},$
and write
$E_{\mathrm{trunc}}(n;s_{n},w_{n}^{\mathrm{in}}):=\Pr[\perp].$
Moreover, let
$S_{n-1}:=\sum_{i=2}^{M}l_{\varepsilon_{n}}(Y_{i};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}}),\qquad S_{n-1}^{\mathrm{tr}}:=\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}},$
where $Z_{i}^{\mathrm{tr}}$ are i.i.d. copies of the truncated random variable used in the FFT construction, and define
$E_{\mathrm{Mtrunc}}:=S_{n-1}^{\mathrm{tr}}-S_{n-1}.$
We control $E_{\mathrm{trunc}}$ and $\Pr[|E_{\mathrm{Mtrunc}}|>c_{n}/6]$ in terms of the blanket divergence.
Control of $E_{\mathrm{trunc}}$.
Denote by
$\tau_{n}:=\max\{|s_{n}|,\ |s_{n}+w_{n}^{\mathrm{in}}|\}$
the truncation radius. Assumption 2.8 guarantees that $l_{\varepsilon_{n}}(Y)$ has finite moments of all integer orders. In particular, for every $p\geq 2$ there exists a constant $C_{p}<\infty$ such that
$\mathbb{E}\bigl[|l_{\varepsilon_{n}}(Y)|^{p}\bigr]\leq C_{p}\qquad\text{for all }n.$
By Markov’s inequality, this implies a polynomial tail bound
(51) $\Pr\bigl[|l_{\varepsilon_{n}}(Y)|>T\bigr]\;\leq\;C_{p}\,T^{-p},\qquad T>0.$
Using the binomial structure of M1M-1 and a union bound, we obtain
$E_{\mathrm{trunc}}(n;s_{n},w_{n}^{\mathrm{in}})$ $=\Pr\bigl[\perp\bigr]$
$\leq\Pr\Bigl[\bigcup_{i=2}^{n}\bigl\{B_{i}=1,\ l_{\varepsilon_{n}}(Y_{i})\notin[s_{n},s_{n}+w_{n}^{\mathrm{in}}]\bigr\}\Bigr]$
$\leq K_{\gamma}\,n\,\Pr\bigl[|l_{\varepsilon_{n}}(Y)|>\tau_{n}\bigr]$
(52) $\leq K_{\gamma}C_{p}\,n\,\tau_{n}^{-p},$
for some constant $K_{\gamma}<\infty$ depending only on $\gamma$.
Dividing (52) by the blanket divergence, we get
$\frac{E_{\mathrm{trunc}}(n;s_{n},w_{n}^{\mathrm{in}})}{\mathcal{D}^{\mathrm{blanket}}}$ $\leq K^{\prime}_{\gamma,p}\,n\,\tau_{n}^{-p}\,\frac{\varepsilon_{n}^{2}n^{3/2}}{\varphi(t_{n})}(1+o(1))$
(53) $=K^{\prime}_{\gamma,p}\,\varepsilon_{n}^{2}n^{5/2}\,\frac{\tau_{n}^{-p}}{\varphi(t_{n})}(1+o(1)).$
In the regime of interest (in particular when $\varepsilon_{n}^{2}n$ and $t_{n}$ grow at most logarithmically in $n$), the prefactor $\varepsilon_{n}^{2}n^{3/2}/\varphi(t_{n})$ grows at most polynomially in $n$. Hence we can choose $p$ large enough and a polynomially growing sequence $\tau_{n}$ such that the right-hand side of (53) is arbitrarily small. For concreteness, fix $p\geq 4$ and set
(54) $\tau_{n}:=n^{\beta}(\log n)^{1/p}$
with some $\beta>0$ sufficiently large (depending on $p$ and on the growth rate of $\varepsilon_{n}^{2}n^{3/2}/\varphi(t_{n})$). Then there exists $n_{0}$ such that, for all $n\geq n_{0}$,
(55) $\frac{E_{\mathrm{trunc}}(n;s_{n},w_{n}^{\mathrm{in}})}{\mathcal{D}^{\mathrm{blanket}}}\leq\frac{\eta_{\mathrm{trunc}}}{4},$
where $\eta_{\mathrm{trunc}}>0$ is the prescribed accuracy knob.
Control of $\Pr\left[|E_{\mathrm{Mtrunc}}|>c_{n}/6\right]$. Recall that $Z_{i}^{\mathrm{tr}}$ is defined as $l_{\varepsilon_{n}}(Y_{i})$ conditioned on the truncation interval. Under the natural coupling where $Z_{i}^{\mathrm{tr}}=l_{\varepsilon_{n}}(Y_{i})$ whenever $l_{\varepsilon_{n}}(Y_{i})\in[s_{n},s_{n}+w_{n}^{\mathrm{in}}]$, we have
$S_{n-1}^{\mathrm{tr}}=S_{n-1}\quad\text{whenever the event }\perp^{c}\text{ occurs}.$
Consequently, on the event $\perp^{c}$, the error $E_{\mathrm{Mtrunc}}=S_{n-1}^{\mathrm{tr}}-S_{n-1}$ is identically zero. This implies the inclusion
$\bigl\{|E_{\mathrm{Mtrunc}}|>c_{n}/6\bigr\}\;\subset\;\bigl\{E_{\mathrm{Mtrunc}}\neq 0\bigr\}\;\subset\;\perp.$
Therefore, regardless of the value of $c_{n}$, we have the simple bound
(56) $\Pr\bigl[|E_{\mathrm{Mtrunc}}|>c_{n}/6\bigr]\;\leq\;\Pr[\perp]\;=\;E_{\mathrm{trunc}}(n;s_{n},w_{n}^{\mathrm{in}}).$
Using the bound (55) derived in the previous step, we immediately obtain
(57) $\Pr\bigl[|E_{\mathrm{Mtrunc}}|>c_{n}/6\bigr]\;\leq\;\frac{\eta_{\mathrm{trunc}}}{4}\,\mathcal{D}^{\mathrm{blanket}}.$
Combining (55) and (57), and recalling that $\delta_{\mathrm{trunc}}$ in (43) is a linear combination of $E_{\mathrm{trunc}}$ terms, we conclude that the total truncation contribution is bounded by
$\delta_{\mathrm{trunc}}+\Pr\bigl[|E_{\mathrm{Mtrunc}}|>c_{n}/6\bigr]\;\leq\;C_{\mathrm{trunc}}\,\eta_{\mathrm{trunc}}\,\mathcal{D}^{\mathrm{blanket}}$
for a suitable constant $C_{\mathrm{trunc}}$. Thus, the truncation contribution to the overall relative error is $O(\eta_{\mathrm{trunc}})$.
Step 3: Discretization error and choice of $h_{n}$. On the truncated interval $s_{n},s_{n}+w_{n}^{\mathrm{in}}]$ we discretize the privacy amplification random variable $Z_{i}^{\mathrm{tr}}$ into a grid with mesh size $h_{n}>0$, obtaining $Z_{i}^{\mathrm{di}}$ such that
$|Z_{i}^{\mathrm{tr}}-Z_{i}^{\mathrm{di}}|\leq\frac{h_{n}}{2}\quad\text{almost surely.}$
We denote
$S_{n-1}^{\mathrm{tr}}:=\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}},\qquad S_{n-1}^{\mathrm{di}}:=\sum_{i=2}^{M}Z_{i}^{\mathrm{di}},$
and write the discretization error of the sum as
$E^{\mathrm{disc}}:=S_{n-1}^{\mathrm{di}}-S_{n-1}^{\mathrm{tr}}=\sum_{i=2}^{M}\bigl(Z_{i}^{\mathrm{di}}-Z_{i}^{\mathrm{tr}}\bigr).$
Following the notation in the proof of Theorem 5, we set
$\Delta_{\mathrm{ES}}:=\mathbb{E}[E^{\mathrm{disc}}],\qquad U_{n}:=E^{\mathrm{disc}}-\Delta_{\mathrm{ES}}.$
By the hypothesis of Theorem 4.3 we assume that there exists a constant $K_{\mathrm{ES}}<\infty$ such that
(58) $|\Delta_{\mathrm{ES}}|\leq K_{\mathrm{ES}}\,c_{n}\qquad\text{for all sufficiently large }n.$
Reducing the constant in the definition of $c_{n}$ if necessary, we may and do assume that for $n$ large enough
(59) $|\Delta_{\mathrm{ES}}|\leq\frac{c_{n}}{12}.$
Control of $E_{\mathrm{disc}}$. Recall that in Step 2 of the proof of Theorem 5 the discretization error for a fixed hypothesis $x\in\{x_{1},x_{1}^{\prime}\}$ was defined as
$E_{\mathrm{disc}}(x;c_{n},h_{n},s_{n},w_{n}^{\mathrm{in}}):=\Pr\Bigl[\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}-\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}-\Delta_{\mathrm{ES}}>c_{n}\Bigr].$
Equivalently, if we set
$R_{n}:=\sum_{i=2}^{M}\bigl(Z_{i}^{\mathrm{tr}}-Z_{i}^{\mathrm{di}}-\Delta_{\mathrm{ES}}\bigr),$
then, we have $E_{\mathrm{disc}}(x;c_{n},h_{n},s_{n},w_{n}^{\mathrm{in}})=\Pr[R_{n}>c_{n}]$. The summands of $R_{n}$ are conditionally independent, mean zero, and bounded in absolute value by a constant multiple of $h_{n}$, while M1M-1 has mean $(n-1)\gamma$ and sub-Gaussian tails. Hence $R_{n}$ is sub-Gaussian with variance proxy of order $nh_{n}^{2}$. By Hoeffding’s inequality there exists a constant $c>0$ such that
(60) $\Pr\left[R_{n}>c_{n}\right]\leq\exp\!\left(-c\,\frac{c_{n}^{2}}{nh_{n}^{2}}\right).$
Fix a discretization accuracy knob $\eta_{\mathrm{disc}}>0$ and define
$\Lambda_{h}:=\log\frac{4}{\eta_{\mathrm{disc}}\mathcal{D}^{\mathrm{blanket}}}.$
We choose
(61) $h_{n}:=\frac{c_{n}}{\sqrt{c\,n\,\Lambda_{h}}}.$
Then (60) yields
$E_{\mathrm{disc}}(x;c_{n},h_{n},s_{n},w_{n}^{\mathrm{in}})\leq\exp(-\Lambda_{h})=\frac{\eta_{\mathrm{disc}}}{4}\,\mathcal{D}^{\mathrm{blanket}}.$
In particular, since $\delta_{\mathrm{disc}}$ in Theorem 5 is a linear combination of $E_{\mathrm{disc}}(x_{1};\cdot)$ and $E_{\mathrm{disc}}(x_{1}^{\prime};\cdot)$ with nonnegative coefficients depending only on $\varepsilon_{n}$, we obtain
$\delta_{\mathrm{disc}}\;\leq\;C_{\mathrm{disc},1}\,\eta_{\mathrm{disc}}\,\mathcal{D}^{\mathrm{blanket}}$
for some constant $C_{\mathrm{disc},1}>0$.
Under our standing regime $\varepsilon_{n}^{2}n=\Theta(\log n)$ we have $\log(1/\mathcal{D}^{\mathrm{blanket}})=\Theta(\log n)$. Consequently
$\Lambda_{h}=\log\frac{4}{\eta_{\mathrm{disc}}\mathcal{D}^{\mathrm{blanket}}}=\Theta(\log n),$
and in particular $h_{n}$ in (61) satisfies
$h_{n}=\Theta\left(\frac{c_{n}}{\sqrt{n\log n}}\right)$
up to multiplicative constants.
Control of $E_{\mathrm{Mdisc}}$. In the decomposition (42) we introduced the random variable
$E_{\mathrm{Mdisc}}:=S_{n-1}^{\mathrm{di}}-S_{n-1}^{\mathrm{tr}}=E^{\mathrm{disc}},$
which measures the effect of discretization at the level of the sum $S_{n-1}$ when passing from $S_{n-1}^{\mathrm{tr}}$ to $S_{n-1}^{\mathrm{di}}$. We control the tail probability $\Pr\left[|E_{\mathrm{Mdisc}}|>c_{n}/6\right]$ using the decomposition $E^{\mathrm{disc}}=U_{n}-\Delta_{\mathrm{ES}}$. Here, $U_{n}:=E_{\mathrm{Mdisc}}-\mathbb{E}[E_{\mathrm{Mdisc}}]=E_{\mathrm{Mdisc}}+\Delta_{\mathrm{ES}}$.
By (59), for all sufficiently large $n$ we have $|\Delta_{\mathrm{ES}}|\leq c_{n}/12$, and therefore
$\bigl\{|E_{\mathrm{Mdisc}}|>c_{n}/6\bigr\}=\bigl\{|U_{n}-\Delta_{\mathrm{ES}}|>c_{n}/6\bigr\}\subset\bigl\{|U_{n}|>c_{n}/12\bigr\}.$
Consequently,
(62) $\Pr\bigl[|E_{\mathrm{Mdisc}}|>c_{n}/6\bigr]\leq\Pr\bigl[|U_{n}|>c_{n}/12\bigr].$
The random variable $U_{n}$ is a sum of conditionally independent, mean-zero increments, each bounded by $h_{n}$ in absolute value. Applying Hoeffding’s inequality once more, we obtain
(63) $\Pr\bigl[|U_{n}|>c_{n}/12\bigr]\leq 2\exp\!\Bigl(-\kappa\,\frac{c_{n}^{2}}{nh_{n}^{2}}\Bigr),$
for some constant $\kappa>0$. With $h_{n}$ chosen as in (61), the right-hand side of (63) is bounded by
$2\exp(-\kappa\Lambda_{h})\leq\frac{\eta_{\mathrm{disc}}}{4}\,\mathcal{D}^{\mathrm{blanket}}$
for all sufficiently large $n$, after possibly adjusting the numerical constant in the definition of $\Lambda_{h}$. Combining this with (62), we conclude that
(64) $\Pr\bigl[|E_{\mathrm{Mdisc}}|>c_{n}/6\bigr]\leq\frac{\eta_{\mathrm{disc}}}{4}\,\mathcal{D}^{\mathrm{blanket}}.$
Putting together the bounds for $E_{\mathrm{disc}}$ and $E_{\mathrm{Mdisc}}$, and recalling that the discretization contribution in (43) is a linear combination of these quantities with constants depending only on $\varepsilon_{n}$ and $\gamma$, we obtain that there exists a constant $C_{\mathrm{disc}}>0$ such that
$\delta_{\mathrm{disc}}+\Pr\bigl[|E_{\mathrm{Mdisc}}|>c_{n}/6\bigr]\leq C_{\mathrm{disc}}\,\eta_{\mathrm{disc}}\,\mathcal{D}^{\mathrm{blanket}}.$
In particular, the discretization contribution to the overall relative error is $O(\eta_{\mathrm{disc}})$.
Step 4: Aliasing error and choice of $w_{n}^{\mathrm{out}}$. We now control the aliasing (wrap-around) error arising from the finite FFT window. Recall that after truncation and discretization we consider the sum
$S_{n-1}^{\mathrm{di}}:=\sum_{i=2}^{M}Z_{i}^{\mathrm{di}},$
where the $Z_{i}^{\mathrm{di}}$ are i.i.d. copies of the discretized summand on $s_{n},s_{n}+w_{n}^{\mathrm{in}}]$. By construction, each $Z_{i}^{\mathrm{di}}$ is supported in an interval of radius at most $\tau_{n}$, hence $|Z_{i}^{\mathrm{di}}|\leq\tau_{n}$ almost surely. Assumption 2.8 further implies that $\mathrm{Var}(Z_{i}^{\mathrm{di}})$ is bounded away from zero and infinity uniformly in $n$, so that
$v_{n}:=\mathrm{Var}\bigl(S_{n-1}^{\mathrm{di}}\bigr)=\mathrm{Var}\Bigl(\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}\Bigr)=\Theta(n).$
Let $\mu_{n}^{\mathrm{di}}:=\mathbb{E}[S_{n-1}^{\mathrm{di}}]$. By Bernstein’s inequality, for any $u>0$,
(65) $\Pr\bigl[|S_{n-1}^{\mathrm{di}}-\mu_{n}^{\mathrm{di}}|>u\bigr]\leq 2\exp\!\left(-\frac{u^{2}}{2\bigl(v_{n}+\tau_{n}u/3\bigr)}\right).$
From the choice of $p$ in Step 2 we have
(66) $\tau_{n}=o(\sqrt{n}),$
so that for $u$ of order $\sqrt{v_{n}\Lambda_{w}}$ the second term in the denominator of (65) is negligible compared to $v_{n}$. More precisely, define
$\Lambda_{w}:=\log\frac{4}{\eta_{\mathrm{alias}}\mathcal{D}^{\mathrm{blanket}}},$
where $0<\eta_{\mathrm{alias}}<1$ is a prescribed accuracy parameter and $\mathcal{D}^{\mathrm{blanket}}$ is the blanket divergence at level $\varepsilon_{n}$. Set
$u_{n}:=\sqrt{C_{w}\,v_{n}\,\Lambda_{w}}$
for a sufficiently small constant $C_{w}>0$. Then, using $v_{n}=\Theta(n)$ and (66), there exist constants $c_{1},c_{2}>0$ such that for all large $n$,
$v_{n}+\tau_{n}u_{n}/3\leq c_{1}v_{n}\qquad\text{and}\qquad\frac{u_{n}^{2}}{2\bigl(v_{n}+\tau_{n}u_{n}/3\bigr)}\geq c_{2}\Lambda_{w}.$
Inserting this into (65) yields
(67) $\Pr\bigl[|S_{n-1}^{\mathrm{di}}-\mu_{n}^{\mathrm{di}}|>u_{n}\bigr]\leq 2\exp(-c_{2}\Lambda_{w}).$
Choice of the outer window and control of $E_{\mathrm{alias}}$. In the FFT computation we embed the PMF of $S_{n-1}^{\mathrm{di}}$ into a periodic grid of period $w_{n}^{\mathrm{out}}$. Aliasing can occur only if $S_{n-1}^{\mathrm{di}}$ exits the FFT window. We choose the outer window so as to contain, with high probability, the interval $\mu_{n}^{\mathrm{di}}-u_{n},\mu_{n}^{\mathrm{di}}+u_{n}]$. Concretely, we set
(68) $w_{n}^{\mathrm{out}}:=2u_{n}+h_{n}=\Theta\bigl(\sqrt{v_{n}\Lambda_{w}}\bigr)=\Theta\bigl(\sqrt{n\Lambda_{w}}\bigr).$
By construction, if $|S_{n-1}^{\mathrm{di}}-\mu_{n}^{\mathrm{di}}|\leq u_{n}$ then $S_{n-1}^{\mathrm{di}}$ lies in the central interval of length $w_{n}^{\mathrm{out}}$ and no aliasing occurs. Thus the aliasing probability satisfies
(69) $E_{\mathrm{alias}}(n;h_{n},s_{n},w_{n}^{\mathrm{in}},w_{n}^{\mathrm{out}}):=\Pr\bigl[\text{aliasing occurs}\bigr]\leq\Pr\bigl[|S_{n-1}^{\mathrm{di}}-\mu_{n}^{\mathrm{di}}|>u_{n}\bigr].$
Combining (67) and (69) and absorbing the factor $2$ into the constants, we obtain
$E_{\mathrm{alias}}(n;h_{n},s_{n},w_{n}^{\mathrm{in}},w_{n}^{\mathrm{out}})\leq\exp(-c_{2}\Lambda_{w})=\frac{\eta_{\mathrm{alias}}}{4}\,\mathcal{D}^{\mathrm{blanket}}$
for all $n$ sufficiently large. In particular, since the aliasing error term $\delta_{\mathrm{alias}}$ in Theorem 5 is a linear combination of $E_{\mathrm{alias}}(x_{1};\cdot)$ and $E_{\mathrm{alias}}(x_{1}^{\prime};\cdot)$ with nonnegative coefficients depending only on $\varepsilon_{n}$, we have
(70) $\delta_{\mathrm{alias}}\;\leq\;C_{\mathrm{alias},1}\,\eta_{\mathrm{alias}}\,\mathcal{D}^{\mathrm{blanket}}$
for some constant $C_{\mathrm{alias},1}>0$.
Moreover, by the asymptotic expansion of $\mathcal{D}^{\mathrm{blanket}}$ in Lemma 3.2 and our regime $\varepsilon_{n}^{2}n=\Theta(\log n)$, the quantity $\log(1/\mathcal{D}^{\mathrm{blanket}})$ grows on the order of $\log n$. Hence
$\Lambda_{w}=\log\frac{4}{\eta_{\mathrm{alias}}\mathcal{D}^{\mathrm{blanket}}}=\Theta(\log n),$
and we obtain
(71) $w_{n}^{\mathrm{out}}=\Theta(\sqrt{n\log n}).$
Control of $E_{\mathrm{Malias}}$. In the decomposition (42) we introduced the wrap-around error at the level of the sum,
$E_{\mathrm{Malias}}:=\widetilde{S}_{n-1}-S_{n-1}^{\mathrm{di}},$
where $\widetilde{S}_{n-1}$ denotes the periodic extension of $S_{n-1}^{\mathrm{di}}$ induced by the FFT grid. By definition, $E_{\mathrm{Malias}}=0$ whenever no aliasing occurs, so that
$\bigl\{|E_{\mathrm{Malias}}|>0\bigr\}\subset\bigl\{\text{aliasing occurs}\bigr\}.$
In particular, for any $c_{n}>0$,
(72) $\Pr\bigl[|E_{\mathrm{Malias}}|>c_{n}/6\bigr]\leq\Pr\bigl[|E_{\mathrm{Malias}}|>0\bigr]\leq E_{\mathrm{alias}}(n;h_{n},s_{n},w_{n}^{\mathrm{in}},w_{n}^{\mathrm{out}}).$
Combining (72) with the bound on $E_{\mathrm{alias}}$ above, we obtain
(73) $\Pr\bigl[|E_{\mathrm{Malias}}|>c_{n}/6\bigr]\leq\frac{\eta_{\mathrm{alias}}}{4}\,\mathcal{D}^{\mathrm{blanket}}.$
Putting together (70) and (73), and recalling that the aliasing contribution in (43) is built from $\delta_{\mathrm{alias}}$ and $\Pr[|E_{\mathrm{Malias}}|>c_{n}/6]$ with constants depending only on $\varepsilon_{n}$ and $\gamma$, we conclude that there exists a constant $C_{\mathrm{alias}}>0$ such that
$\delta_{\mathrm{alias}}+\Pr\bigl[|E_{\mathrm{Malias}}|>c_{n}/6\bigr]\;\leq\;C_{\mathrm{alias}}\,\eta_{\mathrm{alias}}\,\mathcal{D}^{\mathrm{blanket}}.$
In particular, the aliasing contribution to the overall relative error is $O(\eta_{\mathrm{alias}})$.
Step 5: Relative error bound. Combining the four steps above yields
$\bigl|\widehat{D}_{n}-\mathcal{D}^{\mathrm{blanket}}\bigr|\;\leq\;C\bigl(\eta_{\mathrm{main}}+\eta_{\mathrm{trunc}}+\eta_{\mathrm{disc}}+\eta_{\mathrm{alias}}\bigr)\,\mathcal{D}^{\mathrm{blanket}},$
for some constant $C<\infty$, which proves the first claim of the theorem.
Step 6: FFT grid size and running time. By construction, the FFT grid has period $2w_{n}^{\mathrm{out}}$ and spacing $h_{n}$. Using the bounds $w_{n}^{\mathrm{out}}=\Theta(\sqrt{n\log n})$ and $h_{n}=\Theta(c_{n}/\sqrt{n\log n})$, the grid size $N$ satisfies
$N=\frac{2w_{n}^{\mathrm{out}}}{h_{n}}=O\left(\frac{\sqrt{n\log n}}{c_{n}/\sqrt{n\log n}}\right)=O\left(\frac{n\log n}{c_{n}}\right).$
Recalling from Step 1 that $c_{n}=\Theta(\eta_{\mathrm{main}}/\log n)$ in the regime $\varepsilon_{n}=\Theta(\sqrt{(\log n)/n})$, we obtain
$N=O\left(\frac{n(\log n)^{2}}{\eta_{\mathrm{main}}}\right).$
Since the FFT-based algorithm runs in time $O(N\log N)$, we conclude that the overall time complexity is
$O(N\log N)=O\left(\frac{n}{\eta_{\mathrm{main}}}\left(\log\frac{n}{\eta_{\mathrm{main}}}\right)^{3}\right),$
where the hidden constants depend only on $\gamma,\sigma,\eta_{\mathrm{trunc}},\eta_{\mathrm{disc}},\eta_{\mathrm{alias}}$, and we have omitted $\log\log$ factors. This completes the proof.
A.7.1. Proof of Lemma A.10
See A.10
Proof.
Fix $x\in\mathcal{X}$ and suppress the dependence on $x$ in the notation. We work with the sequence $(\varepsilon_{n})$ from Lemma 3.2.
Step 1: Standardization and reduction to a window event. Define
$T_{n-1}:=\frac{S_{n-1}-(n-1)\mu_{n}}{\sigma_{\varepsilon_{n}}\sqrt{n-1}},\qquad t_{n}:=-\frac{\mu_{n}\sqrt{n-1}}{\sigma_{\varepsilon_{n}}}.$
Write
$\eta_{n}(x):=\frac{x}{\sigma_{\varepsilon_{n}}\sqrt{n-1}},$
and
$m_{n}(Z_{1}):=t_{n}-\eta_{n}(Z_{1})-\eta_{n}(\Delta_{\mathrm{ES}}).$
Exactly as before we obtain
(74) $\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c_{n}\,\big|\,Z_{1}\bigr]=\Pr\bigl[m_{n}(Z_{1})-\eta_{n}\left(\frac{3c_{n}}{2}\right)<T_{n-1}\leq m_{n}(Z_{1})+\eta_{n}\left(\frac{3c_{n}}{2}\right)\bigr].$
By the moderate-deviation scaling assumed for $\varepsilon_{n}$ (the same as in Lemma 3.2), we have $t_{n}=O(\sqrt{\log n})$ and $\eta_{n}\left(\frac{3c_{n}}{2}\right)=O(c_{n}/\sqrt{n})$.
Step 2: Gaussian window upper bound for $T_{n-1}$. We claim that there exist constants $C_{1},C_{2}<\infty$ (depending only on Assumption 2.8) such that, for all sufficiently large $n$ and all $m,\eta$ with
(75) $|m|\leq C_{1}\sqrt{\log n},\qquad|\eta|\leq C_{1}/\sqrt{n},$
we have
(76) $\Pr\left[m-\eta<T_{n-1}\leq m+\eta\right]\;\leq\;C_{2}\,\eta\,\varphi(m).$
Under the non-lattice assumption (i.e., the assumption of Theorem 4.3) [^10] The non-lattice assumption is essential in our setting because the band width $c_{n}$ tends to zero, so that the standardized window $(m-\eta,m+\eta)$ has width $2\eta=\Theta(c_{n}/(\sigma_{\varepsilon_{n}}\sqrt{n}))$. In the lattice case the mesh of the normalized sum is of order $1/\sqrt{n}$, and when $c_{n}\to 0$ the window contains at most $O(1)$ lattice points. Consequently the window probability remains of order $n^{-1/2}\varphi(m)$, rather than shrinking linearly with the window width $\eta=\Theta(c_{n}/(\sigma_{\varepsilon_{n}}\sqrt{n}))$, and it cannot in general be bounded by a constant multiple of $\eta\,\varphi(m)=\Theta(c_{n}n^{-1/2}\varphi(m))$. The non-lattice condition rules out this obstruction and allows us to invoke local limit/Edgeworth theory down to such small scales, yielding a window probability of order $\eta\,\varphi(m)$. in addition to Assumption 2.8 and the truncation argument of Lemma A.6, the triangular array $(Z_{i,n}^{(h)})$ satisfies the hypotheses of the triangular-array Edgeworth expansion (Theorem A.1) uniformly in $n$. In particular, for any fixed order $s\geq 3$ we can write the distribution function of $T_{n-1}$ as
$F_{n-1}(u):=\Pr\bigl[T_{n-1}\leq u\bigr]=\Phi(u)+\sum_{j=1}^{s-2}(n-1)^{-j/2}P_{j,n-1}(u)\,\varphi(u)+R_{n,s}(u),$
where $P_{j,n-1}$ are polynomials of degree at most s2s-2 whose coefficients are uniformly bounded in $n$, and the remainder satisfies
$\sup_{u\in\mathbb{R}}\bigl|R_{n,s}(u)\bigr|=O\bigl((n-1)^{-(s-2)/2}\bigr).$
Fix $m,\eta\in\mathbb{R}$ satisfying the moderate-deviation condition
(77) $|m|\leq C_{1}\sqrt{\log n},\qquad|\eta|\leq\frac{C_{1}}{\sqrt{n}}$
for some fixed constant $C_{1}>0$. We then have
$\Pr\bigl[m-\eta<T_{n-1}\leq m+\eta\bigr]$ $=F_{n-1}(m+\eta)-F_{n-1}(m-\eta)$
(78) $=I_{0,n}+\sum_{j=1}^{s-2}I_{j,n}+I_{R,n},$
where
$I_{0,n}$ $:=\Phi(m+\eta)-\Phi(m-\eta),$
$I_{j,n}$ $:=(n-1)^{-j/2}\Bigl[P_{j,n-1}(m+\eta)\,\varphi(m+\eta)-P_{j,n-1}(m-\eta)\,\varphi(m-\eta)\Bigr],$
$I_{R,n}$ $:=R_{n,s}(m+\eta)-R_{n,s}(m-\eta).$
Gaussian main term. By the mean-value form of Taylor’s theorem applied to $\Phi$ at the point $m$, there exists $\theta_{n}\in(-1,1)$ such that
$I_{0,n}=\Phi(m+\eta)-\Phi(m-\eta)=2\eta\,\varphi(m+\theta_{n}\eta).$
Under (77) we have $|m+\theta_{n}\eta|\leq C_{1}\sqrt{\log n}+o(1)$, and on this region the Gaussian density is comparable at $m$ and $m+\theta_{n}\eta$: there exists $C^{\prime}_{1}<\infty$ such that
$\varphi(m+\theta_{n}\eta)\leq C^{\prime}_{1}\,\varphi(m)$
for all $n$ large enough. Hence
(79) $|I_{0,n}|\leq 2C^{\prime}_{1}\,\eta\,\varphi(m).$
Edgeworth correction terms. For each $j\geq 1$, the polynomial $P_{j,n-1}$ has at most polynomial growth in $u$; in particular, there exist constants $C_{j}<\infty$ and integers $k_{j}\geq 0$ such that
$|P_{j,n-1}(u)|\leq C_{j}\bigl(1+|u|^{k_{j}}\bigr)\qquad\text{for all }u\in\mathbb{R},\,n\geq 2.$
Using $|m|\leq C_{1}\sqrt{\log n}$ and $|\eta|\leq C_{1}/\sqrt{n}$ from (77), we have $|m\pm\eta|\leq C_{2}\sqrt{\log n}$ and hence
$|P_{j,n-1}(m\pm\eta)|\leq C_{j}^{\prime}\bigl(1+(\log n)^{k_{j}/2}\bigr)\leq C_{j}^{\prime\prime}(\log n)^{K_{j}}$
for some $K_{j}$ and all $n$ large enough. Moreover $\varphi(m\pm\eta)\leq C\,\varphi(m)$ on this region. Therefore
$|I_{j,n}|\leq(n-1)^{-j/2}\cdot 2|\eta|\cdot C_{j}^{\prime\prime}(\log n)^{K_{j}}\,\sup_{u\in[m-\eta,m+\eta]}\varphi(u)\leq C_{j}^{\prime\prime\prime}\,\eta\,\varphi(m)\,(n-1)^{-j/2}(\log n)^{K_{j}}.$
Summing over $1\leq j\leq s-2$ and using $j\geq 1$ gives
(80) $\sum_{j=1}^{s-2}|I_{j,n}|\leq C_{2}^{\prime}\,\eta\,\varphi(m)\,n^{-1/2}(\log n)^{K}\leq C_{3}^{\prime}\,\eta\,\varphi(m),$
where $K:=\max_{j}K_{j}$ and the polynomial factor in $\log n$ has been absorbed into the constant $C_{3}^{\prime}$.
Remainder term. For the remainder we simply use the uniform bound on $R_{n,s}$:
$|I_{R,n}|\leq 2\sup_{u\in\mathbb{R}}|R_{n,s}(u)|=O\bigl((n-1)^{-(s-2)/2}\bigr).$
On the other hand, under the scaling (77) we have
$\eta\,\varphi(m)=\Theta\left(\frac{1}{\sqrt{n}}\,\exp\!\bigl(-\tfrac{1}{2}m^{2}\bigr)\right)\geq c\,n^{-1/2-C_{4}},$
for some $c,C_{4}>0$ depending only on $C_{1}$, since $|m|\leq C_{1}\sqrt{\log n}$ implies $\varphi(m)\geq n^{-C_{4}}$. Thus, by choosing $s$ large enough we ensure
$(n-1)^{-(s-2)/2}=o\bigl(\eta\,\varphi(m)\bigr),$
and therefore, for all sufficiently large $n$,
(81) $|I_{R,n}|\leq C_{4}^{\prime}\,\eta\,\varphi(m).$
Conclusion. Combining (78), (79), (80), and (81), we obtain
$\Pr\bigl[m-\eta<T_{n-1}\leq m+\eta\bigr]\leq C\,\eta\,\varphi(m)$
for some finite constant $C<\infty$ (independent of $n$, $m$, and $\eta$) and all sufficiently large $n$, whenever (77) holds. This proves (76) in the continuous case.
Step 3: Applying the window bound to $m_{n}(Z_{1}),\eta_{n}\left(\frac{3c_{n}}{2}\right)$. By Assumption 2.8 we have $\mathbb{E}[|Z_{1}|^{r}]<\infty$ for all integers $r\geq 1$. Fix $\beta\in(0,1/2)$ and write
$A_{n}:=\{|Z_{1}|\leq n^{\beta}\},\qquad A_{n}^{c}:=\{|Z_{1}|>n^{\beta}\}.$
On $A_{n}$ we have
$|\eta_{n}(Z_{1})|=\frac{|Z_{1}|}{\sigma_{\varepsilon_{n}}\sqrt{n-1}}\leq Cn^{\beta-1/2}=o(1),$
and the assumption $|\Delta_{\mathrm{ES}}|=O\left(\frac{\sigma_{\varepsilon_{n}}\sqrt{n}}{\sqrt{\log n}}\right)$ implies $|\eta_{n}(\Delta_{\mathrm{ES}})|=O\left(\frac{1}{\sqrt{\log n}}\right)\to 0$. Hence on $A_{n}$,
$|m_{n}(Z_{1})|=\bigl|t_{n}-\eta_{n}(Z_{1})-\eta_{n}(\Delta_{\mathrm{ES}})\bigr|\leq|t_{n}|+o(1)=O(\sqrt{\log n}),$
and for all large $n$ we may assume that (77) holds with $m=m_{n}(Z_{1})$ and $\eta=\eta_{n}\left(\frac{3c_{n}}{2}\right)$.
Therefore, by (76),
$\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c_{n}\,\big|\,Z_{1}\bigr]\mathbf{1}\{A_{n}\}$
$\qquad\leq C_{2}\,\eta_{n}\left(\frac{3c_{n}}{2}\right)\,\varphi\bigl(m_{n}(Z_{1})\bigr)\mathbf{1}\{A_{n}\}.$
Since $|m_{n}(Z_{1})-t_{n}|=o(1)$ on $A_{n}$ and $|t_{n}|=O(\sqrt{\log n})$, a direct calculation of the likelihood ratio yields
$\varphi\bigl(m_{n}(Z_{1})\bigr)\mathbf{1}\{A_{n}\}\leq C_{5}\,\varphi(t_{n})\mathbf{1}\{A_{n}\}$
for some constant $C_{5}<\infty$ and all large $n$. Thus
(82) $\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c_{n}\,\big|\,Z_{1}\bigr]\mathbf{1}\{A_{n}\}\leq C_{2}C_{5}\,\eta_{n}\left(\frac{3c_{n}}{2}\right)\,\varphi(t_{n})\mathbf{1}\{A_{n}\}.$
On the complement $A_{n}^{c}$ we only use the trivial bound
$0\leq\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c_{n}\,\big|\,Z_{1}\bigr]\mathbf{1}\{A_{n}^{c}\}\leq\mathbf{1}\{A_{n}^{c}\},$
and by Markov’s inequality and the moment bounds,
$\Pr[A_{n}^{c}]=\Pr\bigl[|Z_{1}|>n^{\beta}\bigr]=O(n^{-K})$
for any prescribed $K>0$ if we choose the moment order large enough.
Step 4: Taking expectation over $Z_{1}$. Taking expectation of (82) and adding the contribution of $A_{n}^{c}$, we obtain
$\mathbb{E}_{Z_{1}\sim\mathcal{R}_{x}}\Bigl[\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c_{n}\,\big|\,Z_{1}\bigr]\Bigr]$
$\qquad\leq C_{2}C_{5}\,\eta_{n}\left(\frac{3c_{n}}{2}\right)\,\varphi(t_{n})+O\bigl(\Pr[A_{n}^{c}]\bigr).$
Since $\eta_{n}\left(\frac{3c_{n}}{2}\right)=\Theta(c_{n}/(\sigma_{\varepsilon_{n}}\sqrt{n-1}))$ and $\Pr[A_{n}^{c}]=O(n^{-K})$ can be made of smaller order than $\eta_{n}\left(\frac{3c_{n}}{2}\right)\varphi(t_{n})$ by choosing $K$ large enough, we may absorb the tail contribution into the constant. Thus there exists $C,C^{\prime}<\infty$ (independent of $x$ and $n$) such that for all large $n$,
$\mathbb{E}_{Z_{1}\sim\mathcal{R}_{x}}\Bigl[\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c_{n}\,\big|\,Z_{1}\bigr]\Bigr]\leq C\,\eta_{n}\left(\frac{3c_{n}}{2}\right)\,\varphi(t_{n})=C^{\prime}\,\frac{c_{n}}{\sigma_{\varepsilon_{n}}\sqrt{n-1}}\,\varphi(t_{n}).$
This is exactly (44), which completes the proof. ∎
Appendix B The Computation of Shuffle Indices
B.1. Generalized Gaussian Mechanisms
Here, we consider the generalized Gaussian mechanism (i.e., Definition 2.9).
Lower Shuffle Index
From the definition, the variance term and the blanket mass in the lower shuffle index are
$\sigma_{\mathrm{BG},\beta}^{2}(x_{1},x_{1}^{\prime}):=\gamma_{\beta,\alpha}\left\{\int_{-\infty}^{1/2}\frac{\bigl(\phi_{x_{1}}^{\beta}(y)-\phi_{x_{1}^{\prime}}^{\beta}(y)\bigr)^{2}}{\phi_{1}^{\beta}(y)}\,dy\;+\;\int_{1/2}^{\infty}\frac{\bigl(\phi_{x_{1}}^{\beta}(y)-\phi_{x_{1}^{\prime}}^{\beta}(y)\bigr)^{2}}{\phi_{0}^{\beta}(y)}\,dy\right\},$
$\gamma_{\beta,\alpha}=\frac{1}{\Gamma(1/\beta)}\int_{(1/(2\alpha))^{\beta}}^{\infty}t^{1/\beta-1}e^{-t}\,dt=\frac{\Gamma\!\bigl(1/\beta,\,(1/(2\alpha))^{\beta}\bigr)}{\Gamma(1/\beta)},$
where $\Gamma(s,x)$ denotes the upper incomplete gamma function. We can see integrands are smooth and rapidly decaying. In particular, for each fixed $\beta>0$ and each pair $(x_{1},x_{1}^{\prime})$, the quantity $\sigma_{\mathrm{BG},\beta}^{2}(x_{1},x_{1}^{\prime})$ (and hence $\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})$) can be evaluated efficiently by standard one-dimensional numerical quadrature (in the special cases $\beta=2$ (Gaussian mechanism) and $\beta=1$ (Laplace mechanism), the relevant integrals admit closed-form expressions). As a consequence, searching over $(x_{1},x_{1}^{\prime})$ for the lower shuffle index is computationally inexpensive [^11] In all our numerical experiments for generalized Gaussian mechanisms (including the cases $\beta=1$ and $\beta=2$), the pair $(x_{1},x_{1}^{\prime})=(0,1)$ consistently attains (or appears to attain) the infimum of the lower shuffle index $\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})$. This is also intuitively plausible, since $(0,1)$ maximizes the separation between two inputs in the unit interval. However, a general theoretical proof that $(0,1)$ is the unique global minimizer of $\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})$ for all admissible parameters $(\alpha,\beta)$ would require a delicate analysis beyond the scope of this work, and we therefore do not claim such a statement as a theorem..
Figure 2(a) illustrates the lower shuffle index $\underline{\chi}_{\mathrm{lo}}$ for several values of the shape parameter $\beta$. Even when the underlying local randomizers are calibrated to have the same variance (and hence the same accuracy for mean estimation), they can exhibit different shuffle indices.
(a) Lower shuffle index for several generalized Gaussian mechanisms.
(b) Ratio of lower and upper shuffle indices.
Figure 2. Shuffle indices for generalized Gaussian mechanisms with different shape parameters $\beta$.
Upper Shuffle Index
For the upper shuffle index, we instead take the reference distribution to be one of the local randomizers $\mathcal{R}_{x}$ as in Su et al. (Su et al., 2025), and only the variance is needed in our analysis. From the definition, the variance is
(83) $\sigma_{x,\beta}^{2}(x_{1},x_{1}^{\prime}):=\int_{-\infty}^{\infty}\frac{\bigl(\phi_{x_{1}}^{\beta}(y)-\phi_{x_{1}^{\prime}}^{\beta}(y)\bigr)^{2}}{\phi_{x}^{\beta}(y)}\,dy\;=\;\alpha\frac{\beta}{2\alpha\,\Gamma(1/\beta)}\,F_{\beta}(a_{1},a_{1}^{\prime}),$
where
$F_{\beta}(a_{1},a_{1}^{\prime}):=\int_{-\infty}^{\infty}\frac{\bigl(e^{-|z-a_{1}|^{\beta}}-e^{-|z-a_{1}^{\prime}|^{\beta}}\bigr)^{2}}{e^{-|z|^{\beta}}}\,dz,\qquad a_{1}:=\frac{x_{1}-x}{\alpha},\qquad a_{1}^{\prime}:=\frac{x_{1}^{\prime}-x}{\alpha}$
Although the definition of $\underline{\chi}_{\mathrm{up}}$ involves a three-point search over $(x_{1},x_{1}^{\prime},x)$, the fact that $\{\phi_{\mu}^{\beta}\}_{\mu\in\mathbb{R}}$ forms a location family implies that $\sigma_{x,\beta}^{2}(x_{1},x_{1}^{\prime})$ in (83) depends on $(x_{1},x_{1}^{\prime},x)$ only through the two offsets $(a_{1},a_{1}^{\prime})=((x_{1}-x)/\alpha,(x_{1}^{\prime}-x)/\alpha)$. In particular, the optimization can be reduced to a two-dimensional search over $(a_{1},a_{1}^{\prime})$. As in the case of the lower shuffle index, the variance $\sigma_{x,\beta}^{2}(x_{1},x_{1}^{\prime})$ can be efficiently evaluated by standard one-dimensional numerical quadrature (or even in closed form for the special cases $\beta=1,2$), so this search remains computationally inexpensive.
B.2. $k$-$R$
We next specialize the shuffle indices to the $k$-$R$. For an input $x\in[k]:=\{1,\dots,k\}$, the output distribution is
$\phi_{x}(y)=\Pr[\mathcal{R}(x)=y]=\begin{cases}p,&y=x,\\[2.0pt] q,&y\neq x,\end{cases}\qquad p=\dfrac{\mathrm{e}^{\varepsilon_{0}}}{\mathrm{e}^{\varepsilon_{0}}+k-1},\quad q=\dfrac{1}{\mathrm{e}^{\varepsilon_{0}}+k-1}.$
The blanket distribution is uniform on $[k]$: $\phi_{\mathrm{BG}}(y)=\frac{1}{k},\gamma=\sum_{y}\min_{x}\phi_{x}(y)=kq.$
Lower shuffle index. For any distinct pair $x_{1}\neq x_{1}^{\prime}\in[k]$, the privacy amplification random variable with respect to the blanket distribution is
$l_{0}(y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{BG}})=\frac{\phi_{x_{1}}(y)-\phi_{x_{1}^{\prime}}(y)}{\phi_{\mathrm{BG}}(y)}=k\bigl(\phi_{x_{1}}(y)-\phi_{x_{1}^{\prime}}(y)\bigr).$
A simple case distinction shows that, for $Y\sim\mathcal{R}_{\mathrm{BG}}$,
$l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{BG}})=\begin{cases}k(p-q),&Y=x_{1},\\ -k(p-q),&Y=x_{1}^{\prime},\\ 0,&\text{otherwise},\end{cases}$
so that $\mathbb{E}[l_{0}(Y)]=0$ and $\sigma_{\mathrm{BG}}^{2}(x_{1},x_{1}^{\prime}):=\mathrm{Var}\bigl(l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{BG}})\bigr)=2k(p-q)^{2}.$ By definition,
$\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})=\sqrt{\frac{\gamma}{\sigma_{\mathrm{BG}}^{2}(x_{1},x_{1}^{\prime})}}=\sqrt{\frac{q}{2(p-q)^{2}}}=\sqrt{\frac{\mathrm{e}^{\varepsilon_{0}}+k-1}{2\bigl(\mathrm{e}^{\varepsilon_{0}}-1\bigr)^{2}}}.$
In particular, $\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})$ does not depend on the particular choice of the distinct pair $(x_{1},x_{1}^{\prime})$, and hence
$\underline{\chi}_{\mathrm{lo}}=\inf_{x_{1}\neq x_{1}^{\prime}}\chi_{\mathrm{lo}}(x_{1},x_{1}^{\prime})=\sqrt{\frac{q}{2(p-q)^{2}}}.$
Upper shuffle index. For the upper shuffle index we take the reference distribution to be one of the local randomizers $\mathcal{R}_{x}$. Fix $x_{1}\neq x_{1}^{\prime}$, and consider
$l_{0}(y;x_{1},x_{1}^{\prime},\mathcal{R}_{x})=\frac{\phi_{x_{1}}(y)-\phi_{x_{1}^{\prime}}(y)}{\phi_{x}(y)},\qquad Y\sim\mathcal{R}_{x}.$
By symmetry, the variance only depends on how $x$ is positioned relative to $\{x_{1},x_{1}^{\prime}\}$.
  • If $x=x_{1}$ (or symmetrically $x=x_{1}^{\prime}$), then
    $l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{x_{1}})=\begin{cases}\dfrac{p-q}{p},&Y=x_{1},\\[4.0pt] -\dfrac{p-q}{q},&Y=x_{1}^{\prime},\\[4.0pt] 0,&\text{otherwise},\end{cases}$
    so that $\mathrm{Var}\bigl(l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{x_{1}})\bigr)=(p-q)^{2}\Bigl(\frac{1}{p}+\frac{1}{q}\Bigr).$
  • If $k\geq 3$ and $x\notin\{x_{1},x_{1}^{\prime}\}$, then $\mathrm{Var}\bigl(l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{x})\bigr)=2\,\frac{(p-q)^{2}}{q}.$
Since for $p>q$ we have $2\,\frac{(p-q)^{2}}{q}\;>\;(p-q)^{2}\Bigl(\frac{1}{p}+\frac{1}{q}\Bigr),$ the variance is maximized (and hence $\chi$ is minimized) by choosing $x\notin\{x_{1},x_{1}^{\prime}\}$ whenever $k\geq 3$. Because $\gamma=1$ for $\mathcal{R}_{x}$, we obtain, for $k\geq 3$,
$\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime})=\inf_{x\in[k]}\sqrt{\frac{1}{\mathrm{Var}(l_{0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{x}))}}=\sqrt{\frac{q}{2(p-q)^{2}}}$
Thus, $\chi_{\mathrm{up}}(x_{1},x_{1}^{\prime})$ also does not depend on the particular choice of the distinct pair $(x_{1},x_{1}^{\prime})$, and hence, for $k$-randomized response with $k\geq 3$, the lower and upper shuffle indices coincide:
$\underline{\chi}_{\mathrm{up}}=\underline{\chi}_{\mathrm{lo}}=\sqrt{\frac{q}{2(p-q)^{2}}}=\sqrt{\frac{\mathrm{e}^{\varepsilon_{0}}+k-1}{2\bigl(\mathrm{e}^{\varepsilon_{0}}-1\bigr)^{2}}}.$
Appendix C Implementations of FFT algorithms
C.1. Parameters
The error terms in Theorem 5 admit explicit and rigorous upper bounds. Let
$M\sim 1+\mathrm{Bin}(n-1,\gamma)\quad\text{as}\quad M=1+\sum_{i=2}^{n}B_{i},$
where $B_{2},\dots,B_{n}$ are i.i.d. $\mathrm{Bernoulli}(\gamma)$ variables, independent of the blanket samples. In particular,
$\sum_{i=2}^{M}Z_{i}\;\stackrel{{\scriptstyle d}}{{=}}\;\sum_{i=2}^{n}B_{i}Z_{i},$
so every error event can be expressed as a tail event of a sum of independent variables of the form $B_{i}Z_{i}$.
(i) Truncation error $E_{\mathrm{trunc}}$. Let $q:=\Pr[l(Y)\notin[s,s+w^{\mathrm{in}}]]=F_{\mathrm{ref}}(s)+1-F_{\mathrm{ref}}(s+w^{\mathrm{in}})$, where $F_{\mathrm{ref}}$ is the blanket CDF. Then the truncation error is
$E_{\mathrm{trunc}}:=\Pr\bigl[\perp\bigr]=1-\mathbb{E}\bigl[(1-q)^{M-1}\bigr]=1-(1-\gamma q)^{n-1},$
which is computable exactly from $F_{\mathrm{ref}}$ and the numerical parameters $(n,\gamma,s,w^{\mathrm{in}})$. In particular, $E_{\mathrm{trunc}}\leq(n-1)\gamma q$ by the union bound.
(ii) Discretization error $E_{\mathrm{disc}}$. Let $D_{i}:=Z_{i}^{\mathrm{tr}}-Z_{i}^{\mathrm{di}}$ and observe that, by construction, $|D_{i}|\leq h/2$ almost surely. Using the selector representation we may write
$\sum_{i=2}^{M}D_{i}\;\stackrel{{\scriptstyle d}}{{=}}\;\sum_{i=2}^{n}B_{i}D_{i},$
where $B_{2},\dots,B_{n}$ are i.i.d. $\mathrm{Bernoulli}(\gamma)$ variables independent of $(D_{i})_{i\geq 2}$. Define
$X_{i}:=B_{i}D_{i}-\mathbb{E}[B_{i}D_{i}],\qquad S_{\mathrm{disc}}:=\sum_{i=2}^{n}X_{i}.$
Then the variables $X_{i}$ are independent and bounded as
$|X_{i}|\;\leq\;K_{\mathrm{disc}}:=\frac{h}{2}\quad\text{a.s.}$
Moreover, since $|D_{i}|\leq h/2$ we have
$\mathrm{Var}(X_{i})=\mathrm{Var}(B_{i}D_{i})\leq\gamma\,\mathbb{E}[D_{i}^{2}]\leq\gamma\,\frac{h^{2}}{4}=:v_{\mathrm{disc}}^{2}.$
Hence Bernstein’s inequality yields, for any $a>0$,
$\Pr\bigl[|S_{\mathrm{disc}}|>a\bigr]\;\leq\;2\exp\!\left(-\frac{a^{2}}{2(n-1)v_{\mathrm{disc}}^{2}+\frac{2}{3}K_{\mathrm{disc}}a}\right)=2\exp\!\left(-\frac{a^{2}}{2(n-1)\gamma h^{2}/4+\frac{2}{3}(h/2)a}\right).$
Since
$E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}}):=\Pr\Bigl[S_{\mathrm{disc}}>c\Bigr]\leq\Pr\bigl[|S_{\mathrm{disc}}|>c\bigr],$
we obtain the explicit upper bound.
(iii) Aliasing error $E_{\mathrm{alias}}$. Similarly, the aliasing error corresponds to the event that the sum of the discretized blanket variables leaves the outer FFT window. Using the selector representation,
$\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}\;\stackrel{{\scriptstyle d}}{{=}}\;\sum_{i=2}^{n}B_{i}Z_{i}^{\mathrm{di}},$
where the truncated discretized variables $Z_{i}^{\mathrm{di}}$ lie in the bounded interval $s,s+w^{\mathrm{in}}]$. Setting $Y_{i}:=B_{i}Z_{i}^{\mathrm{di}}-\mathbb{E}[B_{i}Z_{i}^{\mathrm{di}}]$ we obtain independent, bounded variables with $|Y_{i}|\leq K_{\mathrm{alias}}$ and variance at most $v_{\mathrm{alias}}^{2}$, both computable from the truncated blanket distribution and $\gamma$. Bernstein’s inequality applied to $S_{\mathrm{alias}}:=\sum_{i=2}^{n}Y_{i}$ then yields an explicit exponentially small upper bound of the form
$E_{\mathrm{alias}}:=\Pr\Bigl[\sum_{i=2}^{n}B_{i}Z_{i}^{\mathrm{di}}\notin[\,-w^{\mathrm{out}}/2,\,w^{\mathrm{out}}/2\,]\Bigr]\;\leq\;2\exp\!\left(-\frac{(w^{\mathrm{out}}/2)^{2}}{2(n-1)v_{\mathrm{alias}}^{2}+\frac{2}{3}K_{\mathrm{alias}}w^{\mathrm{out}}/2}\right).$
That is, the aliasing error can also be explicitly controlled.
In particular, all three error contributions $E_{\mathrm{trunc}},E_{\mathrm{disc}},E_{\mathrm{alias}}$ admit explicit, computable upper bounds in terms of the CDFs $F_{\mathrm{ref}}$ and the numerical parameters $(n,\gamma,s,w^{\mathrm{in}},h,w^{\mathrm{out}})$. Hence the total error terms $\delta_{\mathrm{err}}^{\mathrm{low}}$ and $\delta_{\mathrm{err}}^{\mathrm{up}}$ in Theorem 5 can be made as small as desired by a suitable choice of these knobs.
Here, we outline a systematic procedure for selecting the FFT parameters given $\alpha,\eta_{\mathrm{trunc}},\eta_{\mathrm{disc}},\eta_{\mathrm{alias}},\eta_{\mathrm{main}}$. By Lemma 3.2, we may write the blanket divergence in the form [^12] We use (13) to be more precise.
$D^{\mathrm{blanket}}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}=D_{n}(\varepsilon)\,\bigl(1+o(1)\bigr)\qquad\text{as }n\to\infty.$
Set
$\varepsilon_{n}=\frac{\alpha}{\chi}\sqrt{\frac{\log n}{n}}.$
Moreover, Lemma A.10 yields
$E_{\mathrm{main}}(c_{n})=\eta_{\mathrm{main}}\,D_{n}(\varepsilon_{n})$
for the choice
$c_{n}:=\frac{\sigma^{2}}{4\alpha^{2}}\,\frac{\eta_{\mathrm{main}}}{\log n},$
which is obtained by solving for $c_{n}$ from the leading term in (74). Note that $E_{\mathrm{main}}$ is not a numerical approximation error term (in particular, it is not included in $\delta_{\mathrm{err}}^{\mathrm{up}}$ or $\delta_{\mathrm{err}}^{\mathrm{low}}$), so we do not need to enforce a strict upper bound on $E_{\mathrm{main}}$; it suffices to match its leading term to the target scale $D_{n}(\varepsilon_{n})$.
We then choose the remaining FFT parameters in the following order.
  1. (1)
    Truncation window. First we fix truncation parameters $s_{n}$ and $w_{n}^{\mathrm{in}}$ and let
    $q_{n}:=\Pr\bigl[l_{\varepsilon}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})\notin[s_{n},s_{n}+w_{n}^{\mathrm{in}}]\bigr].$
    Using the exact expression
    $E_{\mathrm{trunc}}=1-(1-\gamma q_{n})^{n-1},$
    we choose $(s_{n},w_{n}^{\mathrm{in}})$ so that
    $E_{\mathrm{trunc}}\leq\eta_{\mathrm{trunc}}\,D_{n}(\varepsilon_{n})$
    for a prescribed relative budget $\eta_{\mathrm{trunc}}>0$.
  2. (2)
    Discretization step size. Next we fix a grid width $h_{n}>0$. By the Bernstein-type bound above, the discretization error satisfies
    $E_{\mathrm{disc}}\leq 2\exp\!\left(-\frac{c_{n}^{2}}{2(n-1)v_{\mathrm{disc}}^{2}+\frac{2}{3}K_{\mathrm{disc}}c_{n}}\right),$
    with $v_{\mathrm{disc}}^{2}\leq\gamma h_{n}^{2}/4$ and $K_{\mathrm{disc}}=h_{n}/2$. Hence
    $E_{\mathrm{disc}}\leq 2\exp\!\left(-\frac{c_{n}^{2}}{\frac{\gamma}{2}\,nh_{n}^{2}+\frac{1}{3}\,c_{n}h_{n}}\right),$
    We then choose $h_{n}$ so that this upper bound satisfies
    $E_{\mathrm{disc}}\leq\eta_{\mathrm{disc}}\,D_{n}(\varepsilon_{n})$
    for a prescribed $\eta_{\mathrm{disc}}>0$.
  3. (3)
    Outer FFT window. Finally, for the aliasing (wrap-around) error we have the bound
    $E_{\mathrm{alias}}\leq 2\exp\!\left(-\frac{(w_{n}^{\mathrm{out}/2})^{2}}{2(n-1)v_{\mathrm{alias}}^{2}+\frac{2}{3}K_{\mathrm{alias}}w_{n}^{\mathrm{out}}/2}\right),$
    where $v_{\mathrm{alias}}^{2}$ and $K_{\mathrm{alias}}$ depend only on the truncated blanket distribution and $\gamma$. Once $h_{n}$ and $w_{n}^{\mathrm{in}}$ are fixed as above, $v_{\mathrm{alias}}$ and $K_{\mathrm{alias}}$ can be computed explicitly from the truncated distribution. We then choose $w_{n}^{\mathrm{out}}$ so that
    $E_{\mathrm{alias}}\leq\eta_{\mathrm{alias}}\,D_{n}(\varepsilon_{n})$
    for a prescribed $\eta_{\mathrm{alias}}>0$.
C.2. Computation of CDFs
The algorithms 12 require as input the CDFs of the privacy amplification random variable. In the discrete case (e.g., $k$-randomized response), these CDFs can be computed by direct enumeration.
In the generalized Gaussian setting, the map $l_{\varepsilon}(y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$ is strictly decreasing on $\mathbb{R}$ whenever [^13] This restriction on the parameter range is not essential. When $y\mapsto l_{\varepsilon}(y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$ is not monotone, one can still represent the CDF of $l_{\varepsilon}(Y)$ as a finite sum of terms $F_{\mathrm{ref}}(b_{j}(u))-F_{\mathrm{ref}}(a_{j}(u))$ over the finitely many intervals $a_{j}(u),b_{j}(u)]$ on which $l_{\varepsilon}(y)\leq u$, where each endpoint $a_{j}(u),b_{j}(u)$ can be computed by one-dimensional root finding. However, in the moderate-deviation regime $\varepsilon_{n}\to 0$ considered in this paper, the monotonicity condition is satisfied for all sufficiently large $n$, so we impose it in the main text for simplicity.
$\varepsilon\leq\ln 2-\left(\frac{x_{1}-x_{1}^{\prime}}{2\alpha}\right)^{\beta},$
Under this condition the privacy amplification map is invertible, and for any reference distribution $Y\sim\mathcal{R}_{\mathrm{ref}}$ with CDF $F_{\mathrm{ref}}$, the CDF of
$X:=l_{\varepsilon}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$
is given by the simple composition formula
$F_{X}(x)=\Pr[X\leq x]=\Pr\bigl[Y\geq l_{\varepsilon}^{-1}(x)\bigr]=1-F_{\mathrm{ref}}\!\bigl(l_{\varepsilon}^{-1}(x)\bigr).$
Thus the only remaining technical task is to evaluate the inverse map $l_{\varepsilon}^{-1}$.
For generalized Gaussian mechanisms with shape parameters $\beta=1$ (Laplace) and $\beta=2$ (Gaussian), the inverse $l_{\varepsilon}^{-1}$ can be expressed in closed form, so that $F_{X}(x)$ is available analytically (in terms of $\Phi$ in the Gaussian case). For general $\beta\in(1,2)$, the inverse is no longer available in closed form, but it can be computed numerically by solving a one-dimensional root-finding problem for each $x$, which yields arbitrarily accurate values of $F_{X}(x)$ at a negligible cost compared to the FFT step.